A Complete Bibliography of The Econometrics Journal

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

25 May 2024 Version 1.09

Title word cross-reference

α [LC08]. b [YV11]. C_{p} [LO13]. F [OY06, Sen03, Sun13]. I(1) [Han05]. I(2) [Kon03, KNR11]. K [JP22]. L [GK18]. l_{0} [CL21]. L_{1,2} [CM18]. M [LHK15, Xu21]. N [HKR15]. R [LM22]. S [PS07]. T [DD12, QWfL16, WK22, Ard09, Die01, FHP10, MP08]. U [HH18].

-class [JP22]. -estimation [PS07]. -estimators [LHK15, LM22, Xu21]. -penalised [CL21]. -ratio [Die01, MP08]. -ratios [FHP10]. -Stable [LC08]. -statistic-based [HH18]. -statistics [GK18]. -test [OY06]. -tests [Sen03].

1 [Bår01]. 19 [Bil22, BP21, BM22, CCGV22, Cho20, Kor22, Sto22]. 1973-2003 [Nie08].

2 [BCMM23, DSBG22, Han22]. 2.2 [Oom99]. 2013Special [Smi16]. 2018 [Ano21c]. 2021 [Abb23b]. 2022 [Abb23a]. 2LASSO [GMMS19]. 2nd [Sil11, Wil11].

4 [ÖD11, PY11].

500 [QP13].

ability [GM04]. absence [Kom13]. academic [CLW15]. Accelerated [LY20, ZfL04]. Accuracy [HHK04, Dah02, Sto22]. ACD [Wil09]. achievement [EH08]. achievements [CLW15]. activity [CMKSW01]. Adaptive [BZ18, CHS09, Rag05, GS18, SZW10, WS05]. additive [AYZ14]. adequacy [Wri10]. ADF [XP98]. adjusted [PUY08]. Adjustment [KP98, DO04, Fan06]. adjustments [BR04]. adolescents [LPR17]. Advances [CLLT14]. affine [FP12, Rag05]. against [BT02, GW12, HIN22, KK05, LKSN03]. aggregate [DHK23, GMMS19, JM05]. aggregated [Cho06, Haf09]. agricultural [KN20]. al [Sch12a]. Algorithms [Rea22, EKS08, KL09, KSD99, Oom99]. Alpha [Han22]. AlphaGo [Iga20]. alternative [All07, AC02, BSJ02, DP06, GW12, Kiv13, KN08, LN99, Sen03]. alternatives [BMRSEA21, HIN22]. among [LPR17, Rün04]. analyse [DP06]. Analysis [HLN01, Wil11, BFHM08, CCG07, CIM^{+}20, Cho23, DGSS22, DV07, Emv20a, Emv20b, FHL^{+}22, Han22, JMN00, JK15, Kon03, MY00, MN11, Nie04a, Ulr08, YV11]. Anderson [DM14]. Anderson-Rubin [DM14]. Andrews [NS12]. Annual [Abb23b, Abb24, Ano19, Ano21c, Ano23, FS09, PS16, PS11a, Smi15, Smi16, Smi17, Smi18]. apparent [KP01b]. Application [DT06, AH04, AT04, BHKY17, BMRSEA21, CLTZ04, iSDLB05, CL22, CLW15, CSY23, DGSS22, DRB07, Frö07, GvV20, GS18, GL07, IMY24, MT99, SHD99, DHT09, Sto22, WW19]. Applications [BT13, XP98, BLLW11, DHB12, Gal17, LM22, QP13]. Applied [PS11a]. approach [Abr13, AS22, BG19, BM17, Can11, FK21, GY22, HP10, Han99a, HP99a, Ino06, KS10, MNT03, PY11, RVH05, TW19, YP01]. approaches [KN20, Woo23]. approximate [AD15, BKG^{+}22, LP05]. Approximation [Lar98, KP05]. approximations [FVFW21, HHK04, PV01]. arbitrage [CDR09]. arbitrage-free [CDR09]. arbitrariness [Das05]. Archimedean [Smi03]. area [ACMG^{+}11]. ARFIMA [GM04, May07]. ARIMA [MCM23]. ARMA [KK03, LV22, Med03, NS12, XP98]. Artificial [Iga20, LR05, BK03]. Assessing [PS09]. assessment [ALP^{+}21, Kau00]. asset [KRH^{+}09]. assuming [Kru21]. assumption [Frö07]. assumptions [MP23]. asymmetric [Ard09, AM07, Kaw07, Kom13]. asymmetries [DV07]. Asymptotic [BKL08, Cam16, DM04, Gos04, LL23, MP08, PV01, Tak08, Das05, DP06, HH18, OY06]. asymptotically [Myn11]. Asymptotics [Cav03, YZ13, AD15, LM22, Wri10]. attitude [CSY23]. Attractive [vdS98]. auctions [HK19, Kom13]. Augmented [ZW24, AT11, HW07, KW21, Mao16, YC22]. autocorrelation [Sun13]. automatic [CMKSW01, MFB03]. Autoregression [Xu08, CS07b, EN12, Hua08, KPR17, PA23]. Autoregressions [KY21, Pro98, RR14]. Autoregressive [Ano21b, CK98, Lar98, BW02, DO04, DLS09, EN09, GM00, HH00, JL01, JS04, KMVW02, LMNS21, NV02, PP04, San08, ST06, WL13]. average [DHM23, HM09, KW21, SZW10, SC21]. average-augmented [KW21]. averages [MW23]. Averaging [KY21, KP04, LS21, LO13, LK16, ZZMZ21]. away [BM17].

B [PY11]. back [PT11]. Backfitting [LMP14]. backward [Eve13]. band [NF11, YP13]. bandits [KT23]. bandwidth [CCF20]. bank [Büt99, CSY23]. based [AD13, AH21c, AS22, Ble15, BB11, Dah02, Das05, DM14, DFGK98, EKS08, FZ14, GOS06, HH18, HC19, Ioa05, JK15, Kal12, Kaw07, KM04, Kna22, KY15, LLL01, MS98, PR16, SY22, Sør00, Tse02, ÖD11]. basis [BK03]. Bayesian [Mar12, Ard09, BL98b, BR07, DV07, GY22, GFCK09, Kau00, KWTZ18, KP04, Koo10, MNT03, MY00, MFB03]. BCG [BP21]. bedtime [LPR17]. Behaviour [LN00, Fan06, Gre04, Juh05, LN99, Sen03]. BEKK [PR14]. better [CCGV22]. Between [Sen98, BK19, GL21, HK01, fLY20]. Beyond [vKR18]. bias [AD15, BKOV20, CCF20, GvV20, HHK04, HB20, PUY08]. bias-adjusted [PUY08]. bias-corrected [CCF20, HB20]. biased [vBRV24]. Bimodal [FHP10]. binarised [AH21c]. Binary [CL21, BK13b, Ble15, Cha17, CR14, CS22, FJ19, Frö06, HMY11, HdP21, JM17, JPX16, MZ18, RVH05, Tho06]. binary-choice [CS22, Tho06]. bivariate [CLTZ04]. block [Ioa05, MTV17, PP02]. block-wise [MTV17]. Blockwise [Bra09]. BLP [GMMS19]. BLP-2LASSO [GMMS19]. Blue [Iga20]. Bonanza [Iga20]. Bootstrap [DM08, MP05, BZ18, CO12, Cam16, Du16, HB20, HH18, Ino06, Ioa05, Kap08, MW18, PP02, Tch15]. Bootstrapping [Xu08]. both [GS13]. boundary [NR24]. Bounding [Sto22]. Bounds [AC02, BK22, BH07a, Gar20, Lec99]. Box [Shi08]. Break [Yan12, BCCW08, iSS06, Cho03, HK01, KY15, LN00, Yan02, ZSS20]. break-point [BCCW08, Cho03]. Breakdown [CO12]. Breaking [iSDLB05]. breaks [GM04, HLRZ22, Ioa05, JMN00, Kim14, Pit04]. bridge [FS04]. brief [CK16]. Bubble [WY23]. bubbles [BCK20, EN12]. BUGS [MY00]. business [VST03, Kau00].

C [PY11]. capita [iSDLB05]. capital [Büt99]. care [DT06]. Carlo [KLS21, Par02, SHD99]. case [BKL08, CCGV22, Cho20, DSBG22, PP04]. Causal [FHL^{+}22, AI22, HS04, KLS21, SC21]. Causality [Haf09]. cautions [BMO04]. CCE [MW23, Wes18, WK22]. censored [Bri11, HS18, LG21, ZL15]. censoring [BTV24]. Central [DSW18]. Change [AHHK06, BHKY17, BP03, DW19, DP06, VST03, KN20, Kri12, LKSN03, LR06, RWW23, Xu15, vBRV24]. Change-point [AHHK06]. changes [Hog17, YP13]. changing [WX18]. characteristics [KjL03, SS24]. Chetan [Gör12]. China [MW08]. Choice [Abb23b, BK13b, BC10, CCF20, CS22, GMMS19, HMY11, KR03, MZ18, Tho06, ZfL04]. Choosing [MP23]. class [AD15, DV07, FK21, JP22]. classes [Cam16]. classification [CL21]. Cliff [MP11]. climate [KN20]. Clive [Osb12]. closed [Han05]. closed-form [Han05]. clusters [MW18]. co [GS13]. co-trending [GS13]. Coefficient [XP98, Che19, GL21, Juh05, LMP14, LCG11, RPS14, ZLS23]. coefficients [CR14, DHK23, Zam02]. coexplosive [EN12]. COGARCH [HKLZ07]. coherency [BK19]. Cointegrated [BL98a, CL22, CPP01, HJ99, JS04, Nie08, Pét00, PS01]. cointegrating [CS04, DLS09, HP04, LST01, Mad05, RG06]. Cointegration [Cha11, JMN00, Nie04a, RM99, BiS13, EM02, HKR15, JLLN08, Kon03, KNR11, LLL01, NF11, ÖD11, PP04, PR16, Qu07]. collection [CLW20]. collinear [Myn11]. collinearity [MJ19]. combination [DSBG22]. Combining [MCM23]. commodity [CL22]. Common [Kim14, BiS13, HLRZ22, ZSP12, ZSS20]. comparative [BSJ02, Tak08]. Comparing [KZ23, KN20, TW19]. Comparison [CK98, DP06, Joc13, Mad10, Pal22, PP04]. competing [CCG07]. Complete [LS21]. complex [ZW24]. complier [SS24]. component [BT02, BSJ02, Cho23]. components [BT02, HK00, NV02, Smi08a]. composite [JJS14]. compromised [HK21]. computation [MFB03, ST21]. Computationally [vdS98, GL07]. Computing [DHS21]. concave [LY20]. concentration [PS09]. condition [DSW18, PV01]. Conditional [NT09, AS22, AM07, BR07, Ber22, Bra12, CLW15, CHS09, DSW18, Frö07, FHP21, GK18, GM00, GL07, Hsu17, HS17, JL01, KMVW02, KN08, Mof01, SC21, Tse02, Zha13]. Conditionally [Sen98]. Conditions [SHD99, Bra09, HLT08, fLY20, Pal22, YZ13]. Conference [Abb23b, Abb24, Ano19, Ano21c, Ano23, FS09, PS16, PS11a, Smi15, Smi16, Smi17, Smi18]. Confidence [DM14, KY15, Gos04, HB20]. consequences [DO04]. Consistency [DD02, BCCW08, Cho03]. Consistent [GS13, Hsu17, Tho06]. constancy [HJ99]. Constant [NT09, DW19, JS04]. constrained [DHS21]. constraints [Sca16]. Constructing [NS24]. Constructive [CE99]. consumer [CE99, Van10]. consumption [EJN02]. contagiousness [Han22]. Continuous [CDF98, EKS08, Lec99, PY11, YP01, YLW23]. continuous-time [EKS08]. contract [GvV20]. contrasts [IRS20b]. contribution [Han99a]. contributions [Phi15]. Control [CDF98, BF19, CLW20, GY22]. Controlling [GO00, GL07, GS18]. controls [Joc13]. conventional [Wri10]. Convergence [MW08, Ble15, SHD99, iK07]. convex [GY22, Kuo08]. convexity [Sca16]. cooperation [ALK23]. copula [CLTZ04]. copulas [JM05, Smi03]. corrected [CCF20, HB20]. correction [AD15, BR04, CM24, DRB07, EM02, KMT03, LP09]. correction-volatility [LP09]. correlated [AH21a, BiS13, BKL08, MPP14, PP04, Xu14]. Correlation [NT09, DW19, GL21, ST21]. Corrigendum [Ano05, ÖD11, PY11]. cost [Fan06]. count [AT04, Kim20]. counted [CLTZ04]. counterfactual [MCM23]. countries [CMKSW01]. country [Kon03]. Counts [RVH05, CLTZ04]. Course [Koo10]. CoV [BCMM23, DSBG22, Han22]. covariance [ADXBA21, All07, DUM24, FLL16, MP05, WZQ07, ZZMZ21]. covariate [CL21, YZLC21]. covariate-dependent [YZLC21]. Covariates [BLV08, BF19, DHB12, GMMS19, HH10, JW19, KR23]. COVID [BP21, BM22, CCGV22, Cho20, Kor22, Sto22, Bil22]. COVID-19 [BP21, BM22, CCGV22, Cho20, Kor22, Sto22, Bil22]. Cox [BM06, Shi08]. crash [Sen03]. crashes [BCK20]. criteria [MW23]. criterion [AT11, Ble15, HP04]. criterion-function-based [Ble15]. Critical [BP03, San08, WZO06]. Cross [OLL06, Wil11, CPT11, HKR15, HLRZ22, Hua08, Kap08, KW21, Mad05, Mao16, PUY08, PS03, SR09, DHT09, SRPR22]. cross-dependent [HLRZ22]. cross-section [CPT11, HKR15, KW21, PUY08]. cross-sectional [Hua08, Kap08, Mao16, SR09, DHT09, SRPR22]. Cross-validation [OLL06]. curve [DF99]. curved [BC14]. curves [BHKY17, KR13]. cycle [VST03]. cycles [Kau00, OTP04, Rün04]. cyclical [BK19].

D [HK99]. Dag [Osb12]. Danish [Nie08]. Data [GT99, Han99b, HP99a, Abr13, AH21a, AYZ14, AT04, AS19, AB16, AI22, BFK11, BMO04, BG19, BW02, BC07, BRS16, BW10, CE99, Can11, CLW20, Cha17, Che19, CH21, CNSY22, DD12, DSBG22, DRB07, Eve13, GvV20, Had00, HL05a, HMY11, HdP21, HS18, Joc14, Kap08, KZ23, Kim14, Kim20, fLY20, LG21, LCG11, LPR17, Mad10, OY20, PR16, Rod02, RPS14, San08, SRPR22, Tho06, WL13, WC20, Woo23, ZSP12, ZW24, Han99a, HK99, Wil11]. date [KY15]. Dave [Gör12]. David [Gör12]. Davidson [Sch12a]. day [AM07]. days [CG05]. Dealing [KK11]. Debiased [CNS22, SC21, Cha20, CCD^{+}18, LZZ21]. decision [HK21]. decision-making [HK21]. decisions [LPR17]. decomposition [CL22]. deep [KN20, Iga20]. Defaults [BLV08]. defiers [DHM23]. Degrees [KP98]. DeJong [Gör12]. Delta [Han22]. demand [Agu23, DHK23, Nie08, Pét00]. Denis [Abb23a, Abb22, Ano11a]. densities [BH07b, LW17, WS05, iK07]. density [JK10, Tak08, YLW23, iK07]. dependence [BY13, BK19, BB11, CJKL17, CPT11, HKR15, Hal08, Hua08, OTP04, PS03, SR09, DHT09, SRPR22]. Dependent [MS98, Sil11, BKOV20, Frö06, Gre04, HLRZ22, LK05, LX17, Mag10, Smi08a, WH07, WL13, YZLC21]. derivative [AN08, SZW10]. derivatives [HM09]. Designed [BKG^{+}22]. designs [CCF20, HB20, KR23, SY22]. Detecting [HLRZ22]. detection [Kri12]. Determination [HP04, Kon03]. Determining [FT00, LP09, ST06]. deterministic [AT18, DLS09, GS13, JMN00, NV02]. deterministically [CPP01]. detrended [PR16]. developments [Rob10]. deviations [Ots11]. diagnostic [MV07]. diagnostics [Tse02]. Dickey [LN99, LN00, LV06, Sen03]. difference [Cha20, CH01, LKSN03, LP05, OM22, Woo23, Ano23]. difference-in-differences [Cha20, Woo23, Ano23]. difference-stationarity [LKSN03]. difference-stationary [CH01]. differences [CLTZ04, Cha20, GK18, Woo23, Yan12, dCD23, Ano23]. differences-in-differences [dCD23]. differencing [FT00, Joc14]. different [AS19, Mad10]. differentiable [LHK15]. differential [Nic02]. differentiated [Agu23, DHK23]. differentiation [MFB03]. diffusion [SHD99]. diffusions [Hua11]. Dijk [Mar12]. dimension [HL05a, Mad10]. Dimensional [PS16, HLRZ22, JW12, KR23, PA23, SHR14, UT19, ZZMZ21]. dimensions [BC14]. Direct [RPS14]. Dirichlet [GL07]. Dirichlet-multinomial [GL07]. disaggregated [ABS03]. disaggregation [MS05b, Pro06]. discontinuity [CCF20, HB20, KR23]. discontinuous [BR04]. discovery [GS18]. Discrete [Abb23b, JPX12, KR03, ALK23, BC10, GMMS19, LK05, Xu14, ZfL04]. discretely [Hua11]. discussants [HP99b]. Discussion [Han99a, Han99b, Sch12a, GT99]. Disentangling [BCMM23]. dissimilarity [ABDW15]. distance [KS10, May07, Win19]. distancing [Bil22]. Distinguishing [HK01, PST08, CM07]. distributed [PP04]. Distribution [ABS03, DGSS22, Lar98, Dan05, Die01, Geo07, HL00, LR06, Mag02, OY06]. Distributional [JP22, ALP^{+}21]. Distributions [EM02, LC08, KRH^{+}09, Sch12b, WW19]. Disturbance [KP98]. disturbances [KK05]. do [Ros21]. dominance [LPW14]. dominances [BLLW11]. Donald [Sch12a]. Double [Cha20, CCD^{+}18, Kna22, LZZ21, SS24, BHL22, FHL^{+}22]. Double/debiased [Cha20, CCD^{+}18, LZZ21]. Doubly [AI22]. drawn [LR06]. drift [EKS08]. driven [Ver11]. DSGE [CDH^{+}21]. DTARCH [HJ05, JJS14]. due [KP01b]. dummies [KjL03]. dummy [MZ18, vGS02]. duration [DGSS22, GM00, KN08, OTP04]. durations [Bri11]. Durbin [KK05]. during [CCGV22, Cho20]. Dynamic [Abb23b, Agu23, Fan06, KP98, PS03, Pro06, Aba23, BiS13, BH16, BHL22, BCMM23, Bra09, BH07a, BC10, BW10, CL22, sCK10, CK12, CSY23, Dem02, Eve13, HW07, HS18, Isk10, JK15, Kau00, Khe15, KP05, KP04, Kru21, fLY20, LG21, OP23, PP04, Qu07, QWfL16, RV02, SR09, SHR14, ZfL04]. dynamics [DMN20, HP10, HH00, MN11, OY20].

early [Sto22]. earnings [AH04, Ulr08]. earnings-equation [AH04]. East [Lec99]. EC [AM01]. ECF [JK10]. Echelon [BL98a]. Echelon-form [BL98a]. Econometric [BJN03, BU09, CPP01, CK16, KN20, Wil11]. Econometrics [AdP17, Abb22, Abb23a, Abb23b, Ano10, Ano11a, Ano11b, Ano12b, Ano13, Ano14, Ano15, Ano16, Ano17, Ano18, Ano19, BT13, CLLT14, CD09, FS09, GM98, HS98, Mar13, NS01, ÖD11, PY11, Smi08b, Smi15, Smi16, Smi17, AM02, Gal17, IRS20b, Phi15, Sen09, Mar12, Sil11]. Economic [Abb23b, Abb24, Ano19, Ano21c, Ano23, EM98, FS09, Gia15, HS98, MW08, Osb12, PS16, PS11a, Ros16, Smi08b, Smi15, Smi16, Smi17, Smi18, BK19, CMKSW01, KP01b, MNT03, NS24, OTP04]. Economy [XP98]. Ed [Sil11, Wil11]. Edgeworth [HV16]. edited [Mar12]. Editorial [AdP17, Abb22, Ano21a, IRS20a, BT13, LS12]. Editors [GM98, NS01].

Edmond [Phi15]. effect [BCMM23, CG05, GS18, HP10, JW12, Kor22]. Effects [Bil22, AH21a, Agu23, AYZ14, AS19, AM07, BKL08, BFK11, BK22, BHL22, BC07, BM22, CLW15, Cha17, CNSY22, CSY23, DHM23, DDF09, VST03, Gar20, Gre04, GFCK09, HMY11, HLL20, Hsu17, KT18, KLS21, Lec99, LCG11, LPR17, LMW20, MP08, OY06, RPS14, Ros17, Ros21, SC21, SY22, Tug21, YC22, YZ23, ZSP12, dCD23]. efficiency [BH07a, DSW18, DG11, Emv20a, Emv20b, LPW14, Sen09]. Efficient [Nie04b, vdS98]. EGARCH [KK03]. eigenvalue [LST01]. elasticity [vBRV24]. eliminating [BKOV20]. elliptical [BH07a, KRH^{+}09]. emending [HKR15]. emerging [Han22]. Empirical [KLS21, XP98, Ber22, Bra02, Bra09, Bra12, CL21, EN09, HK01, Ma17, Ots11]. employment [Lec99, PS01]. encompassing [Han99a, HP99a]. endogeneity [AS22, HSW15, KK11, TD14, Van10]. endogenous [JPX12, JPX16, Kan14, KL09, Kru21, MZ18, QWfL16, RVH05, Sil03]. England [BCMM23]. entropy [WS05]. epochs [RWW23]. equalities [PT11]. equation [AH04, DM08, Kiv13]. equations [DRB07, LK05, fLY20, Nic02, PS09, ZW24, vGS02]. equicontinuity [Hag14]. equilibria [Kas15]. Equilibrium [OP23, BR04]. Errata [Ano09]. Erratum [Ano12a, Ano21b, Emv20a]. error [AH04, BSJ02, BY13, CM24, EM02, GO00, HSW15, KMT03, LP09, Mao16, PP04, PUY08, SR09, Smi08a, WZO06, WC20]. errors [ABS03, BCCW08, BKL08, Cho03, DD12, Du16, LY20, MS05a, MPP14, NS12, PS17, WH07]. errors-in-variables [WH07]. estimated [DM08]. estimates [Zam02]. Estimating [BL98a, BH07b, Che19, CG05, DSBG22, FP12, FS04, HWY08, JW12, JPX16, Mad05, Mon98, YP13, YG02, HP10, LHK15, Ros21, Sør00, WZO06, ZW24]. Estimation [AYZ14, BK13a, CS07b, CSW22, CM18, CK12, DUM24, Dan05, HHK04, HH10, Kam98, Kan14, KN08, LG21, LX17, LC08, Mag02, McK01, PA23, SY22, WL13, Xu14, ADXBA21, Abr99, AH04, AS11, All07, Ana13, AH21c, Ard09, AB16, BKG^{+}22, BK13b, Bre02, CDH^{+}21, CLS14, CPT11, CHS09, CS22, CM24, CNSY22, CSY23, DT06, DRB07, EKS08, FG03, FPR15, FLL16, FF10, GH06, Gør06, GFCK09, HKLZ07, HM09, HH18, HSW15, HS18, Hua11, HC19, Iga20, JK10, JJS14, Joc13, KMT03, Kaw07, KK11, KLS21, KS10, KjL03, Kor22, Kri12, Kru21, fLLL10, LG08, LV22, MP05, May07, MTV17, MZ18, MT99, NF11, NR24, OY20, OP23, OLL06, Par02, PS03, Pit04, PS07, PS17, QWfL16, RG06, RPS14, SR09, SZW10, Shi08, Sil03, Spe09, Tho06, Van10, Ver11]. estimation [WH07, WC20, WS05, ZfL04, Zha13, ZL15, ZZMZ21]. estimator [AN08, BW10, CH21, Cho03, Esc18, Gre04, LL23, Mag07, OM22, ST21, Sun13, iK07, vBRV24]. Estimators [KP98, AD15, BCCW08, BSJ02, BKL08, BU09, Ble15, BW02, BC07, Cam16, DD02, DM07, HH18, HV16, HW07, JP22, KP05, LHK15, LP05, LM22, Ots11, PP04, Tak08, WZQ07, Xu21, Yan12]. euro [ACMG^{+}11]. European [CMKSW01]. Evaluating [BHL22]. evaluation [GS18, Kna22, MCM23]. events [BM22, LG21]. evidence [Akr04, BM22, HK01, KLS21]. Exact [McE08, ST21, vGS02, JS04, Van10]. example [FF10]. excessive [YG02]. excessive-zeros [YG02]. exchange [Akr04, JLLN08]. excluded [Tsy23]. exclusion [Ada16, AH21c, LMW20]. exogeneity [FHP21, MP23]. exogenous [Ana13, Esc18, vKR18]. expansion [RVH05]. Expansions [LP05, HV16]. expectation [BMRSEA21]. Expectations [Ros07, BFHM08, JS04]. expenditure [ABS03, CE99]. experiment [BP21, BM22]. experimenter [HK21]. experiments [HK21]. Explicit [Aba23]. Exploring [MNT03]. explosiveness [GSW19]. exponential [BC14, BM17]. expression [Han05]. extended [MFB03]. extent [ALP^{+}21]. extraction [HK00]. extreme [Hog17, TfL14]. extremum [Cam16, FPR15].

Factor [BFHM08, PS11a, Sen98, YC22, AT11, BHKY17, BH16, CL22, Cho23, DUM24, Geo07, HS04, JK15, Kau00, Kim14, KP04, LP09, MN11, SHR14, VDP03]. Factor-augmented [YC22, AT11]. factors [BiS13, LP09, Wes18, ZSS20]. failure [CH02, EJN02, LY20]. false [GS18]. families [BC14]. family [Liu09]. far [BM17]. fat [JL01]. fat-tailed [JL01]. fatality [DSBG22]. Feasible [Cho23, Tsy23]. female [DRB07]. few [MW18]. field [BM17]. fields [Dah02]. filter [Lin05, McE08, MFB03]. filtered [DW19]. filters [Kaw07]. Financial [FS09, GL21, GL09]. findings [KP01b]. Finite [Ada16, DFGK98, KK05, HL05a, HK21, HLL20, HL00, DHT09, Xu21]. finite-population [HLL20, Xu21]. Finite-sample [KK05, HK21]. firm [BKOV20]. First [Joc14, Koo10, AGMR23, DSBG22, KM04, Yan12]. First-differencing [Joc14]. first-stage [AGMR23]. Fiscal [AM01]. fit [WW19]. Fixed [YV11, Agu23, AYZ14, BFK11, BC07, Cha17, Gre04, HKR15, HMY11, LCG11, Mad10, MP08, RPS14, Tug21, WK22, ZSP12, dCD23]. Fixed- [YV11, WK22]. fixed-effects [Agu23]. flash [BCK20]. Flexible [CLLT14]. fluctuations [VST03]. football [BM22]. Forecast [CK98, KY21, CH02, EJN02]. Forecasting [CH01, EM98, KW21, KP04, NS01, NV02, Ros16, Smi15, AM01, BCK20, CMKSW01, GM04, Gia15, Haf09, IMY24, JK15, UT19, YC22]. forecasts [ACMG^{+}11, HLN01, HC04, Pal22]. Foreword [HS98]. form [BL98a, Han05, Kor22, YT08]. formation [PS01]. formulas [McE08]. formulation [HK00]. forward [Fan06]. forward-looking [Fan06]. Fractional [LV06, LP05, LV22, NF11]. fractionally [KK05, Nie04b]. Fragility [FJ19]. Framework [EM98]. frameworks [DP06]. free [ALP^{+}21, CDR09]. Freedom [KP98]. frequency [Cha11, GH21, NS24]. friend [BM17]. frontier [Smi08a, TK04]. full [VDP03]. full-factor [VDP03]. Fuller [LN99, LN00, LV06, Sen03]. Fuller-type [LN00]. Fully [NF11]. function [BK03, Ble15, EJN02, GW12]. Functional [Juh05, BHKY17, YT08]. Functional-coefficient [Juh05]. functionals [ZW24]. functions [BK13a, BH07b, CS07b, Gør06, GM00, Joc14, LHK15, LG08, SC21, Sør00, YG02]. Further [WZO06, GOS06]. fuzzy [HB20].

G [Sch12a]. game [LX17]. games [ALK23, KWTZ18, OP23, Xu14, Ano19]. gap [Frö07]. GARCH [Ard09, BL98b, BPR10, DV07, DG11, Haf09, HWY08, LS05, Liu09, LM22, LR05, NT09, PR14, PS17, RWW23, Sen98, VDP03]. GARCH-in-mean [LR05]. Gary [Mar12]. Gaussian [Emv20a, PY11, Emv20b, Kaw07, YP01]. GDP [ACMG^{+}11, Bha01, CMKSW01, iSDLB05]. gender [Frö07]. general [BK19, Han99a, HP99a, PP02, Wes18, YZ13]. general-to-specific [Han99a, HP99a]. generalised [HP04]. Generalized [BH16, Bra12, KY21, SHR14, YT08, Bra09, Cav01, CDR09, Das05, Du16, KRH^{+}09, Ots11, PR16, WC20, Zha13]. generated [CLS14]. Generic [BCCW08, Cho03]. German [BM22]. Germany [BP21, DSBG22, Lec99]. Geweke [Mar12]. Gibbs [BL98b]. Gini [YLW23]. given [LPW14]. global [CNS22]. GMM [BW10, GW12, Wri10, vBRV24]. GNP [CK98]. goodness [WW19]. goodness-of-fit [WW19]. Granger [Osb12, Han05]. graph [BM17]. Graphical [Mof01, CM18, MNT03, TW19]. gravity [YZ23]. Great [Gar20]. group [Che19, DDF09, SS06]. grouped [GL07]. Growth [MW08, ACMG^{+}11, CMKSW01]. guide [HK21]. guided [WW19].

Handbook [Mar12]. Hausman [MP11]. hazard [Gør06, GM00]. hazards [Bri11, Gør06]. health [DT06, Hal08, Isk10]. heavy [Dav12, PS17, Sch12b]. heavy-tailed [PS17, Sch12b]. Hedonic [KjL03]. Herman [Mar12]. Heterogeneity [BC10, Hal08, Smi16, Agu23, HK19]. heterogeneous [Had00, HL05a, KLS21, LLL01, LMW20, OY20, ÖD11, SS06, Tho06, WK22, dCD23]. heteroscedastic [GFCK09, MZ18]. Heteroscedasticity [LO13, CD22, CHS09, Tse02, YT08, ZL15]. Heteroscedasticity-robust [LO13]. Heteroskedastic [Sen98, JL01]. Heteroskedasticity [Ano21b, BR07, GOS06, HH00, KMVW02, LMNS21, Sun13]. Hidden [BLV08]. hierarchical [MN11]. High [UT19, HLRZ22, JW12, KR23, NS24, PA23, SHR14, ZZMZ21]. High-dimensional [UT19, JW12, KR23, PA23, SHR14, ZZMZ21]. Higher [DMN20, AD15, HHK04]. Higher-order [DMN20, AD15, HHK04]. highly [LS05, San08]. Hildreth [Zam02]. hitting [CG05]. Hodrick [McE08]. homework [EH08]. homogeneity [BRS16, PS03]. Honour [CLLT14]. Hoover [Han99a, HK99, GT99]. Horizons [Ros07]. Horowitz [CLLT14]. Houck [Zam02]. household [Pét00]. housing [MN11]. huge [MTV17]. hull [GY22]. hyperbolic [AH21b]. hypergeometric [BH07b]. hypotheses [BMRSEA21, Ros07]. hypothesis [iSS06, Cav01, LHK15, LN99, LN00, Psa01].

Identification [Ano21b, BT13, DDF09, FK21, HS04, HM09, HdP21, HSW15, HS18, Kiv13, KT18, Kru21, Man13, MZ18, Ros17, TD14, ZL15, AI22, Bri11, CM24, DHK23, FJ19, Kim20, Kom13, LMNS21, Mag10].

Identification-robust [TD14]. identified [FHP21, Rea22]. Identifying [GvV20, Ros21, vBRV24]. ignorant [Koe20]. ill [Sca16]. ill-posedness [Sca16]. IMF [AM01]. impact [AM09, Cho20, EH08, FHP10, KN20, SR09]. imperfect [BK22]. implication [Aba23]. implied [BH07b, CO12]. import [Kon03]. importance [ZfL04]. Improved [RR14]. Improving [sCK10, HK99]. impulse [BK13a, CS07b, Gos04]. imputation [NS24]. incidental [Joc14]. income [DMN20, Lec99, WW19]. inconsistency [Mag07]. Incorporating [DHB12]. increasing [BC14]. independence [Du16, Frö07, Kom13, KS10, KjL03, Mao16, PUY08]. Index [Ano10, Ano11b, Ano12b, Ano13, Ano14, Ano15, Ano16, Ano17, Ano18, CD09, AS11, ABDW15, CS22, CM24, FK21, Gør06, HSW15, KjL03]. indicator [CMKSW01, Isk10]. indicators [NS24]. Indices [BL98a, GL21, QP13, YLW23]. Indirect [CDF98, FZ14, KPR17, LC08, Mon98, AD15]. individual [AS19, BKL08, OY06]. industrial [VST03]. inequality [DG19, DHB12, DHS21, KZ23, PT11, Sch12b]. infection [BM22, DSBG22, Sto22]. infections [BCMM23]. Inference [BL98b, CDF98, CLLT14, DF99, HLL20, KR23, Mag10, MS98, Mon98, Smi18, AD13, ABDW15, AC02, Ana13, AD15, BK13a, BR07, BPR10, BC14, Bra09, CCF20, Dah02, DM08, Dav12, DM04, FHP10, FZ14, GK18, HK21, HM13, HJJX21, Kas15, Kiv13, KWTZ18, KPR17, LX12, Mad10, Nie04b, Rag05, RM99, Rea22, Sch12b, SY22, TD14, Xu21, YLW23]. Inferential [DG19]. Inflation [JLLN08, BJN03, IMY24]. influence [LX17]. Influential [Nie08]. Information [BKOV20, AT11, Kiv13, Mag07, MW23]. infrequent [GM04]. inhomogeneous [CHS09]. Initial [fLY20, HL05b, HLT08, LR06, PV01]. innovations [Ard09]. instability [GL21, KP01b, LX12]. institutional [VST03]. Instrument [AH21c, BW10, LX12]. Instrument-based [AH21c]. Instrumental [Ana13, BTV24, FJV12, QWfL16, RG06, AGMR23, CR14, CNSY22, DM08, FHP21, Kim20, Sca16]. instruments [BK22, CD22, Esc18, HHK04, LS21, Tch15, Tsy23]. insuring [Büt99]. integrals [BKG^{+}22]. integrated [BMO04, CJKL17, Gos04, HW07, KK05, Med03, Nie04b, PV01]. integration [Kaw07]. integration-based [Kaw07]. intelligence [Iga20]. intensity [BCMM23]. interaction [fLLL10, TfL14]. interactions [Man13, NT09]. interactive [CNSY22, CSY23, Tug21]. Interest [CDF98, PY11, YP01]. international [Bha01]. interpretation [vGS02]. interquantile [GK18]. interval [Bri11, HH10]. interval-censored [Bri11]. intervals [Gos04, HB20]. interventions [Cho20, Kor22]. intraday [KR13]. Introduction [GM98, NS01]. invariance [Pal22]. inverse [AH21b, FF10]. inverses [Das05]. inverting [DM14]. investigation [Dah02]. irrelevant [LG08]. Issue [AdP17, Abb23b, Ano19, Ano21c, Ano23, CLLT14, FS09, PS16, PS11a, Smi15, Smi16, Smi17, Smi18]. issues [AH21c]. Italy [CCGV22]. IV [CD22, CSY23, DM07, GW12, Tsy23].

J [HK99, ÖD11, Osb12, PY11]. jackknife [CD22]. jae [Sil11]. Jeffrey [Wil11]. JIVE [DM07]. job [KP01a]. Joel [CLLT14]. John [Mar12]. Joint [iSS06, KK11]. Journal [HS98, ÖD11, PY11, Smi08b, Ano10, Ano11b, Ano12b, Ano13, Ano14, Ano15, Ano16, Ano17, Ano18, CD09]. jumps [Rag05].

K-nearest-neighbour [LG08]. Kalman [Lin05, MFB03]. Kernel [OY20, DD02]. kernels [BNHLS09]. Kingdom [Frö07]. kink [YZLC21]. known [Dan05, Mag02]. Koop [Mar12]. KPSS [HL00]. Kronecker [BL98a]. kurtosis [BU09, Wil09].

L [CLLT14]. labour [HP10]. Lag [HW07, Yan02, PP04]. Lag-augmented [HW07]. Lagrange [BRS16, RR14]. land [GvV20]. land-lease-contract [GvV20]. Large [BM22, GH21, Ots11, PS16, CPT11, CSY23, DUM24, FLL16, FP12, Kim14, LX17]. Large-scale [BM22]. Larsson [ÖD11]. LATE [DHM23]. latent [Che19, CK12, HM13, Isk10, KZ23, LNN09]. leading [CMKSW01]. learned [Ros21]. learning [BHL22, Cha20, CCD^{+}18, CNS22, FHL^{+}22, IRS20b, KLS21, Kna22, LZZ21, NSS22, SC21]. learning-based [Kna22]. lease [GvV20]. Least [Pit04, PS17, HW07, KP05, Kuo08, NF11, PR16, WZQ07, Win19]. Least-squares [PS17, KP05, PR16]. Lecture [Abb24]. Lee [Sil11]. length [Yan02]. lessons [Gia15]. level [Psa01, QP13]. levels [GO00, Yan12]. leverage [AM09]. Lévy [Ver11]. Lévy-driven [Ver11]. Likelihood [JK15, Kam98, LLL01, MS98, Ver11, BKG^{+}22, Ber22, Bra02, Bra09, Bra12, Cam16, Dah02, DT06, Gre04, Hua11, Kaw07, Kru21, LL23, LP05, Ma17, MT99, Nic02, NR24, Ots11, Zam02, ÖD11]. Likelihood-based [JK15, LLL01, ÖD11]. likelihoods [FS04]. limit [CG05, DSW18]. limit-hitting [CG05]. Limited [MS98, Sil11, Gre04, HLL20, Mag10, WH07]. Limiting [Sen03, Die01]. limits [CG05]. Linear [DFGK98, AS11, AH21a, AYZ14, AS22, AHHK06, BKL08, BY13, BR04, Bra02, Bra09, BH07a, CM07, CNSY22, DD02, DM08, DP06, Esc18, FJV12, GO00, GOS06, GS13, HH00, JW12, JS04, Kri17, KNR11, LS05, LZZ21, TD14, Tho06, YP13, ZL15]. Linearity [Kil04, CS04, Dah02, HL07, San08]. linkages [CIM^{+}20]. linked [DMN20]. liquidity [Büt99]. liquidity-insuring [Büt99]. literature [Gia15]. lives [CCGV22]. LM [BY13, PUY08, Rod02, YT08, YV11]. LM-type [YV11]. Local [BK13b, MV07, CNS22, DHM23, GT99, HM09, MP08]. location [PS11b]. lockdown [CCGV22]. lockdowns [Bil22]. log [AN08, LY20]. log-concave [LY20]. logistic [CJP12, LZZ21]. logit [ADXBA21, GL07]. Long [Ros07, CJKL17, CS07b, DM04, PST08]. long-memory [DM04]. long-run [CJKL17]. look [DHM23]. Looking [GL09, Fan06]. Lorenz [DF99]. Löthgren [ÖD11]. Low [FVFW21]. Low-rank [FVFW21]. LSTUR [GL21]. Lyhagen [ÖD11].

M [ÖD11, Wil11]. Machine [IRS20b, KLS21, BHL22, Cha20, CCD^{+}18, CNS22, FHL^{+}22, Kna22, LZZ21, NSS22, SC21]. Macroeconometrics [Ano21c, Gör12]. macroeconomic [BMO04, HLN01, UT19]. magnitude [PS09]. making [HK21]. Malinvaud [Phi15]. many [Ana13, CD22, Kap08, LS21]. marginal [FS04]. market [BH16, HP10, MN11]. Markov [Ard09, BLV08, BPR10, Cav03, CK98, EKS08, FS04, JK10, Kan14, Kau00, Psa01, TK04]. Markov-switching [Ard09, CK98]. Markowitz [BLLW11]. Matching [KT23, Smi17, Frö07]. matrices [Aba23, DUM24, FLL16, MP05, ZZMZ21]. matrix [All07, Mag07, PA23]. matters [All07]. Maximization [FPR15, KR03]. Maximum [CLS14, Kam98, LST01, Zam02, BKG^{+}22, GY22, Gre04, Hua11, Kaw07, Kru21, LL23, LP05, MT99, ST21, WS05]. MCMC [Rag05]. Mean [SS06, AT18, Dan05, Dem02, Eve13, KjL03, Kru21, LR05, Mag02, Pit04, Sen09, Yan02, YV11]. mean-independence [KjL03]. mean-variance [Sen09]. means [HLRZ22]. measure [BK19, DG19]. Measurement [JM05, ABS03, AH21b, DHB12, HSW15, JK15]. measures [BCMM23, DSW18, Sch12b, Ver11]. Measuring [Kau00]. mediation [FHL^{+}22]. memory [DM04, PST08]. Method [HKLZ07, vdS98, BF19, GY22, HK01, MP05, Ots11, PS07]. methodology [CH02]. Methods [GM98, Sil11, AD13, BMO04, DHB12, Gal17, GS18, Kau00, KN08, Pro06, Rag05, Ulr08, Koo10]. Micro [Sil11]. Micro-Econometrics [Sil11]. microeconometrics [Rob10]. Migration [Gar20]. minimal [Aba23]. minimax [HIN22]. minimisation [CL21]. Minimum [May07, KS10, Win19]. mining [CE99, GT99, Han99a, Han99b, HK99, HP99a]. Misclassification [CS22]. Misclassification-robust [CS22]. mismeasured [CLTZ04]. missing [WL13]. missingness [BKOV20]. Misspecification [Kal12, PT11, Pit04, WZQ07]. misspecified [BMRSEA21]. mixed [BR07, Bri11, DUM24, GH21, MT99]. mixed-frequency [GH21]. mixed-Poisson [MT99]. mixture [ADXBA21, Ada16, AT04, FS04, Geo07, Kaw07, Sil03]. mixture-distribution [Geo07]. mixtures [CK16]. mobility [Ulr08]. mode [Kri17]. Model [HS17, LK16, RV02, Smi18, ZLS23, ZZMZ21, Abr99, AH04, AS19, Ard09, BCCW08, BT02, BSJ02, BKL08, BFK11, BR07, BPR10, BM06, BFHM08, BCMM23, BC07, Bri11, CS07a, CSW22, CR14, sCK10, Cho06, CDR09, CM24, DT06, Dem02, DG11, Gar20, Geo07, GM00, GT99, GFCK09, HZ12, HKLZ07, HK21, HH10, Isk10, JLLN08, JL01, JW12, JPX16, KK03, Kau00, KK11, KN08, KP04, KNR11, LP09, LO13, LPR17, LZZ21, LR05, MP11, NT09, NR24, PR14, PV01, Pét00, PS01, PS09, PS07, QP13, QWfL16, Sen03, Shi08, ST06, TfL14, TK04, VDP03, Wil09, YZLC21, Zam02, Zha13]. Model-selection [HS17]. modeling [CE99, SHD99]. Modelling [BLV08, CLTZ04, CH02, DV07, Rün04, Smi03, HJ05, Osb12]. Models [BLV08, BL98a, BL98b, CDF98, CK98, Kam98, KP98, Lar98, Mon98, PS16, PS11a, Sen98, XP98, Aba23, ADXBA21, AS11, AH21a, Ada16, AN08, AYZ14, AT04, AT11, AS22, AI22, AM07, AHHK06, BHKY17, BH16, BR04, Ber22, BK13b, Ble15, BW02, BM17, Bra02, Bra09, Bra12, BRS16, BC10, BW10, CDH^{+}21, CLTZ04, CL22, CPP01, Cha20, Cha17, CGLL15, Che19, CH21, CM18, sCK10, Cho23, CHS09, CS22, CH01, CK16, CK12, CNSY22, DSW18, DV07, DHS21, DHK23, DRB07, EKS08, Esc18, Eve13, Fan06, FVFW21, FP12, FF10, FJV12, FJ19, FK21, FS04, FHP21, GM04, GMMS19, GO00, GOS06, Gre04, GL07, Hag14, HJ99, HK00, HS04, HM13, HM09, HMY11, HdP21, HW07, HK01, HS17, HSW15, HS18, HWY08, HJ05, JK10, JJS14, Joc14]. models [JM17, JS04, Juh05, JPX12, Kan14, KS06, Kaw07, KL09, Khe15, Kim20, KP05, KMVW02, KS10, KP01a, KP04, KSD99, Kru21, LS05, fLLL10, fLY20, LG21, LCG11, LL23, LY20, LMW20, LM22, LV22, LMNS21, Mad10, Mag10, MP23, McK01, MCM23, MY00, MS05b, MZ18, MT99, NS12, NSS22, Nie04b, Nie08, Oom99, PY11, PT11, PS17, Rag05, RM99, RV02, RPS14, San08, Sil03, Smi08a, SHR14, TD14, Tho06, Tse02, Tug21, Ver11, WH07, WL13, WC20, YP13, Yan02, YZ23, YP01, YLW23, ZfL04, ZSP12, ZL15, Ano21b]. moderate [GSW19]. modified [NF11]. Moment [EKS08, KP05, AS22, Ber22, Bra09, Bra12, DHS21, HKLZ07, HS17, Ino06, Mof01, PT11]. moment-based [AS22]. Moments [DM07, Dem02, KK03, Sil11, vdS98, CK12, Ots11]. monetary [Ros21]. money [Nie08, Pét00]. monitoring [AHHK06]. monopolist [Koe20]. monotonic [GM00]. monotonicity [HM09]. Monte [KLS21, Par02, SHD99]. MSE [HHK04]. Multi [KRH^{+}09]. Multi-tail [KRH^{+}09]. multilateral [ALP^{+}21]. Multilayer [CIM^{+}20]. Multinomial [Bre02, GL07]. Multiple [Cha17, BP03, BCCW08, TW19, YP13]. multiple-break [BCCW08]. multiplicative [CSW22, YZ23]. multiplicity [OP23]. multiplier [BRS16, RR14]. Multivariate [AM09, PR14, AT04, DV07, DM04, Geo07, Haf09, JLLN08, LP09, MS05b, MT99, Nie04b, VDP03]. My [BM17]. Myoung [Sil11]. Myoung-jae [Sil11].

N [Gör12]. narrow [NF11]. narrow-band [NF11]. NASDAQ [QP13]. natural [BP21]. Near [KY21, Gos04, Mag07, PV01]. near-integrated [Gos04, PV01]. nearest [LG08, OLL06]. nearest-neighbour [OLL06]. nearly [CJKL17]. neglected [BK03]. neighbour [LG08, OLL06]. Nelson [CDR09]. Nelson-Siegel [CDR09]. nested [KMT03]. network [BK03, CIM^{+}20, KN20, fLLL10, LX17, LPR17, ZZMZ21]. Networks [AdP17, ABHK17, MTV17, Ros17]. Neumann [HP04]. neural [BK03, KN20]. neutral [BH07b]. never [DHM23]. news [AM09]. NIG [JL01, Wil09]. NLLS [BK13b]. Non [AM07, Ble15, Bri11, CJP12, Frö06, GK18, GM00, HZ12, HMY11, Kas15, KWTZ18, Kri12, LS05, LCG11, LNN09, Mar13, PS11b, Van10, Xu08, AH21a, BT02, BKL08, BR04, BZ18, Bra09, BH07a, CS07a, CM07, CK12, DSW18, DT06, FP12, FF10, GW12, GS13, HP04, HM09, Joc13, KS10, LHK15, LG08, May07, NS12, OLL06, PV01, RG06, SHR14, Spe09, Tak08, Xu15]. non-affine [FP12]. non-differentiable [LHK15]. Non-linear [LS05, AH21a, BR04, Bra09, BH07a, CM07, GS13]. Non-monotonic [GM00]. non-normal [DT06]. non-orthogonal [BT02]. Non-Parametric [Mar13, Bri11, Frö06, GK18, HMY11, Kas15, KWTZ18, Kri12, LCG11, LNN09, PS11b, Van10, CK12, DSW18, FF10, GW12, HZ12, Joc13, LG08, OLL06, Spe09, Tak08]. non-separable [HM09, KS10]. Non-standard [Ble15, NS12]. Non-stationary [CJP12, HZ12, Xu08, BKL08, BZ18, HP04, May07, RG06, SHR14, Xu15]. Non-trading [AM07]. non-zero [CS07a, PV01]. nonfundamental [LV22]. Nonlinear [AB16, CPP01, Osb12, BMRSEA21, Hag14, Khe15, Ma17, NSS22, RV02, Woo23]. nonlinearity [BK03, KP01b]. Nonparametric [BK22, CJKL17, CLW15, Du16, DHK23, Lec99, SRPR22, ALK23, Cam16, CSW22, FHP21, JPX16, Kuo08, Sca16, WW19, ZLS23]. nonparametrically [CLS14]. nonpharmaceutical [Cho20]. nonseparable [FVFW21, Kim20]. nonsmooth [HIN22]. nonstationary [CGLL15, CSW22, LV22, WC20]. norm [CM18]. normal [Dan05, DT06, Mag02]. Normality [DFGK98]. Norwegian [Akr04, EJN02]. Notation [AM02]. note [JM17, Sil03, Spe09]. Novel [HKR15]. nuisance [AC02, Bre02, ZW24]. null [LN00, LKSN03]. number [Kas15, LP09, ST06]. Numerical [Kaw07].

observation [HL05b, HH10, LR06]. observations [Nie08]. observed [Abr99, Hua11]. obtaining [HB20]. occur [GM04]. OECD [AM01]. Oil [Akr04, CIM^{+}20]. older [HP10]. OLS [PP04]. Omicron [Han22]. one [Ada16, GS18]. one-sided [GS18]. Online [CDH^{+}21]. only [AC02]. Optimal [CCF20, CLW20, HIN22, JW19, LV06, BT02, Büt99, Gal17, KY15, MPP14, WZO06]. optimality [KP01a]. optimization [DHS21]. option [BH07b, FP12]. options [FP12]. Oracle [GS18]. Ord [MP11]. Order [XP98, AD15, DMN20, FT00, HHK04, Ma17, TfL14]. ordered [KL09, LK05]. orderings [DF99]. ordinal [KZ23]. Orthogonal [Eve13, BT02, NSS22]. orthonormal [Sun13]. other [SC21]. our [HP99b]. outbreak [CCGV22, Cho20]. outcome [DT06, HdP21, MP23]. outcomes [CLTZ04, CR14, MCM23, Xu21]. outliers [Geo07, Nie04a]. outsourcing [BKOV20]. overdispersion [GL07]. overidentifying [CD22]. overlap [HLL20]. overview [FLL16]. Oxford [Mar12].

P [PY11]. Pairwise [Joc13, Mao16]. Pairwise-comparison [Joc13]. pandemic [Kor22]. Panel [Hua08, DHT09, Tug21, Wil11, YZLC21, Abr13, AH21a, AYZ14, AS19, AB16, AI22, BiS13, BKL08, BFK11, BY13, BMO04, BG19, BW02, BC07, BRS16, BW10, Can11, Cha17, Che19, CH21, CNSY22, DD12, DRB07, Eve13, FVFW21, FP12, FJ19, Had00, HL05a, HKR15, HLN01, HMY11, HdP21, HS18, JLLN08, Joc14, JM17, JW19, Kap08, Kim14, Kru21, fLY20, LCG11, Mad10, MPP14, MP11, OY20, ÖD11, PS03, QWfL16, RPS14, SR09, SRPR22, Tho06, WC20, Woo23, ZfL04, ZSP12, vBRV24]. panels [iSDLB05, CPT11, CSY23, HLRZ22, LLL01, McK01, MP08, SS06, Wes18, WK22, ZSS20]. parameter [Dem02, GH21, HJ99, KK11, LX12, LP05, NR24, PS09, RWW23]. parameters [AC02, Bre02, CCD^{+}18, CNS22, Kan14, NR24, SS24, TD14]. Parametric [Mar13, BK13b, Bri11, Cho23, CK12, DSW18, DG11, Emv20b, FF10, Frö06, GK18, GW12, HZ12, HMY11, HS18, Joc13, Kas15, KWTZ18, KS10, Kri12, LG08, LCG11, LNN09, OLL06, PS11b, RPS14, SS24, SRPR22, SHR14, Spe09, Tak08, Ulr08, Van10, YLW23, ZSP12, Zha13, Emv20a]. parsimonious [GH21]. Partial [AH21a, Kim20, Kom13]. Partially [AS22, WS05, AS11, AYZ14, FJV12, HP04, LZZ21, ZL15]. participation [KT18]. parts [FPR15]. patterns [VST03]. PcGets [Bår01]. Peer [LPR17, DDF09, LX17, Ros17]. penalised [CL21]. Penalized [NR24, UT19]. percentile [MP05]. Perez [GT99, Han99a, HK99]. Performance [CK98, DHT09, Tak08]. Periodic [Pro98, DO04]. periods [HS18]. Permanent [CL22]. Permanent-Transitory [CL22]. Perron [LN99]. persistence [LKSN03]. persistent [BK13a, LS05, San08]. Perspectives [PS11a]. phase [Rün04]. Phillips [LN99]. Phillips-Perron [LN99]. piecewise [KNR11]. plausibly [vKR18]. Ploberger [NS12]. Point [MPP14, AHHK06, BCCW08, BHKY17, CO12, Cho03, Yan12]. Point-optimal [MPP14]. pointwise [CHS09]. Poisson [MT99]. policy [Kor22, MCM23, Ros21]. polynomial [Cho06, WY23]. pooled [MP08]. Pooling [HC04]. population [HLL20, Xu21]. populations [YLW23]. Portfolio [BLV08, LPW14]. posedness [Sca16]. possibly [LHK15]. Posterior [BC14]. posteriori [GY22]. posteriors [IMY24]. Potential [Xu21, MP23]. power [KK05, MP08, Par02]. PPP [JLLN08]. practice [BNHLS09, FF10]. pre [WZO06]. pre-test [WZO06]. precision [FLL16, PA23]. Predictability [KR13]. predicted [Spe09]. predicting [KN20]. Prediction [Sør00, GO00]. Prediction-based [Sør00]. predictive [LW17, LK16]. predictors [AT11]. preferences [ABS03, GvV20]. Prescott [McE08]. presence [Ana13, Dav12, Gre04, Ioa05, JMN00, LG08, LX12, NV02, Nie04a, DHT09, YG02]. present [AC02, GvV20, GS13]. pressure [LX17]. prevalence [Sto22]. price [CG05, KR13, KjL03, Kon03, Van10]. prices [Akr04, CL22]. pricing [FP12]. principal [Cho23]. priors [IMY24]. Prize [Abb23a, Ano11a]. Prizes [Abb22]. Probabilistic [BCK20]. probabilities [LL23]. probability [CO12]. probit [Bre02, GFCK09, KL09]. problem [BW10, Sch12b]. problems [FF10, YG02]. procedure [Kap08, ST06]. procedures [KK11]. process [DLS09, EN09, LST01]. Processes [LC08, MS98, BK13a, DO04, CM07, DD02, DM04, Emv20b, FT00, Gos04, Haf09, Han05, HP04, Liu09, May07, Emv20a]. product [SHD99]. production [DSW18, VST03]. productivity [BKOV20, Emv20a, Emv20b]. products [Agu23, DHK23]. professional [BM22]. programme [Kna22]. Progress [EJN02]. projected [Cho23]. Projection [BW02, Abr13]. Propensity [Frö07]. properties [BKL08, LL23]. proportional [Bri11, Gør06]. proposal [AM02]. prospect [BLLW11]. pseudo [McK01]. pseudo-panels [McK01]. PULSE [JP22]. quadrature [BKG^{+}22]. qualitative [Tak08]. quality [KjL03]. Quantifying [Cho20]. Quantile [AT11, BK19, HC19, AD13, AGMR23, AB16, BG19, BTV24, Can11, CH21, CG23, FHP21, GK18, JJS14, Kal12, KM04, LMP14, Zha13]. Quantile-based [HC19]. quarterly [VST03, Rod02]. Quasi [Hua11, Cam16, Kru21, NR24]. quasi-likelihood [Cam16, NR24]. Quasi-maximum [Hua11, Kru21]. quotes [BNHLS09].

R [ÖD11]. R. [Sch12a]. radial [BK03]. random [AH21a, BT02, BKL08, BTV24, BM17, CR14, Dah02, DHK23, GFCK09, KMVW02, QP13, Zam02]. random-coefficients [CR14]. randomized [CLW20, KT18]. randomly [WL13]. range [Cav01]. ranges [GK18]. Rank [Abr99, AS11, Pal22, DLS09, FVFW21, GS13, HP04, LST01, RM99, ST21]. Rank-invariance [Pal22]. Rate [CDF98, DSBG22, GS18, PY11, YP01, iK07]. rates [Akr04, Ble15, HIN22, JLLN08]. ratio [Cam16, Die01, Ma17, MP08, YLW23]. rational [BMRSEA21, BFHM08, EN12, JS04]. ratios [FHP10]. Realized [BNHLS09, HV16, Med03, Ver11]. Reconsideration [BG19]. reconsidered [Han99a, Han99b, HK99, HP99a]. record [AM01]. recovering [BH16]. recurrent [LG21]. reduced [Kor22]. Reduction [CDF98, KMVW02]. redux [Koe20]. refinement [Ma17]. refinements [Cam16]. Reform [MW08]. regime [Cav03, Kan14, KS06, LL23]. regime-switching [Cav03, Kan14, LL23]. regimes [ST06]. Regression [KP98, AD13, AYZ14, AGMR23, AT11, BSJ02, BKL08, BG19, BTV24, BB11, CJKL17, CCF20, Can11, CJP12, CSW22, DGSS22, DMN20, EKS08, Esc18, FJV12, Frö06, GL21, GO00, GOS06, GL07, HB20, HJJX21, JJS14, KM04, KP05, KR23, Kri17, LMP14, LG08, LNN09, LR05, MT99, PS11b, Pro06, Sca16, SRPR22, Tsy23, UT19, WZO06, YZLC21, YZ13, Zha13, ZL15]. regression-based [BB11]. Regressions [DFGK98, Myn11, AB16, BY13, CH21, CS04, KW21, LK16, Mao16, RG06, YP13, YC22]. regressor [LK05, RVH05]. regressor-dependent [LK05]. regressors [Ana13, BiS13, BKL08, CJKL17, CPP01, CLS14, JPX16, KL09, Kru21, MZ18, Myn11, RM99]. Regularised [NSS22, CD22]. regularized [CNS22]. Relation [Sen98]. relations [Mad05]. Relative [Han22, Par02]. reliable [ABDW15]. Repeated [Lin05, KWTZ18]. Reply [HP99b]. Representation [Kuo08, Aba23, AGMR23, CM24, Han05, Med03]. representative [Pét00]. representers [CNS22]. resampling [AD13]. Residual [Ioa05, Tse02]. Residual-based [Ioa05, Tse02]. Residuals [PR16, BKL08, EN09, Kal12]. Residuals-based [PR16]. Response [AH04, BK13a, Ble15, CS07b, Cha17, FJ19, Man13]. responses [Gos04, LK05]. Restricted [JS04, WZQ07]. restriction [Ada16, AH21c, Ber22, HS17, Zha13]. restrictions [BMRSEA21, Bra02, Bra12, CD22, FHP21, GY22, LMW20, Ma17, Rea22, Ros17]. results [CK16, DG19, DM04, GOS06, WZO06]. retirement [Isk10]. return [GL21, QP13]. returns [KRH^{+}09]. reunified [BP21]. Revealing [IMY24]. Review [Bår01, Gör12, Koo10, Mar13, Mar12, Oom99, Osb12, Ros16, Sil11, Tay13, Wil11, CK16]. revisited [Pro06]. rich [GMMS19]. Riesz [CNS22]. right [BTV24]. Risk [FG03, BH07b, CL21, CSY23, HC19, JM05, Wil09]. risk-neutral [BH07b]. risks [CCG07]. Robust [AT18, BC07, CLLT14, HJ05, LHK15, AI22, BM06, CCF20, sCK10, CS22, Esc18, HB20, LO13, Mag10, Sun13, TD14]. robustness [JP22, SS24]. role [BF19, HLT08]. Root [KY21, Lar98, Tay13, XP98, BZ18, iSS06, Cav01, Cav03, CM07, DD12, Die01, HLT08, Ioa05, Juh05, JW19, KS06, KP05, LR06, Mad10, Mag07, MPP14, Psa01, DHT09]. roots [HL05b, LV22, MP08, Rod02]. Roy [Gar20]. Roy-model [Gar20]. Royal [Abb23b, Abb24, Ano19, Ano21c, Ano23, FS09, HS98, PS16, PS11a, Smi08b, Smi15, Smi16, Smi17, Smi18]. Rubin [DM14]. run [CJKL17].

S [HK99]. S&ARCH [JL01]. S&P [QP13]. S. [Sch12a]. Sample [DFGK98, CG23, CSY23, GL21, HK21, HL00, KK05, DHT09, Smi03]. sampled [LPR17]. Sampler [BL98b]. samplers [ZfL04]. samples [KMT03]. sampling [Cha11, FS04, Kiv13, Sil03, SY22]. Sargan [Abb22, Abb23a, Abb24, Ano11a]. SARS [BCMM23, DSBG22, Han22]. SARS-CoV-2 [BCMM23, DSBG22, Han22]. satisficing [HK21]. saved [CCGV22]. saving [YG02]. scale [BM22]. schemes [Kiv13]. schooling [CLW15]. score [CLS14, Frö07, FZ14]. search [Han99a, HP99a, KP01a]. Searching [Qu07]. Seasonal [HLT08, AT18, DO04, VST03, FT00, Rod02]. seasonality [AT18]. Second [Ma17]. Second-order [Ma17]. Section [Wil11, CPT11, HKR15, KW21, Mad05, PUY08, PS03]. sectional [Hua08, Kap08, Mao16, SR09, DHT09, SRPR22]. segregation [ABDW15]. select [MW23]. Selection [DRB07, Smi18, BM06, CL21, sCK10, CG23, DT06, GS13, HS17, Ino06, NR24, PS07, RV02, Smi03, UT19, ZLS23]. selective [BKOV20, KT18]. selectivity [Sto22]. Semi [Emv20b, KS10, Kri17, YLW23, Zha13, BK13b, Cho23, DG11, HS18, KT23, RPS14, SS24, SHR14, Ulr08, ZSP12, Emv20a]. semi-bandits [KT23]. semi-continuous [YLW23]. Semi-linear [Kri17]. Semi-parametric [Emv20b, KS10, YLW23, Zha13, BK13b, Cho23, DG11, HS18, RPS14, SS24, SHR14, Ulr08, ZSP12, Emv20a]. semilogarithmic [vGS02]. seminonparametric [iK07]. Semiparametric [AT04, BH07a, CCG07, FG03, GH06, Gør06, Shi08, WC20, AN08, Bra12, CS22, DD02, HJJX21, NSS22]. sensitivity [MV07, WZQ07]. separability [DSW18]. separable [HM09, JPX12, KS10]. Separating [AS19]. sequences [PP02, SHD99]. sequential [LW17, ST06]. serially [BKL08, MPP14, PP04]. Series [Tay13, BMO04, Bha01, CL22, CGLL15, DW19, VST03, FT00, GH06, GL09, Hag14, HL07, HS04, KW21, KP01b, LNN09, Mad10, MNT03, MJ19, MS05b, NV02, Nie04b, RWW23, RVH05, SHR14, Sun13, Osb12]. Set [HM13, Ble15]. SETAR [KS06]. sets [DM14, Kap08, Kiv13, KY15]. settings [GW12]. sex [CLW15]. shapes [KR13]. shift [YV11, Yan12]. shifts [AT18, PS11b, Psa01, QP13, Rün04]. shocks [BH16]. Short [ACMG^{+}11, HS18, PY11, PST08, SR09, YP01]. Short-term [ACMG^{+}11, PY11, YP01]. sided [GS18]. Siegel [CDR09]. sign [Rea22]. Signal [HK00]. significance [GO00]. similarity [iK07]. Simple [BB11, Woo23, BG19, Can11, CH21, Esc18, GL07, HdP21, TW19]. Simpler [HH18]. Simulated [Kam98, MT99, BKG^{+}22, CK12, DT06]. simulating [Nic02]. Simulation [DFGK98, GM98, GOS06, MS98, ZfL04, Kau00]. Simulation-based [DFGK98, GOS06, MS98]. Simultaneous [LK05, Kiv13, PS09]. sine [AH21b]. Single [LV22, CLW15, CS22, Gør06, HSW15]. single-index [CS22, Gør06, HSW15]. single-sex [CLW15]. Size [All07, DDF09]. skewness [BU09, GL09, Wil09]. slope [BRS16, Yan12]. small [DD12, Kon03, QWfL16]. small-country [Kon03]. smooth [CS04, HC19, LMP14, San08]. smoothed [CH21]. Smoothness [SZW10]. social [Bil22, fLLL10, LX17, LPR17, Man13, Ros17, TfL14]. social-influence-dependent [LX17]. Society [Abb23b, Abb24, Ano19, Ano21c, Ano23, FS09, HS98, PS16, PS11a, Smi08b, Smi15, Smi16, Smi17, Smi18]. solving [FF10]. Some [BMO04, GOS06, HJ99, CK16, Das05, Gal17]. sooner [CCGV22]. space [Kan14, KSD99, NR24, Oom99, Pro06]. spaced [McK01]. Sparse [ADXBA21, MTV17, PA23]. spatial [BY13, BB11, CSY23, KPR17, QWfL16, RR14, WL13]. Special [AdP17, Abb23b, Ano19, Ano21c, Ano23, CLLT14, FS09, PS16, PS11a, Smi15, Smi17, Smi18]. specific [Han99a, HP99a]. Specification [CGLL15, DT06, Khe15, fLLL10, BCCW08, Bha01, iSS06, Cho03, GT99, Han99a, HP99a, LW17, Tch15]. specifications [PST08]. spectral [YP13]. spells [DGSS22]. sphericity [BFK11]. spline [BH07b]. sport [BM22]. spread [BP21]. Spurious [Pro98]. squares [HW07, KP05, Kuo08, NF11, PR16, Pit04, PS17, WZQ07, Win19]. SsfPack [KSD99, Oom99]. Stability [vdS98]. Stable [LC08]. stage [AGMR23, CNSY22, DSW18, HW07, KMT03, KM04, WH07, Win19]. Standard [KMT03, AM02, Ble15, LN00, NS12]. Standardized [BY13]. state [Hal08, Kan14, KSD99, Oom99, Pro06]. state-space [Kan14]. Stationarity [Liu09, Had00, HL05a, Kil04, Kru21, LKSN03, SS06]. stationary [BKL08, BZ18, CJP12, CM07, CH01, HZ12, HP04, HS04, May07, PP02, RM99, RG06, SHR14, Xu08, Xu15, Yan02]. statistic [HH18, OY06, TfL14]. Statistical [Dav12, Koo10, KSD99, Oom99]. statistics [BLLW11, Cav01, Das05, GK18, PP02]. step [KK11, LV22, ZW24]. Stochastic [Bha01, Hag14, MFB03, Mon98, Smi08a, AM07, AM09, BLLW11, Dem02, GS13, JL01, Kaw07, KN08, KR03, LPW14, MY00, Nic02, QP13, Rag05, Rün04, TK04, Ver11]. strategies [KWTZ18]. strictly [Esc18]. strong [CPT11]. strongly [BK13a]. Structural [Ano21b, Ano21c, Isk10, BP03, iSS06, CCD^{+}18, DW19, DP06, HP10, HW07, Iga20, IRS20b, JMN00, KP01b, Kri12, LR06, LMNS21, MS05b, Rob10, TD14, Xu15, YP13, ZSS20, vBRV24, Gör12]. structure [Büt99, Che19, CSW22, CDR09, Dem02, GH21, Kim14, MTV17, ZZMZ21]. structured [DV07]. structures [DUM24, fLLL10]. student [EH08, Ard09]. Student- [Ard09]. study [BSJ02, Tak08]. subset [LS21]. substitution [vBRV24]. sufficiency [JM17]. sums [DD02]. surfaces [AM09]. surplus [Van10]. survey [Gal17, Sen09, ZW24, dCD23]. surveys [Lin05]. SVARs [Rea22]. Sweden [Cho20]. switching [Ard09, BPR10, Cav03, CK98, EKS08, FS04, Kan14, Kau00, LL23, TK04]. symmetric [LN00]. symmetry [BH07a]. synergies [IRS20b]. Synthetic [GY22, BF19]. system [ABS03, BW10, Qu07]. systems [BH07a].

tail [KRH^{+}09]. tailed [JL01, PS17, Sch12b]. tails [Dav12, FHP10]. tapered [PP02]. targeting [BJN03, PR14]. taxation [HP10]. technical [vBRV24]. technique [Nic02]. techniques [FS04]. technological [VST03]. technology [BKOV20]. temperature [GH06]. Temporal [MS05b, Pro06]. temporally [Haf09]. Ten [Abb22]. Teräsvirta [Osb12]. term [ACMG^{+}11, CDR09, DLS09, JS04, PY11, YP01]. Test [BLLW11, XP98, ALK23, BT02, BK03, Das05, Die01, HL00, KK05, LW17, MP11, OY06, PUY08, SS24, Sun13, WZO06, WW19, ZSS20]. Testing [BiS13, BFK11, Ber22, Bra02, CD22, Cav01, CS04, CG23, DLS09, DW19, EN12, FHP21, GW12, GSW19, Had00, HL05a, HH00, HL07, Hog17, KMVW02, KP01a, LPW14, LMNS21, Mao16, MJ19, NT09, NS12, OTP04, OP23, RWW23, Wri10, WX18, Xu15, ZSP12, BMRSEA21, Bra12, CGLL15, FP12, GH06, GS18, HL05b, Ioa05, JS04, JW19, LR05, PS03, WY23, YP13, iK07, Ano21b]. Tests [BMRSEA21, DFGK98, Lar98, LKSN03, Tay13, vdS98, AH21c, AT18, BP03, BY13, BHKY17, BB11, BZ18, BRS16, Cam16, iSS06, Cav03, CLW15, CM07, sCK10, Dah02, DSW18, DM14, DD12, Du16, EM02, GO00, GOS06, HKR15, HJ99, HLT08, Hsu17, HS17, Kal12, KS06, Khe15, Kil04, KY15, LLL01, LHK15, LN99, LN00, LR06, LV06, LST01, Ma17, Mad10, MV07, Mof01, MP08, MPP14, NS12, ÖD11, PR16, Pit04, RV02, RR14, Rod02, Ros07, San08, Sen03, Sen09, SS06, DHT09, Sto22, Tch15, YT08, YV11]. Theil [MS05a]. theorem [Han05, Kuo08]. theorems [DSW18]. Theoretical [PS11a]. Theory [BPR10, BT13, CO12, CSY23, Gia15, GL21, GS18]. there [LN00]. thick [FHP10]. Three [YZ23, HW07, KS06]. three-regime [KS06]. three-stage [HW07]. Three-way [YZ23]. Threshold [CK98, Ard09, CS07a, ST06, YC22, YZLC21, YZ13, Zha13]. thresholds [LK05]. Time [CDF98, Osb12, Tay13, CL22, CGLL15, Che19, CHS09, Dem02, DP06, EKS08, FT00, GvV20, GH21, GL09, HL05a, Hag14, HL07, HS04, HH10, HS18, KW21, KK11, Kim14, KjL03, KP01b, LCG11, LL23, LNN09, LY20, Mad10, MNT03, MJ19, MS05b, NV02, Nie04b, PY11, RWW23, San08, SHR14, Wil09, YP01]. time-inhomogeneous [CHS09]. time-series [Mad10]. time-varying [Che19, Dem02, GH21, HH10, KK11, LCG11, LL23, San08]. Timo [Osb12]. Tjøstheim [Osb12]. Tobit [CS07a]. too [DHM23]. topological [FK21]. trace [LST01]. track [AM01]. trades [BNHLS09]. Trading [ABHK17, AM07]. training [Lec99]. transformation [Abr99, AH04, AN08, AH21b, Eve13, FF10, HH10, Shi08, WC20, ZLS23]. Transition [Kam98, CJP12, CS04, HC19, JK10, LL23, San08]. Transitory [CL22]. transport [Gal17]. treated [MW18]. treatment [BK22, BHL22, CLW15, CCD^{+}18, DHM23, DT06, GS18, HLL20, Hsu17, KT18, LMW20, Man13, SC21, SY22, dCD23]. treatment-based [SY22]. treatment-outcome [DT06]. treatments [AH21c, TW19]. tree [DV07]. trees [DMN20]. trend [CH01, DLS09, DP06, GH06, HK01, Ioa05, JMN00]. trend-break [HK01]. trend-stationary [CH01]. trending [CPP01, GS13]. trends [GS13, Kim14, KNR11, Tho06, WY23, ZSP12]. trial [KT18]. trials [CLW20]. triangular [JPX16]. tribute [Phi15]. triple [OM22]. truncation [Abr99, HH10]. Turkey [Bil22]. Two [CNSY22, KM04, LMW20, WH07, Win19, dCD23, BSJ02, DSW18, HW07, KMT03, KK11, YLW23, ZW24, iK07]. two- [HW07]. Two-stage [CNSY22, KM04, WH07, Win19, DSW18, KMT03]. two-step [KK11, ZW24]. Two-way [LMW20, dCD23, BSJ02]. type [BRS16, LN00, Rod02, YV11]. types [Xu14].

U.S. [MN11]. UK [HLN01, IMY24]. unbalanced [Abr13, BSJ02, FP12]. uncertain [DLS09]. uncertainty [GT99]. unconditional [LR06]. unconfoundedness [Kna22]. unconventional [Ros21]. uncorrelated [NS12]. unemployment [DGSS22]. unequally [McK01]. Uniform [iK07]. Unifying [HJJX21]. Unit [DD12, KS06, KY21, Lar98, LR06, Mad10, Tay13, XP98, ALP^{+}21, BZ18, iSS06, Cav01, Cav03, CM07, Die01, HL05b, HLT08, Ioa05, Juh05, JW19, KP05, Mag07, MP08, MPP14, Psa01, Rod02, DHT09]. unit-root [CM07, Psa01]. United [Frö07]. Units [AH21b, Kap08]. univariate [BK13a, Dan05, Mag02]. Unknown [XP98, JK10, Wes18, ZL15, iK07]. unobservable [BiS13]. Unobserved [HK19, Agu23, BT02, HK00, YZ23]. unrestricted [Mag07]. use [BMO04, HWY08]. useful [CM07]. Using [BLV08, BL98b, CK16, HK21, MW23, Ulr08, Ber22, BH07b, CLTZ04, CD22, Cav01, CS07b, CH21, CNS22, CM18, CK12, Dah02, DDF09, DHS21, Emv20a, Emv20b, FS04, GK18, KW21, Kau00, KP04, KSD99, Mao16, MFB03, MS05b, Ros17, Smi03, Sun13, YP13]. utility [KR03]. utilization [DT06].

vaccine [BP21]. vaccines [BCMM23]. validation [OLL06]. Validity [HV16, Tch15]. Value [FG03, HC19, Wil09]. values [BP03, San08, WZO06]. VAR [HJ99, LST01, Nie08, Pét00, PS01, RM99, Tug21, Yan02, Hog17]. VAR-models [HJ99]. variable [BTV24, BKOV20, CR14, CK12, CNSY22, Gre04, JW12, Kim20, Mag10, QWfL16, Sca16, UT19, WH07, WL13]. variables [AGMR23, Ana13, BK19, DM08, Frö06, FHP21, HM13, JPX12, LG08, RG06, Spe09, WH07, vGS02, Sil11]. Variance [CDF98, KP98, Aba23, Dan05, DD02, HH18, Mag02, PR14, Pit04, Ros17, Sen09, Sun13, Ver11, WZO06, Wil09]. variances [Med03, Xu15]. variants [BCMM23, Han22]. Variates [CDF98]. variation [ALP^{+}21, DDF09]. VARMA [BL98a]. VARs [GH21]. varying [Che19, Dem02, GH21, HH10, KK11, LMP14, LCG11, LL23, RPS14, San08, Wil09, ZLS23]. Vector [Ano21b, BR04, CM24, DLS09, EN12, Hua08, JS04, LMNS21, MJ19, PA23]. Venezuela [CE99]. versus [LST01, Yan12]. via [Ano21b, AB16, CL22, HP04, KN08, LMNS21, MP05, PA23, UT19, WS05]. virus [Han22]. vocational [Lec99]. volatilities [BH16]. Volatility [Mon98, Xu08, Aba23, AM07, AM09, BH16, BZ18, DV07, Dem02, EKS08, HZ12, HV16, JL01, Kaw07, LS05, LP09, MY00, MFB03, NT09, Pal22, PST08, QP13, Rag05, Ver11, WX18]. Volume [Ano10, Ano11b, Ano12b, Ano13, Ano14, Ano15, Ano16, Ano17, Ano18, CD09]. Volumes [Tay13]. von-Neumann [HP04].

W [Osb12]. Wage [PS01, DRB07, Frö07, Gar20]. walk [BT02, KMVW02]. Watson [KK05]. wave [DSBG22]. way [BSJ02, LMW20, YZ23, dCD23]. Weak [CPT11, LX12, BW10, HHK04, Mag10, NF11, Tch15]. weakly [JPX12]. Weighted [JJS14, BHL22, Cho23]. weights [QWfL16]. welfare [DHB12]. where [HL05a, JK10, Mad10]. Wild [HB20, BZ18, MW18]. Windows [Bår01]. wise [MTV17]. within [BM06, fLY20]. without [Bre02, DHM23, Frö07, Gør06, HM09, Kru21, Tsy23]. Wooldridge [Wil11]. workers [HP10].

years [Abb22]. yield [BHKY17, KN20].

zero [CS07a, PV01, Rea22]. zeros [YG02].

References

Arellano:2016:NPD

[AB16] Manuel Arellano and Stéphane Bonhomme. Nonlinear panel data estimation via quantile regressions. The Econometrics Journal, 19(3):C61–C94, October 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Abadir:2023:EMR

[Aba23] Karim M. Abadir. Explicit minimal representation of variance matrices, and its implication for dynamic volatility models. The Econometrics Journal, 26(1):88–104, January 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

            //academic.oup.com/ectj/article/26/1/88/6675807.

Abbring:2022:TYD

[Abb22] Jaap H. Abbring. Ten years of Denis Sargan Econometrics Prizes: Editorial. The Econometrics Journal, 25(2):i–iii, May 2022. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL

http://academic.oup.com/ectj/article/25/2/i/6588014.

Abbring:2023:DSE

[Abb23a] Jaap H. Abbring. The 2022 Denis Sargan Econometrics Prize. The Econometrics Journal, 26(2):i, May 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/26/2/i/7160642.

Abbring:2023:RES

[Abb23b] Jaap H. Abbring. Royal Economic Society Annual Conference 2021 special issue on econometrics of dynamic discrete choice. The Econometrics Journal, 26(1):Ci–Cii, January 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/26/1/Ci/6982525.

Abbring:2024:RES

[Abb24] Jaap H. Abbring. Royal Economic Society Annual Conference 2022 Sargan Lecture. The Econometrics Journal, 27(1):Ci–Cii, January 2024. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
27/1/Ci/7502794.

Allen:2015:MRI

[ABDW15] Rebecca Allen, Simon Burgess, Russell Davidson, and Frank Windmeijer. More reliable inference for the dissimilarity index of segregation. The Econometrics Journal, 18(1):40–66, February 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Adamic:2017:TN

[ABHK17] Lada Adamic, Celso Brunetti, Jeffrey H. Harris, and Andrei Kirilenko. Trading networks. The Econometrics Journal, 20(3): S126–S149, October 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Abrevaya:1999:RET

[Abr99] Jason Abrevaya. Rank estimation of a transformation model with observed truncation. The Econometrics Journal, 2(2):292–305, December 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Abrevaya:2013:PAU

[Abr13] Jason Abrevaya. The projection approach for unbalanced panel data. The Econometrics Journal, 16(2):161–178, June 2013. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Aasness:2003:DPM

[ABS03] Jørgen Aasness, Erik Biørn, and Terje Skjerpen. Distribution of preferences and measurement errors in a disaggregated expenditure system. The Econometrics Journal, 6(2):374–400, December 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Altissimo:2002:BIN

[AC02] Filippo Altissimo and Valentina Corradi. Bounds for inference with nuisance parameters present only under the alternative. The Econometrics Journal, 5(2):494–519, December 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Angelini:2011:STF

[ACMG^{+}11] Elena Angelini, Gonzalo Camba-Mendez, Domenico Giannone, Lucrezia Reichlin, and Gerhard Rünstler. Short-term forecasts of euro area GDP growth. The Econometrics Journal, 14(1): C25–C44, February 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Aguirre:2013:NIM

[AD13] Víctor M. Aguirre and Manuel A. Domínguez. New inference methods for quantile regression based on resampling. The Econometrics Journal, 16(2):278–283, June 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Arvanitis:2015:CII

[AD15] Stelios Arvanitis and Antonis Demos. A class of indirect inference estimators: higher-order asymptotics and approximate bias correction. The Econometrics Journal, 18(2):200–241, June 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Adams:2016:FMM

[Ada16] Christopher P. Adams. Finite mixture models with one exclusion restriction. The Econometrics Journal, 19(2):150–165, June 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Abbring:2017:SIE

[AdP17] Jaap H. Abbring and Áureo de Paula. Special issue on econometrics of networks: Editorial. The Econometrics Journal, 20 (3):Si–Sii, October 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Aboutaleb:2021:SCE

[ADXBA21] Youssef M. Aboutaleb, Mazen Danaf, Yifei Xie, and Moshe E. Ben-Akiva. Sparse covariance estimation in logit mixture models. The Econometrics Journal, 24(3):377–398, September 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/24/3/
377/6178866.

Alejo:2023:FSR

[AGMR23] Javier Alejo, Antonio F. Galvao, and Gabriel Montes-Rojas. A first-stage representation for instrumental variables quantile regression. The Econometrics Journal, 26(3):350–377, September 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/26/3/
350/7100955.

Aguirregabiria:2023:DDD

[Agu23] Victor Aguirregabiria. Dynamic demand for differentiated products with fixed-effects unobserved heterogeneity. The Econometrics Journal, 26(1):C1–C25, January 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:/

/academic.oup.com/ectj/article/26/1/C1/6761678.

Abrevaya:2004:RET

[AH04] Jason Abrevaya and Jerry A. Hausman. Response error in a transformation model with an application to earnings-equation estimation. The Econometrics Journal, 7(2):366–388, December 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Abrevaya:2021:PEN

[AH21a] Jason Abrevaya and Yu-Chin Hsu. Partial effects in non-linear panel data models with correlated random effects. The Econometrics Journal, 24(3):519–535, September 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/24/3/519/6156618.

Aihounton:2021:UMI

[AH21b] Ghislain B. D. Aihounton and Arne Henningsen. Units of measurement and the inverse hyperbolic sine transformation. The Econometrics Journal, 24(2):334–351, May 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/24/2/334/5948096.

Andresen:2021:IBE

[AH21c] Martin E. Andresen and Martin Huber. Instrument-based estimation with binarised treatments: issues and tests for the exclusion restriction. The Econometrics Journal, 24(3):536– 558, September 2021. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/24/3/536/6126343.

Aue:2006:CPM

[AHHK06] Alexander Aue, Lajos Horváth, Marie Husková, and Piotr Kokoszka. Change-point monitoring in linear models. The Econometrics Journal, 9(3):373–403, November 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Arkhangelsky:2022:DRI

[AI22] Dmitry Arkhangelsky and Guido W. Imbens. Doubly robust identification for causal panel data models. The Econometrics Journal, 25(3):649–674, September 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:/

/academic.oup.com/ectj/article/25/3/649/6617637.

Akram:2004:OPE

[Akr04] Q. Farooq Akram. Oil prices and exchange rates: Norwegian evidence. The Econometrics Journal, 7(2):476–504, December 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Aradillas-Lopez:2023:NTC

[ALK23] Andrés Aradillas-López and Lidia Kosenkova. A nonparametric test for cooperation in discrete games. The Econometrics Journal, 26(2):257–278, May 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/26/2/257/6969418.

Allen:2007:SMC

[All07] Jason Allen. Size matters: covariance matrix estimation under the alternative. The Econometrics Journal, 10(3):637–644, November 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anderson:2021:UFA

[ALP^{+}21] Gordon Anderson, Oliver Linton, Maria Grazia Pittau, Yoon-Jae Whang, and Roberto Zelli. On unit free assessment of the extent of multilateral distributional variation. The Econometrics Journal, 24(3):502–518, September 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.

oup.com/ectj/article/24/3/502/6125966.

Artis:2001:FFT

[AM01] Michael Artis and Massimiliano Marcellino. Fiscal forecasting: The track record of the IMF, OECD and EC. The Econometrics Journal, 4(1):20–36, June 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Abadir:2002:NEP

[AM02] Karim Abadir and Jan Magnus. Notation in econometrics: a proposal for a standard. The Econometrics Journal, 5(1):76–90, June 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Asai:2007:NTD

[AM07] Manabu Asai and Michael McAleer. Non-trading day effects in asymmetric conditional and stochastic volatility models. The Econometrics Journal, 10(1):113–123, March 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Asai:2009:MSV

[AM09] Manabu Asai and Michael McAleer. Multivariate stochastic volatility, leverage and news impact surfaces. The Econometrics Journal, 12(2):292–309, July 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Ai:2008:SDE

[AN08] Chunrong Ai and Edward C. Norton. A semiparametric derivative estimator in log transformation models. The Econometrics Journal, 11(3):538–553, November 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anatolyev:2013:IVE

[Ana13] Stanislav Anatolyev. Instrumental variables estimation and inference in the presence of many exogenous regressors. The Econometrics Journal, 16(1):27–72, February 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anonymous:2005:C

[Ano05] Anonymous. Corrigendum. The Econometrics Journal, 8(1):112– 113, March 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anonymous:2009:E

[Ano09] Anonymous. Errata. The Econometrics Journal, 12(3):447, November 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anonymous:2010:IEJ

[Ano10] Anonymous. Index to The Econometrics Journal volume 13. The Econometrics Journal, 13(3):291–292, October 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anonymous:2011:DSE

[Ano11a] Anonymous. The Denis Sargan Econometrics Prize. The Econometrics Journal, 14(1):Ai, February 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anonymous:2011:IEJ

[Ano11b] Anonymous. Index to The Econometrics Journal volume 14. The Econometrics Journal, 14(3):499–500, October 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anonymous:2012:E

[Ano12a] Anonymous. Erratum. The Econometrics Journal, 15(3):535, October 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anonymous:2012:IEJ

[Ano12b] Anonymous. Index to The Econometrics Journal volume 15. The Econometrics Journal, 15(3):537–538, October 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anonymous:2013:IEJ

[Ano13] Anonymous. Index to The Econometrics Journal volume 16. The Econometrics Journal, 16(3):485–486, October 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anonymous:2014:IEJ

[Ano14] Anonymous. Index to The Econometrics Journal volume 17. The Econometrics Journal, 17(3):395–396, October 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anonymous:2015:IEJ

[Ano15] Anonymous. Index to The Econometrics Journal volume 18. The Econometrics Journal, 18(3):437, October 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anonymous:2016:IEJ

[Ano16] Anonymous. Index to The Econometrics Journal volume 19. The Econometrics Journal, 19(3):291–292, October 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anonymous:2017:IEJ

[Ano17] Anonymous. Index to The Econometrics Journal volume 20. The Econometrics Journal, 20(3):271–272, October 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anonymous:2018:IEJ

[Ano18] Anonymous. Index to The Econometrics Journal volume 21. The Econometrics Journal, 21(3):354, October 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Anonymous:2019:RES

[Ano19] Anonymous. Royal Economic Society Annual Conference 2017 Special Issue on Econometrics of Games. The Econometrics Journal, 22(1):–, January 2019. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/22/1/Ci/5317154.

Anonymous:2021:E

[Ano21a] Anonymous. Editorial. The Econometrics Journal, 24(2):Ci–Civ, May 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
24/2/Ci/6311309.

Anonymous:2021:ETI

[Ano21b] Anonymous. Erratum to: Testing Identification via Heteroskedasticity in Structural Vector Autoregressive Models. The Econometrics Journal, 24(1):198, January 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/24/1/198/5866004. See

[LMNS21].

Anonymous:2021:RES

[Ano21c] Anonymous. Royal Economic Society Annual Conference 2018 special issue on structural macroeconometrics. The Econometrics Journal, 24(1):Ci–Ciii, January 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.

oup.com/ectj/article/24/1/Ci/6161542.

Anonymous:2023:RES

[Ano23] Anonymous. Royal Economic Society Annual Conference 2022 Special Issue on The New Difference-in-Differences. The Econometrics Journal, 26(3):Ci–Cii, September 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/26/3/Ci/7278391.

Ardia:2009:BEM

[Ard09] David Ardia. Bayesian estimation of a Markov-switching threshold asymmetric GARCH model with Student-t innovations. The Econometrics Journal, 12(1):105–126, March 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Abrevaya:2011:REP

[AS11] Jason Abrevaya and Youngki Shin. Rank estimation of partially linear index models. The Econometrics Journal, 14(3):409–437, October 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Amsler:2019:SDI

[AS19] Christine Amsler and Peter Schmidt. Separating different individual effects in a panel data model. The Econometrics Journal, 22(2):173–187, May 2019. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/22/2/173/5487077.

Antoine:2022:PLM

[AS22] Bertille Antoine and Xiaolin Sun. Partially linear models with endogeneity: a conditional moment-based approach. The Econometrics Journal, 25(1):256–275, January 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:/

/academic.oup.com/ectj/article/25/1/256/6325167.

Alfo:2004:SMM

[AT04] Marco Alfò and Giovanni Trovato. Semiparametric mixture models for multivariate count data, with application. The Econometrics Journal, 7(2):426–454, December 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Ando:2011:QRM

[AT11] Tomohiro Ando and Ruey S. Tsay. Quantile regression models with factor-augmented predictors and information criterion. The Econometrics Journal, 14(1):1–24, February 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Astill:2018:RTD

[AT18] S. Astill and A. M. R. Taylor. Robust tests for deterministic seasonality and seasonal mean shifts. The Econometrics Journal, 21(3):277–297, October 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Ai:2014:EFE

[AYZ14] Chunrong Ai, Jinhong You, and Yong Zhou. Estimation of fixed effects panel data partially linear additive regression models. The Econometrics Journal, 17(1):83–106, February 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Baardsen:2001:RPW

[Bår01] Gunnar Bårdsen. Review of PcGets 1 for Windows. The Econometrics Journal, 4(2):311–318, December 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Born:2011:SRB

[BB11] Benjamin Born and Jörg Breitung. Simple regression-based tests for spatial dependence. The Econometrics Journal, 14(2):330– 342, July 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bramati:2007:REF

[BC07] Maria Caterina Bramati and Christophe Croux. Robust estimators for the fixed effects panel data model. The Econometrics Journal, 10(3):521–540, November 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Browning:2010:HDD

[BC10] Martin Browning and Jesus M. Carro. Heterogeneity in dynamic discrete choice models. The Econometrics Journal, 13(1):1–39, February 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Belloni:2014:PIC

[BC14] Alexandre Belloni and Victor Chernozhukov. Posterior inference in curved exponential families under increasing dimensions. The Econometrics Journal, 17(2):S75–S100, June 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bai:2008:GCB

[BCCW08] Jushan Bai, Haiqiang Chen, Terence Tai-Leung Chong, and Seraph Xin Wang. Generic consistency of the break-point estimators under specification errors in a multiple-break model. The Econometrics Journal, 11(2):287–307, July 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Banerjee:2020:PFB

[BCK20] Anurag Banerjee, Guillaume Chevillon, and Marie Kratz. Probabilistic forecasting of bubbles and flash crashes. The Econometrics Journal, 23(2):297–315, May 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.

oup.com/ectj/article/23/2/297/5736054.

Boldea:2023:DEM

[BCMM23] Otilia Boldea, Adriana Cornea-Madeira, and João Madeira. Disentangling the effect of measures, variants, and vaccines on SARS-CoV-2 infections in England: a dynamic intensity model. The Econometrics Journal, 26(3):444–466, September 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/26/3/
444/6998553.

Berger:2022:TCM

[Ber22] Yves G. Berger. Testing conditional moment restriction models using empirical likelihood. The Econometrics Journal, 25(2):384– 403, May 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
25/2/384/6400100.

Botosaru:2019:RCS

[BF19] Irene Botosaru and Bruno Ferman. On the role of covariates in the synthetic control method. The Econometrics Journal, 22 (2):117–130, May 2019. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/22/2/117/5303851.

Beyer:2008:FAM

[BFHM08] Andreas Beyer, Roger E. A. Farmer, Jérôme Henry, and Massimiliano Marcellino. Factor analysis in a model with rational expectations. The Econometrics Journal, 11(2):271–286, July 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Baltagi:2011:TSF

[BFK11] Badi H. Baltagi, Qu Feng, and Chihwa Kao. Testing for sphericity in a fixed effects panel data model. The Econometrics Journal, 14 (1):25–47, February 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Besstremyannaya:2019:RSA

[BG19] Galina Besstremyannaya and Sergei Golovan. Reconsideration of a simple approach to quantile regression for panel data. The Econometrics Journal, 22(3):292–308, September 2019. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/22/3/292/5549841.

Brown:2007:SEB

[BH07a] Bryan W. Brown and Douglas J. Hodgson. Semiparametric efficiency bounds in dynamic non-linear systems under elliptical symmetry. The Econometrics Journal, 10(1):35–48, March 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bu:2007:EOI

[BH07b] Ruijun Bu and Kaddour Hadri. Estimating option implied riskneutral densities using spline and hypergeometric functions. The Econometrics Journal, 10(2):216–244, July 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Barigozzi:2016:GDF

[BH16] Matteo Barigozzi and Marc Hallin. Generalized dynamic factor models and volatilities: recovering the market volatility shocks. The Econometrics Journal, 19(1):C33–C60, February 2016. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bhargava:2001:SSI

[Bha01] Alok Bhargava. Stochastic specification and the international GDP series. The Econometrics Journal, 4(2):273–286, December 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bardsley:2017:CPT

[BHKY17] Patrick Bardsley, Lajos Horváth, Piotr Kokoszka, and Gabriel Young. Change point tests in functional factor models with application to yield curves. The Econometrics Journal, 20(1):86–117, February 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bodory:2022:EWD

[BHL22] Hugo Bodory, Martin Huber, and Lukás Lafférs. Evaluating (weighted) dynamic treatment effects by double machine learning. The Econometrics Journal, 25(3):628–648, September 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/25/3/
628/6604379.

Bilgel:2022:ECL

[Bil22] Firat Bilgel. Effects of Covid-19 lockdowns on social distancing in Turkey. The Econometrics Journal, 25(3):781–805, September 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
25/3/781/6590814.

Bai:2013:TPC

[BiS13] Jushan Bai and Josep Lluís Carrion i Silvestre. Testing panel cointegration with unobservable dynamic common factors that are correlated with the regressors. The Econometrics Journal, 16 (2):222–249, June 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Baardsen:2003:EIT

[BJN03] Gunnar Bårdsen, Eilev S. Jansen, and Ragnar Nymoen. Econometric inflation targeting. The Econometrics Journal, 6(2):430– 461, December 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Blake:2003:RBF

[BK03] Andrew P. Blake and George Kapetanios. A radial basis function artificial neural network test for neglected nonlinearity. The Econometrics Journal, 6(2):357–373, December 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Baillie:2013:EII

[BK13a] Richard T. Baillie and George Kapetanios. Estimation and inference for impulse response functions from univariate strongly persistent processes. The Econometrics Journal, 16(3):373–399, October 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Blevins:2013:LNE

[BK13b] Jason R. Blevins and Shakeeb Khan. Local NLLS estimation of semi-parametric binary choice models. The Econometrics Journal, 16(2):135–160, June 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Barunik:2019:QCG

[BK19] Jozef Baruník and Tobias Kley. Quantile coherency: a general measure for dependence between cyclical economic variables. The Econometrics Journal, 22(2):131–152, May 2019. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/22/2/131/5303852.

Ban:2022:NBT

[BK22] Kyunghoon Ban and Désiré Kédagni. Nonparametric bounds on treatment effects with imperfect instruments. The Econometrics Journal, 25(2):477–493, May 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.

oup.com/ectj/article/25/2/477/6445996.

Bansal:2022:DQA

[BKG^{+}22] Prateek Bansal, Vahid Keshavarzzadeh, Angelo Guevara, Shanjun Li, and Ricardo A. Daziano. Designed quadrature to approximate integrals in maximum simulated likelihood estimation. The Econometrics Journal, 25(2):301–321, May 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/25/2/301/6325166.

Baltagi:2008:APE

[BKL08] Badi H. Baltagi, Chihwa Kao, and Long Liu. Asymptotic properties of estimators for the linear panel regression model with random individual effects and serially correlated errors: the case of stationary and non-stationary regressors and residuals. The Econometrics Journal, 11(3):554–572, November 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Breunig:2020:ITO

[BKOV20] Christoph Breunig, Michael Kummer, Joerg Ohnemus, and Steffen Viete. Information technology outsourcing and firm productivity: eliminating bias from selective missingness in the dependent variable. The Econometrics Journal, 23(1):88–114, January 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/23/1/
88/5572205.

Bartel:1998:EKI

[BL98a] Holger Bartel and Helmut Lütkepohl. Estimating the Kronecker indices of cointegrated echelon-form VARMA models. The Econometrics Journal, 1(1):76–99, June 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bauwens:1998:BIG

[BL98b] Luc Bauwens and Michel Lubrano. Bayesian inference on GARCH models using the Gibbs sampler. The Econometrics Journal, 1(1):23–46, June 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Blevins:2015:NSR

[Ble15] Jason R. Blevins. Non-standard rates of convergence of criterionfunction-based set estimators for binary response models. The Econometrics Journal, 18(2):172–199, June 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bai:2011:TSP

[BLLW11] Zhidong Bai, Hua Li, Huixia Liu, and Wing-Keung Wong. Test statistics for prospect and Markowitz stochastic dominances with applications. The Econometrics Journal, 14(2):278–303, July 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Banachewicz:2008:MPD

[BLV08] Konrad Banachewicz, André Lucas, and Aad Van Der Vaart. Modelling portfolio defaults using hidden Markov models with covariates. The Econometrics Journal, 11(1):155–171, March 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bednarski:2006:RMS

[BM06] Tadeusz Bednarski and Edyta Mocarska. On robust model selection within the Cox model. The Econometrics Journal, 9(2): 279–290, July 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Boucher:2017:MFF

[BM17] Vincent Boucher and Ismael Mourifié. My friend far, far away: a random field approach to exponential random graph models. The Econometrics Journal, 20(3):S14–S46, October 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Breidenbach:2022:LSS

[BM22] Philipp Breidenbach and Timo Mitze. Large-scale sport events and COVID-19 infection effects: evidence from the German professional football ‘experiment’. The Econometrics Journal, 25 (1):15–45, January 2022. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/25/1/15/6318366.

Banerjee:2004:SCU

[BMO04] Anindya Banerjee, Massimiliano Marcellino, and Chiara Osbat. Some cautions on the use of panel methods for integrated series of macroeconomic data. The Econometrics Journal, 7(2):322– 340, December 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bera:2021:TNR

[BMRSEA21] Anil Bera, Gabriel Montes-Rojas, Walter Sosa-Escudero, and Javier Alejo. Tests for nonlinear restrictions under misspecified alternatives with an application to testing rational expectation hypotheses. The Econometrics Journal, 24(1):41–57, January 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/24/1/
41/5835215.

Barndorff-Nielsen:2009:RKP

[BNHLS09] O. E. Barndorff-Nielsen, P. Reinhard Hansen, A. Lunde, and N. Shephard. Realized kernels in practice: trades and quotes. The Econometrics Journal, 12(3):C1–C32, November 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bai:2003:CVM

[BP03] Jushan Bai and Pierre Perron. Critical values for multiple structural change tests. The Econometrics Journal, 6(1):72–78, June 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bluhm:2021:SCB

[BP21] Richard Bluhm and Maxim Pinkovskiy. The spread of COVID-19 and the BCG vaccine: a natural experiment in reunified Germany. The Econometrics Journal, 24(3):353–376, September 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
24/3/353/6271317.

Bauwens:2010:TIM

[BPR10] Luc Bauwens, Arie Preminger, and Jeroen V. K. Rombouts. Theory and inference for a Markov switching GARCH model. The Econometrics Journal, 13(2):218–244, July 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bec:2004:VEC

[BR04] Frédérique Bec and Anders Rahbek. Vector equilibrium correction models with non-linear discontinuous adjustments. The Econometrics Journal, 7(2):628–651, December 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bauwens:2007:BIM

[BR07] L. Bauwens and J. V. K. Rombouts. Bayesian inference for the mixed conditional heteroskedasticity model. The Econometrics Journal, 10(2):408–425, July 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bravo:2002:TLR

[Bra02] Francesco Bravo. Testing linear restrictions in linear models with empirical likelihood. The Econometrics Journal, 5(1):104–130, June 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bravo:2009:BGE

[Bra09] Francesco Bravo. Blockwise generalized empirical likelihood inference for non-linear dynamic moment conditions models. The Econometrics Journal, 12(2):208–231, July 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bravo:2012:GEL

[Bra12] Francesco Bravo. Generalized empirical likelihood testing in semiparametric conditional moment restrictions models. The Econometrics Journal, 15(1):1–31, February 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Breslaw:2002:MPE

[Bre02] Jon A. Breslaw. Multinomial probit estimation without nuisance parameters. The Econometrics Journal, 5(2):417–434, December 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Brinch:2011:NPI

[Bri11] Christian N. Brinch. Non-parametric identification of the mixed proportional hazards model with interval-censored durations. The Econometrics Journal, 14(2):343–350, July 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Breitung:2016:LMT

[BRS16] Jörg Breitung, Christoph Roling, and Nazarii Salish. Lagrange multiplier type tests for slope homogeneity in panel data models. The Econometrics Journal, 19(2):166–202, June 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Baltagi:2002:CSA

[BSJ02] Badi H. Baltagi, Seuck H. Song, and Byoung C. Jung. A comparative study of alternative estimators for the unbalanced two-way error component regression model. The Econometrics Journal, 5(2):480–493, December 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bailey:2002:OTA

[BT02] Ralph W. Bailey and A. M. Robert Taylor. An optimal test against a random walk component in a non-orthogonal unobserved components model. The Econometrics Journal, 5(2):520– 532, December 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bontemps:2013:IET

[BT13] Christian Bontemps and Elie Tamer. Identification in econometrics, theory and applications: EDITORIAL. The Econometrics Journal, 16(1):Si–Sii, February 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Beyhum:2024:IVQ

[BTV24] Jad Beyhum, Lorenzo Tedesco, and Ingrid Van Keilegom. Instrumental variable quantile regression under random right censoring. The Econometrics Journal, 27(1):21–36, January 2024. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/27/1/21/7231783.

Bao:2009:SKE

[BU09] Yong Bao and Aman Ullah. On skewness and kurtosis of econometric estimators. The Econometrics Journal, 12(2):232–247, July 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Buttler:1999:OCS

[Büt99] Hans-Jürg Büttler. The optimal capital structure of a liquidityinsuring bank. The Econometrics Journal, 2(2):268–291, December 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bond:2002:PEA

[BW02] Stephen Bond and Frank Windmeijer. Projection estimators for autoregressive panel data models. The Econometrics Journal, 5 (2):457–479, December 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Bun:2010:WIP

[BW10] Maurice J. G. Bun and Frank Windmeijer. The weak instrument problem of the system GMM estimator in dynamic panel data models. The Econometrics Journal, 13(1):95–126, February 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Baltagi:2013:SLT

[BY13] Badi H. Baltagi and Zhenlin Yang. Standardized LM tests for spatial error dependence in linear or panel regressions. The Econometrics Journal, 16(1):103–134, February 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Boswijk:2018:AWB

[BZ18] H. Peter Boswijk and Yang Zu. Adaptive wild bootstrap tests for a unit root with non-stationary volatility. The Econometrics Journal, 21(2):87–113, June 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Camponovo:2016:ARN

[Cam16] Lorenzo Camponovo. Asymptotic refinements of nonparametric bootstrap for quasi-likelihood ratio tests for classes of extremum estimators. The Econometrics Journal, 19(1):33–54, February 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Canay:2011:SAQ

[Can11] Ivan A. Canay. A simple approach to quantile regression for panel data. The Econometrics Journal, 14(3):368–386, October 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Cavaliere:2001:TUR

[Cav01] Giuseppe Cavaliere. Testing the unit root hypothesis using generalized range statistics. The Econometrics Journal, 4(1):70–88, June 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Cavaliere:2003:AUR

[Cav03] Giuseppe Cavaliere. Asymptotics for unit root tests under Markov regime-switching. The Econometrics Journal, 6(1):193– 216, June 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Chernozhukov:2018:DDM

[CCD^{+}18] Victor Chernozhukov, Denis Chetverikov, Mert Demirer, Esther Duflo, Christian Hansen, Whitney Newey, and James Robins.

Double/debiased machine learning for treatment and structural parameters. The Econometrics Journal, 21(1):C1–C68, February 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Calonico:2020:OBC

[CCF20] Sebastian Calonico, Matias D. Cattaneo, and Max H. Farrell. Optimal bandwidth choice for robust bias-corrected inference in regression discontinuity designs. The Econometrics Journal, 23 (2):192–210, May 2020. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/23/2/192/5625071.

Canals-Cerda:2007:SCR

[CCG07] José Canals-Cerdá and Shiferaw Gurmu. Semiparametric competing risks analysis. The Econometrics Journal, 10(2):193–215, July 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Cerqueti:2022:SBL

[CCGV22] Roy Cerqueti, Raffaella Coppier, Alessandro Girardi, and Marco Ventura. The sooner the better: lives saved by the lockdown during the COVID-19 outbreak. The case of Italy. The Econometrics Journal, 25(1):46–70, January 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.

oup.com/ectj/article/25/1/46/6363683.

Coudin:2009:IEJ

[CD09] Elise Coudin and Jean-Marie Dufour. Index to The Econometrics Journal volume 12. The Econometrics Journal, 12(3):449–450, November 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Carrasco:2022:TOR

[CD22] Marine Carrasco and Mohamed Doukali. Testing overidentifying restrictions with many instruments and heteroscedasticity using regularised jackknife IV. The Econometrics Journal, 25(1):71–97, January 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
25/1/71/6318365.

Calzolari:1998:CVV

[CDF98] Giorgio Calzolari, Francesca Di Iorio, and Gabriele Fiorentini. Control variates for variance reduction in indirect inference: Interest rate models in continuous time. The Econometrics Journal, 1(1):100–112, June 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Cai:2021:OED

[CDH^{+}21] Michael Cai, Marco Del Negro, Edward Herbst, Ethan Matlin, Reca Sarfati, and Frank Schorfheide. Online estimation of DSGE models. The Econometrics Journal, 24(1):C33–C58, January 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/24/1/
C33/5909595.

Christensen:2009:AFG

[CDR09] Jens H. E. Christensen, Francis X. Diebold, and Glenn D. Rudebusch. An arbitrage-free generalized Nelson-Siegel term structure model. The Econometrics Journal, 12(3):C33–C64, November 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Campos:1999:CDM

[CE99] Julia Campos and Neil R. Ericsson. Constructive data mining: modeling consumers’ expenditure in Venezuela. The Econometrics Journal, 2(2):226–240, December 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Chung:2005:EEP

[CG05] Jeff Chung and Li Gan. Estimating the effect of price limits on limit-hitting days. The Econometrics Journal, 8(1):79–96, March 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Corradi:2023:TQS

[CG23] Valentina Corradi and Daniel Gutknecht. Testing for quantile sample selection. The Econometrics Journal, 26(2):147–173, May 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/26/2/
147/6840230.

Chen:2015:STN

[CGLL15] Jia Chen, Jiti Gao, Degui Li, and Zhengyan Lin. Specification testing in nonstationary time series models. The Econometrics Journal, 18(1):117–136, February 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Clements:2001:FDS

[CH01] Michael P. Clements and David. F. Hendry. Forecasting with difference-stationary and trend-stationary models. The Econometrics Journal, 4(1):1–19, June 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Clements:2002:MMF

[CH02] Michael P. Clements and David F. Hendry. Modelling methodology and forecast failure. The Econometrics Journal, 5(2):319– 344, December 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Chen:2021:SEQ

[CH21] Liang Chen and Yulong Huo. A simple estimator for quantile panel data models using smoothed quantile regressions. The Econometrics Journal, 24(2):247–263, May 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/24/2/247/5881299.

Chambers:2011:CSF

[Cha11] Marcus J. Chambers. Cointegration and sampling frequency. The Econometrics Journal, 14(2):156–185, July 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Charbonneau:2017:MFE

[Cha17] Karyne B. Charbonneau. Multiple fixed effects in binary response panel data models. The Econometrics Journal, 20(3):S1–S13, October 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Chang:2020:DDM

[Cha20] Neng-Chieh Chang. Double/debiased machine learning for difference-in-differences models. The Econometrics Journal, 23 (2):177–191, May 2020. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/23/2/177/5722119.

Chen:2019:ELG

[Che19] Jia Chen. Estimating latent group structure in time-varying coefficient panel data models. The Econometrics Journal, 22(3): 223–240, September 2019. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/22/3/223/5497920.

Chong:2003:GCB

[Cho03] Terence Tai-Leung Chong. Generic consistency of the break-point estimator under specification errors. The Econometrics Journal, 6(1):167–192, June 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Chong:2006:PAA

[Cho06] Terence Tai-Leung Chong. The polynomial aggregated AR(1) model. The Econometrics Journal, 9(1):98–122, March 2006. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Cho:2020:QIN

[Cho20] Sang-Wook (Stanley) Cho. Quantifying the impact of nonpharmaceutical interventions during the COVID-19 outbreak: The case of Sweden. The Econometrics Journal, 23(3):323– 344, September 2020. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/23/3/323/5899049.

Choi:2023:FWP

[Cho23] Sung Hoon Choi. Feasible weighted projected principal component analysis for semi-parametric factor models. The Econometrics Journal, 26(2):215–234, May 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.

oup.com/ectj/article/26/2/215/6840231.

Cizek:2009:APE

[CHS09] P. Cízek, W. Härdle, and V. Spokoiny. Adaptive pointwise estimation in time-inhomogeneous conditional heteroscedasticity models. The Econometrics Journal, 12(2):248–271, July 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Casarin:2020:MNA

[CIM^{+}20] Roberto Casarin, Matteo Iacopini, German Molina, Enrique ter Horst, Ramon Espinasa, Carlos Sucre, and Roberto Rigobon.

Multilayer network analysis of oil linkages. The Econometrics Journal, 23(2):269–296, May 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.

oup.com/ectj/article/23/2/269/5717302.

Cai:2017:NRN

[CJKL17] Zongwu Cai, Bingyi Jing, Xinbing Kong, and Zhi Liu. Nonparametric regression with nearly integrated regressors under long-run dependence. The Econometrics Journal, 20(1):118–138, February 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Chang:2012:NSR

[CJP12] Yoosoon Chang, Bibo Jiang, and Joon Park. Non-stationary regression with logistic transition. The Econometrics Journal, 15 (2):255–287, June 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Clements:1998:CFP

[CK98] Michael P. Clements and Hans-Martin Krolzig. A comparison of the forecast performance of Markov-switching and threshold autoregressive models of US GNP. The Econometrics Journal, 1(1):47–75, June 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Creel:2012:EDL

[CK12] Michael Creel and Dennis Kristensen. Estimation of dynamic latent variable models using simulated non-parametric moments. The Econometrics Journal, 15(3):490–515, October 2012. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Compiani:2016:UME

[CK16] Giovanni Compiani and Yuichi Kitamura. Using mixtures in econometric models: a brief review and some new results. The Econometrics Journal, 19(3):C95–C127, October 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Chen:2021:BCC

[CL21] Le-Yu Chen and Sokbae Lee. Binary classification with covariate selection through l_{0}-penalised empirical risk minimisation. The Econometrics Journal, 24(1):103–120, January 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/24/1/103/5860277.

Casoli:2022:PTD

[CL22] Chiara Casoli and Riccardo (Jack) Lucchetti. Permanenttransitory decomposition of cointegrated time series via dynamic factor models, with an application to commodity prices. The Econometrics Journal, 25(2):494–514, May 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/25/2/494/6490126.

Chen:2014:ARF

[CLLT14] Xiaohong Chen, Sokbae Lee, Oliver Linton, and Elie Tamer. Advances in robust and flexible inference in econometrics: a special issue in honour of Joel L. Horowitz. The Econometrics Journal, 17(2):Si–Sii, June 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Chen:2014:MSE

[CLS14] Le-Yu Chen, Sokbae Lee, and Myung Jae Sung. Maximum score estimation with nonparametrically generated regressors. The Econometrics Journal, 17(3):271–300, October 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Cameron:2004:MDC

[CLTZ04] A. Colin Cameron, Tong Li, Pravin K. Trivedi, and David M. Zimmer. Modelling the differences in counted outcomes using bivariate copula models with application to mismeasured counts. The Econometrics Journal, 7(2):566–584, December 2004. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Chang:2015:NTC

[CLW15] Minsu Chang, Sokbae Lee, and Yoon-Jae Whang. Nonparametric tests of conditional treatment effects with an application to single-sex schooling on academic achievements. The Econometrics Journal, 18(3):307–346, October 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Carneiro:2020:ODC

[CLW20] Pedro Carneiro, Sokbae Lee, and Daniel Wilhelm. Optimal data collection for randomized control trials. The Econometrics Journal, 23(1):1–31, January 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/23/1/1/5614972.

Choi:2007:HUT

[CM07] Chi-Young Choi and Young-Kyu Moh. How useful are tests for unit-root in distinguishing unit-root processes from stationary but non-linear processes? The Econometrics Journal, 10(1):82– 112, March 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Chiong:2018:EGM

[CM18] Khai Xiang Chiong and Hyungsik Roger Moon. Estimation of graphical models using the L_{1,2} norm. The Econometrics Journal, 21(3):247–263, October 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Cubadda:2024:VEC

[CM24] Gianluca Cubadda and Marco Mazzali. The vector error correction index model: representation, estimation and identification. The Econometrics Journal, 27(1):126–150, January 2024. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/27/1/
126/7328937.

Camba-Mendez:2001:ALI

[CMKSW01] Gonzalo Camba-Mendez, George Kapetanios, Richard J. Smith, and Martin R. Weale. An automatic leading indicator of economic activity: forecasting GDP growth for European countries. The Econometrics Journal, 4(1):S56–S90, June 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Chernozhukov:2022:DML

[CNS22] Victor Chernozhukov, Whitney K. Newey, and Rahul Singh. Debiased machine learning of global and local parameters using regularized Riesz representers. The Econometrics Journal, 25(3): 576–601, September 2022. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/25/3/576/6572833.

Cui:2022:TSI

[CNSY22] Guowei Cui, Milda Norkute, Vasilis Sarafidis, and Takashi Yamagata. Two-stage instrumental variable estimation of linear panel data models with interactive effects. The Econometrics Journal, 25(2):340–361, May 2022. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/25/2/340/6402892.

Camponovo:2012:BPT

[CO12] Lorenzo Camponovo and Taisuke Otsu. Breakdown point theory for implied probability bootstrap. The Econometrics Journal, 15 (1):32–55, February 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Chang:2001:NEM

[CPP01] Yoosoon Chang, Joon Y. Park, and Peter C. B. Phillips. Nonlinear econometric models with cointegrated and deterministically trending regressors. The Econometrics Journal, 4(1):1–36, June 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Chudik:2011:WSC

[CPT11] Alexander Chudik, M. Hashem Pesaran, and Elisa Tosetti. Weak and strong cross-section dependence and estimation of large panels. The Econometrics Journal, 14(1):C45–C90, February 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Chesher:2014:IVR

[CR14] Andrew Chesher and Adam M. Rosen. An instrumental variable random-coefficients model for binary outcomes. The Econometrics Journal, 17(2):S1–S19, June 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Choi:2004:TLC

[CS04] In Choi and Pentti Saikkonen. Testing linearity in cointegrating smooth transition regressions. The Econometrics Journal, 7(2): 341–365, December 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Carson:2007:TMN

[CS07a] Richard T. Carson and Yixiao Sun. The tobit model with a non-zero threshold. The Econometrics Journal, 10(3):488–502, November 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Chang:2007:EIR

[CS07b] Pao-Li Chang and Shinichi Sakata. Estimation of impulse response functions using long autoregression. The Econometrics Journal, 10(2):453–469, July 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Cizek:2022:MRS

[CS22] P. Cízek and S. Sadikoglu. Misclassification-robust semiparametric estimation of single-index binary-choice models. The Econometrics Journal, 25(2):433–454, May 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/25/2/433/6517309.

Chen:2022:ENN

[CSW22] Likai Chen, Ekaterina Smetanina, and Wei Biao Wu. Estimation of nonstationary nonparametric regression model with multiplicative structure. The Econometrics Journal, 25(1):176–214, January 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
25/1/176/6297266.

Cui:2023:IES

[CSY23] Guowei Cui, Vasilis Sarafidis, and Takashi Yamagata. IV estimation of spatial dynamic panels with interactive effects: large sample theory and an application on bank attitude towards risk. The Econometrics Journal, 26(2):124–146, May 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/26/2/124/6840229.

Dahl:2002:ITL

[Dah02] Christian M. Dahl. An investigation of tests for linearity and the accuracy of likelihood based inference using random fields. The Econometrics Journal, 5(2):263–284, December 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Danilov:2005:EMU

[Dan05] Dmitry Danilov. Estimation of the mean of a univariate normal distribution when the variance is not known. The Econometrics Journal, 8(3):277–291, November 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Dastoor:2005:ASA

[Das05] Naorayex K. Dastoor. On the arbitrariness of some asymptotic test statistics based on generalized inverses. The Econometrics Journal, 8(3):292–305, November 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Davidson:2012:SIP

[Dav12] Russell Davidson. Statistical inference in the presence of heavy tails. The Econometrics Journal, 15(1):C31–C53, February 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). See discussion [Sch12a].

deChaisemartin:2023:TWF

[dCD23] Clément de Chaisemartin and Xavier D’Haultfoeuille. Twoway fixed effects and differences-in-differences with heterogeneous treatment effects: a survey. The Econometrics Journal, 26(3): C1–C30, September 2023. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/26/3/C1/6604378.

Davidson:2002:CKV

[DD02] James Davidson and Robert De Jong. Consistency of kernel variance estimators for sums of semiparametric linear processes. The Econometrics Journal, 5(1):160–175, June 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

DeBlander:2012:URT

[DD12] Rembert De Blander and Geert Dhaene. Unit root tests for panel data with AR(1) errors and small T. The Econometrics Journal, 15(1):101–124, February 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Davezies:2009:IPE

[DDF09] Laurent Davezies, Xavier D’Haultfoeuille, and Denis Fougère. Identification of peer effects using group size variation. The Econometrics Journal, 12(3):397–413, November 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Demos:2002:MDS

[Dem02] Antonis Demos. Moments and dynamic structure of a timevarying parameter stochastic volatility in mean model. The Econometrics Journal, 5(2):345–357, December 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Dardanoni:1999:ILC

[DF99] Valentino Dardanoni and Antonio Forcina. Inference for Lorenz curve orderings. The Econometrics Journal, 2(1):49–75, June 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Dufour:1998:SBF

[DFGK98] Jean-Marie Dufour, Abdeljelil Farhat, Lucien Gardiol, and Lynda Khalaf. Simulation-based finite sample normality tests in linear regressions. The Econometrics Journal, 1(1):154–173, June 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Di:2011:ESP

[DG11] Jianing Di and Ashis Gangopadhyay. On the efficiency of a semiparametric GARCH model. The Econometrics Journal, 14(2): 257–277, July 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Davydov:2019:IRN

[DG19] Youri Davydov and Francesca Greselin. Inferential results for a new measure of inequality. The Econometrics Journal, 22(2): 153–172, May 2019. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/22/2/153/5393360.

Delgado:2022:DRD

[DGSS22] Miguel A. Delgado, Andrés García-Suaza, and Pedro H. C. Sant’Anna. Distribution regression in duration analysis: an application to unemployment spells. The Econometrics Journal, 25 (3):675–698, September 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/25/3/675/6527572.

Donald:2012:ICM

[DHB12] Stephen G. Donald, Yu-Chin Hsu, and Garry F. Barrett. Incorporating covariates in the measurement of welfare and inequality: methods and applications. The Econometrics Journal, 15(1):C1– C30, February 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). See discussion [Sch12a].

Dunker:2023:NIR

[DHK23] Fabian Dunker, Stefan Hoderlein, and Hiroaki Kaido. Nonparametric identification of random coefficients in aggregate demand models for differentiated products. The Econometrics Journal, 26 (2):279–306, May 2023. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/26/2/279/6967904.

Dahl:2023:INT

[DHM23] Christian M. Dahl, Martin Huber, and Giovanni Mellace. It is never too LATE: a new look at local average treatment effects with or without defiers. The Econometrics Journal, 26(3):378– 404, September 2023. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/26/3/378/7223460.

Dong:2021:CMI

[DHS21] Baiyu Dong, Yu-Wei Hsieh, and Matthew Shum. Computing moment inequality models using constrained optimization. The Econometrics Journal, 24(3):399–416, September 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/24/3/399/6262350.

Silva:2009:PUR

[DHT09] S. De Silva, K. Hadri, and A. R. Tremayne. Panel unit root tests in the presence of cross-sectional dependence: finite sample performance and an application. The Econometrics Journal, 12 (2):340–366, July 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Dietrich:2001:LDR

[Die01] Franz K. Dietrich. The limiting distribution of the t-ratio for the unit root test in an AR(1). The Econometrics Journal, 4(2): 242–256, December 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Demetrescu:2009:TCR

[DLS09] Matei Demetrescu, Helmut Lütkepohl, and Pentti Saikkonen. Testing for the cointegrating rank of a vector autoregressive process with uncertain deterministic trend term. The Econometrics Journal, 12(3):414–435, November 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Dolado:2004:AIR

[DM04] Juan J. Dolado and Francesc Marmol. Asymptotic inference results for multivariate long-memory processes. The Econometrics Journal, 7(1):168–190, June 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Davidson:2007:MIJ

[DM07] Russell Davidson and James G. MacKinnon. Moments of IV and JIVE estimators. The Econometrics Journal, 10(3):541–553, November 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Davidson:2008:BIL

[DM08] Russell Davidson and James G. MacKinnon. Bootstrap inference in a linear equation estimated by instrumental variables. The Econometrics Journal, 11(3):443–477, November 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Davidson:2014:CSB

[DM14] Russell Davidson and James G. MacKinnon. Confidence sets based on inverting Anderson-Rubin tests. The Econometrics Journal, 17(2):S39–S58, June 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Druedahl:2020:HOI

[DMN20] Jeppe Druedahl and Anders Munk-Nielsen. Higher-order income dynamics with linked regression trees. The Econometrics Journal, 23(3):S25–S58, September 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.

oup.com/ectj/article/23/3/S25/5899050.

Castro:2004:CSA

[DO04] Tomas Del Barrio Castro and Denise R. Osborn. The consequences of seasonal adjustment for periodic autoregressive processes. The Econometrics Journal, 7(2):307–321, December 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Deng:2006:CAA

[DP06] Ai Deng and Pierre Perron. A comparison of alternative asymptotic frameworks to analyse a structural change in a linear time trend. The Econometrics Journal, 9(3):423–447, November 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Dustmann:2007:SCP

[DRB07] Christian Dustmann and María Engracia Rochina-Barrachina. Selection correction in panel data models: An application to the estimation of females’ wage equations. The Econometrics Journal, 10(2):263–293, July 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Dimpfl:2022:ESC

[DSBG22] Thomas Dimpfl, Jantje Sönksen, Ingo Bechmann, and Joachim Grammig. Estimating the SARS-CoV-2 infection fatality rate by data combination: the case of Germany’s first wave. The Econometrics Journal, 25(2):515–530, May 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/25/2/515/6517310.

Daraio:2018:CLT

[DSW18] Cinzia Daraio, Léopold Simar, and Paul W. Wilson. Central limit theorems for conditional efficiency measures and tests of the ‘separability’ condition in non-parametric, two-stage models of production. The Econometrics Journal, 21(2):170–191, June 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Deb:2006:SSL

[DT06] Partha Deb and Pravin K. Trivedi. Specification and simulated likelihood estimation of a non-normal treatment-outcome model with selection: Application to health care utilization. The Econometrics Journal, 9(2):307–331, July 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Du:2016:NBT

[Du16] Zaichao Du. Nonparametric bootstrap tests for independence of generalized errors. The Econometrics Journal, 19(1):55–83, February 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Dai:2024:ELC

[DUM24] Runyu Dai, Yoshimasa Uematsu, and Yasumasa Matsuda. Estimation of large covariance matrices with mixed factor structures. The Econometrics Journal, 27(1):62–83, January 2024. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/27/1/62/7284383.

Dellaportas:2007:MVA

[DV07] P. Dellaportas and I. D. Vrontos. Modelling volatility asymmetries: a Bayesian analysis of a class of tree structured multivariate GARCH models. The Econometrics Journal, 10(3):503–520, November 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Demetrescu:2019:TCC

[DW19] Matei Demetrescu and Dominik Wied. Testing for constant correlation of filtered series under structural change. The Econometrics Journal, 22(1):10–33, January 2019. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:/

/academic.oup.com/ectj/article/22/1/10/5317156.

Eren:2008:IHS

[EH08] Ozkan Eren and Daniel J. Henderson. The impact of homework on student achievement. The Econometrics Journal, 11(2):326– 348, July 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Eitrheim:2002:PFF

[EJN02] Øyvind Eitrheim, Eilev Jansen, and Ragnar Nymoen. Progress from forecast failure — the Norwegian consumption function. The Econometrics Journal, 5(1):40–64, June 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Elliott:2008:MBR

[EKS08] Robert J. Elliott, Vikram Krishnamurthy, and Jörn Sass. Moment based regression algorithms for drift and volatility estimation in continuous-time Markov switching models. The Econometrics Journal, 11(2):244–270, July 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Ericsson:1998:FEF

[EM98] Neil R. Ericsson and Jaime Marquez. A framework for economic forecasting. The Econometrics Journal, 1(1):203–227, June 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Ericsson:2002:DEC

[EM02] Neil R. Ericsson and James G. MacKinnon. Distributions of error correction tests for cointegration. The Econometrics Journal, 5 (2):285–318, December 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Emvalomatis:2020:ESP

[Emv20a] Grigorios Emvalomatis. Erratum to: Semi-parametric analysis of efficiency and productivity using Gaussian processes. The Econometrics Journal, 23(1):176, January 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/23/1/176/5614492. See

[Emv20b].

Emvalomatis:2020:SPA

[Emv20b] Grigorios Emvalomatis. Semi-parametric analysis of efficiency and productivity using Gaussian processes. The Econometrics Journal, 23(1):48–67, January 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/23/1/48/5561435. See erra-

tum [Emv20a].

Engler:2009:EPA

[EN09] Eric Engler and Bent Nielsen. The empirical process of autoregressive residuals. The Econometrics Journal, 12(2):367–381, July 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Engsted:2012:TRB

[EN12] Tom Engsted and Bent Nielsen. Testing for rational bubbles in a coexplosive vector autoregression. The Econometrics Journal, 15 (2):226–254, June 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Escanciano:2018:SRE

[Esc18] Juan Carlos Escanciano. A simple and robust estimator for linear regression models with strictly exogenous instruments. The Econometrics Journal, 21(1):36–54, February 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Everaert:2013:OBM

[Eve13] Gerdie Everaert. Orthogonal to backward mean transformation for dynamic panel data models. The Econometrics Journal, 16 (2):179–221, June 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Fanelli:2006:DAC

[Fan06] Luca Fanelli. Dynamic adjustment cost models with forwardlooking behaviour. The Econometrics Journal, 9(1):23–47, March 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Feve:2010:PNP

[FF10] Frédérique Fève and Jean-Pierre Florens. The practice of nonparametric estimation by solving inverse problems: the example of transformation models. The Econometrics Journal, 13(3):S1– S27, October 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Fan:2003:SEV

[FG03] Jianqing Fan and Juan Gu. Semiparametric estimation of value at risk. The Econometrics Journal, 6(2):261–290, December 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Farbmacher:2022:CMA

[FHL^{+}22] Helmut Farbmacher, Martin Huber, Lukás Lafférs, Henrika Langen, and Martin Spindler. Causal mediation analysis with double machine learning. The Econometrics Journal, 25(2):277–300, May 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
25/2/277/6517682.

Fiorio:2010:BRI

[FHP10] Carlo V. Fiorio, Vassilis A. Hajivassiliou, and Peter C. B. Phillips. Bimodal t-ratios: the impact of thick tails on inference. The Econometrics Journal, 13(2):271–289, July 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Fu:2021:TEN

[FHP21] Jia-Young Michael Fu, Joel L. Horowitz, and Matthias Parey. Testing exogeneity in nonparametric instrumental variables models identified by conditional quantile restrictions. The Econometrics Journal, 24(1):23–40, January 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:/

/academic.oup.com/ectj/article/24/1/23/5819016.

Forchini:2019:FIP

[FJ19] Giovanni Forchini and Bin Jiang. Fragility of identification in panel binary response models. The Econometrics Journal, 22(3):

282–291, September 2019. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/22/3/282/5541058.

Florens:2012:IRP

[FJV12] Jean-Pierre Florens, Jan Johannes, and Sébastien Van Bellegem. Instrumental regression in partially linear models. The Econometrics Journal, 15(2):304–324, June 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Fosgerau:2021:ICI

[FK21] Mogens Fosgerau and Dennis Kristensen. Identification of a class of index models: a topological approach. The Econometrics Journal, 24(1):121–133, January 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/24/1/121/5858893.

Fan:2016:OEL

[FLL16] Jianqing Fan, Yuan Liao, and Han Liu. An overview of the estimation of large covariance and precision matrices. The Econometrics Journal, 19(1):C1–C32, February 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Lee:2010:SES

[fLLL10] Lung fei Lee, Xiaodong Liu, and Xu Lin. Specification and estimation of social interaction models with network structures. The Econometrics Journal, 13(2):145–176, July 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Lee:2020:ICD

[fLY20] Lung fei Lee and Jihai Yu. Initial conditions of dynamic panel data models: on within and between equations. The Econometrics Journal, 23(1):115–136, January 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:/

/academic.oup.com/ectj/article/23/1/115/5556939.

Ferriani:2012:ETN

[FP12] Fabrizio Ferriani and Sergio Pastorello. Estimating and testing non-affine option pricing models with a large unbalanced panel of options. The Econometrics Journal, 15(2):171–203, June 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Fan:2015:MPE

[FPR15] Yanqin Fan, Sergio Pastorello, and Eric Renault. Maximization by parts in extremum estimation. The Econometrics Journal, 18 (2):147–171, June 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Frolich:2006:NPR

[Frö06] Markus Frölich. Non-parametric regression for binary dependent variables. The Econometrics Journal, 9(3):511–540, November 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Frolich:2007:PSM

[Frö07] Markus Frölich. Propensity score matching without conditional independence assumption — with an application to the gender wage gap in the United Kingdom. The Econometrics Journal, 10 (2):359–407, July 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Fruhwirth-Schnatter:2004:EML

[FS04] Sylvia Frühwirth-Schnatter. Estimating marginal likelihoods for mixture and Markov switching models using bridge sampling techniques. The Econometrics Journal, 7(1):143–167, June 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Fan:2009:RES

[FS09] Jianqing Fan and Richard J. Smith. Royal economic society annual conference 2008 special issue on financial econometrics. The Econometrics Journal, 12(3):ci–ciii, November 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Franses:2000:DOD

[FT00] Philip Hans Franses and A. M. Robert Taylor. Determining the order of differencing in seasonal time series processes. The Econometrics Journal, 3(2):250–264, December 2000. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Fernandez-Val:2021:LRA

[FVFW21] Iván Fernández-Val, Hugo Freeman, and Martin Weidner. Lowrank approximations of nonseparable panel models. The Econometrics Journal, 24(2):C40–C77, May 2021. CODEN ????

ISSN 1368-4221 (print), 1368-423X (electronic). URL http:/

/academic.oup.com/ectj/article/24/2/C40/6177679.

Fuleky:2014:IIB

[FZ14] Peter Fuleky and Eric Zivot. Indirect inference based on the score. The Econometrics Journal, 17(3):383–393, October 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Galichon:2017:SSR

[Gal17] Alfred Galichon. A survey of some recent applications of optimal transport methods to econometrics. The Econometrics Journal, 20(2):C1–C11, June 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Gardner:2020:RMB

[Gar20] John R. Gardner. Roy-model bounds on the wage effects of the Great Migration. The Econometrics Journal, 23(1):68–87, January 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
23/1/68/5556528.

Georgiev:2007:MDF

[Geo07] Iliyan Georgiev. A mixture-distribution factor model for multivariate outliers. The Econometrics Journal, 10(3):605–636, November 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Gu:2009:BER

[GFCK09] Yuanyuan Gu, Denzil G. Fiebig, Edward Cripps, and Robert Kohn. Bayesian estimation of a random effects heteroscedastic probit model. The Econometrics Journal, 12(2):324–339, July 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Gao:2006:SET

[GH06] Jiti Gao and Kim Hawthorne. Semiparametric estimation and testing of the trend of temperature series. The Econometrics Journal, 9(2):332–355, July 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Gotz:2021:LMF

[GH21] Thomas B. Götz and Klemens Hauzenberger. Large mixedfrequency VARs with a parsimonious time-varying parameter structure. The Econometrics Journal, 24(3):442–461, September 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
24/3/442/6102556.

Giacomini:2015:ETF

[Gia15] Raffaella Giacomini. Economic theory and forecasting: lessons from the literature. The Econometrics Journal, 18(2):C22–C41, June 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Goldman:2018:NPI

[GK18] Matt Goldman and David M. Kaplan. Non-parametric inference on (conditional) quantile differences and interquantile ranges, using L-statistics. The Econometrics Journal, 21(2):136–169, June 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Guimaraes:2007:COG

[GL07] Paulo Guimarães and Richard C. Lindrooth. Controlling for overdispersion in grouped conditional logit models: a computationally simple application of Dirichlet-multinomial regression. The Econometrics Journal, 10(2):439–452, July 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Grigoletto:2009:LSF

[GL09] Matteo Grigoletto and Francesco Lisi. Looking for skewness in financial time series. The Econometrics Journal, 12(2):310–323, July 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Ginker:2021:LRT

[GL21] Tim Ginker and Offer Lieberman. LSTUR regression theory and the instability of the sample correlation coefficient between financial return indices. The Econometrics Journal, 24(1):58–82, January 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
24/1/58/5846040.

Gallo:1998:SME

[GM98] Giampiero M. Gallo and Grayham E. Mizon. Simulation methods in econometrics: Editors’ introduction. The Econometrics Journal, 1(1):i–vii, June 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Grammig:2000:NMH

[GM00] Joachim Grammig and Kai-Oliver Maurer. Non-monotonic hazard functions and the autoregressive conditional duration model. The Econometrics Journal, 3(1):16–38, June 2000. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Gabriel:2004:FAA

[GM04] Vasco J. Gabriel and Luis F. Martins. On the forecasting ability of ARFIMA models when infrequent breaks occur. The Econometrics Journal, 7(2):455–475, December 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Gillen:2019:BAD

[GMMS19] Benjamin J. Gillen, Sergio Montero, Hyungsik Roger Moon, and Matthew Shum. BLP-2LASSO for aggregate discrete choice models with rich covariates. The Econometrics Journal, 22(3):262– 281, September 2019. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/22/3/262/5531304.

Godfrey:2000:CSL

[GO00] Leslie G. Godfrey and Chris D. Orme. Controlling the significance levels of prediction error tests for linear regression models. The Econometrics Journal, 3(1):66–83, June 2000. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Gorgens:2006:SES

[Gør06] Tue Gørgens. Semiparametric estimation of single-index hazard functions without proportional hazards. The Econometrics Journal, 9(1):1–22, March 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Gortz:2012:RSM

[Gör12] Christoph Görtz. A review of Structural Macroeconometrics by DeJong (David N.) and Dave (Chetan). The Econometrics Journal, 15(3):B5–B10, October 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Gospodinov:2004:ACI

[Gos04] Nikolay Gospodinov. Asymptotic confidence intervals for impulse responses of near-integrated processes. The Econometrics Journal, 7(2):505–527, December 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Godfrey:2006:SBT

[GOS06] L. G. Godfrey, C. D. Orme, and J. M. C. Santos Silva. Simulationbased tests for heteroskedasticity in linear regression models: Some further results. The Econometrics Journal, 9(1):76–97, March 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Greene:2004:BML

[Gre04] William Greene. The behaviour of the maximum likelihood estimator of limited dependent variable models in the presence of fixed effects. The Econometrics Journal, 7(1):98–119, June 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Guo:2013:CCT

[GS13] Zheng-Feng Guo and Mototsugu Shintani. Consistent co-trending rank selection when both stochastic and non-linear deterministic trends are present. The Econometrics Journal, 16(3):473–484, October 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Gu:2018:OAF

[GS18] Jiaying Gu and Shu Shen. Oracle and adaptive false discovery rate controlling methods for one-sided testing: theory and application in treatment effect evaluation. The Econometrics Journal, 21(1):11–35, February 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Guo:2019:TME

[GSW19] Gangzheng Guo, Yixiao Sun, and Shaoping Wang. Testing for moderate explosiveness. The Econometrics Journal, 22(1):73–95, January 2019. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
22/1/73/5257425.

Granger:1999:DML

[GT99] Clive Granger and Allan Timmermann. Data mining with local model specification uncertainty: a discussion of Hoover and perez. The Econometrics Journal, 2(2):220–225, December 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Gautier:2020:IPB

[GvV20] Pieter A. Gautier and Aico van Vuuren. Identifying present bias and time preferences with an application to land-lease-contract data. The Econometrics Journal, 23(3):363–385, September 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/23/3/
363/5861310.

Gorgens:2012:TPF

[GW12] Tue Gørgens and Allan Würtz. Testing a parametric function against a non-parametric alternative in IV and GMM settings. The Econometrics Journal, 15(3):462–489, October 2012. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Goh:2022:SCM

[GY22] Gyuhyeong Goh and Jisang Yu. Synthetic control method with convex hull restrictions: a Bayesian maximum a posteriori approach. The Econometrics Journal, 25(1):215–232, January 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/25/1/
215/6270888.

Hadri:2000:TSH

[Had00] Kaddour Hadri. Testing for stationarity in heterogeneous panel data. The Econometrics Journal, 3(2):148–161, December 2000. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hafner:2009:CFT

[Haf09] Christian M. Hafner. Causality and forecasting in temporally aggregated multivariate GARCH processes. The Econometrics Journal, 12(1):127–146, March 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hagemann:2014:SEN

[Hag14] Andreas Hagemann. Stochastic equicontinuity in nonlinear time series models. The Econometrics Journal, 17(1):188–196, February 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Halliday:2008:HSD

[Hal08] Timothy J. Halliday. Heterogeneity, state dependence and health. The Econometrics Journal, 11(3):499–516, November 2008. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hand:1999:DCD

[Han99a] David J. Hand. Discussion contribution on ‘Data mining reconsidered: encompassing and the general-to-specific approach to specification search’ by Hoover and Perez. The Econometrics Journal, 2(2):241–243, December 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). See [HP99a].

Hansen:1999:DDM

[Han99b] Bruce E. Hansen. Discussion of ‘data mining reconsidered’. The Econometrics Journal, 2(2):192–201, December 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). See [HP99a].

Hansen:2005:GRT

[Han05] Peter Reinhard Hansen. Granger’s representation theorem: a closed-form expression for I(1) processes. The Econometrics Journal, 8(1):23–38, March 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hansen:2022:RCE

[Han22] Peter Reinhard Hansen. Relative contagiousness of emerging virus variants: an analysis of the Alpha, Delta, and Omicron SARS-CoV-2 variants. The Econometrics Journal, 25(3):739– 761, September 2022. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/25/3/739/6553812.

He:2020:WBF

[HB20] Yang He and Otávio Bartalotti. Wild bootstrap for fuzzy regression discontinuity designs: obtaining robust bias-corrected confidence intervals. The Econometrics Journal, 23(2):211–231, May 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
23/2/211/5715907.

Hendry:2004:PF

[HC04] David F. Hendry and Michael P. Clements. Pooling of forecasts. The Econometrics Journal, 7(1):1–31, June 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hubner:2019:QBS

[HC19] Stefan Hubner and Pavel Cízek. Quantile-based smooth transition value at risk estimation. The Econometrics Journal, 22(3): 241–261, September 2019. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/22/3/241/5511890.

Honore:2021:ISB

[HdP21] Bo E. Honoré and Áureo de Paula. Identification in simple binary outcome panel data models. The Econometrics Journal, 24(2): C78–C93, May 2021. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/24/2/C78/6271316.

Hafner:2000:TLA

[HH00] Christian M. Hafner and Helmut Herwartz. Testing for linear autoregressive dynamics under heteroskedasticity. The Econometrics Journal, 3(2):177–197, December 2000. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Honore:2010:ETM

[HH10] Bo E. Honoré and Luojia Hu. Estimation of a transformation model with truncation, interval observation and time-varying covariates. The Econometrics Journal, 13(1):127–144, February 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Honore:2018:SBE

[HH18] Bo E. Honoré and Luojia Hu. Simpler bootstrap estimation of the asymptotic variance of U-statistic-based estimators. The Econometrics Journal, 21(1):1–10, February 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hahn:2004:EWI

[HHK04] Jinyong Hahn, Jerry Hausman, and Guido Kuersteiner. Estimation with weak instruments: Accuracy of higher-order bias and MSE approximations. The Econometrics Journal, 7(1):272–306, June 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hitomi:2022:OMR

[HIN22] Kohtaro Hitomi, Masamune Iwasawa, and Yoshihiko Nishiyama. Optimal minimax rates against nonsmooth alternatives. The Econometrics Journal, 25(2):322–339, May 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/25/2/322/6380482.

Hansen:1999:STP

[HJ99] Henrik Hansen and Søren Johansen. Some tests for parameter constancy in cointegrated VAR-models. The Econometrics Journal, 2(2):306–333, December 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hui:2005:RMD

[HJ05] Yer Van Hui and Jiancheng Jiang. Robust modelling of DTARCH models. The Econometrics Journal, 8(2):143–158, July 2005. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hong:2021:UIS

[HJJX21] Shaoxin Hong, Jiancheng Jiang, Xuejun Jiang, and Zhijie Xiao. Unifying inference for semiparametric regression. The Econometrics Journal, 24(3):482–501, September 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/24/3/482/6168403.

Hendry:1999:IDM

[HK99] David F. Hendry and Hans-Martin Krolzig. Improving on ‘Data mining reconsidered’ by K. D. Hoover and S. J. Perez. The Econometrics Journal, 2(2):202–219, December 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). See [HP99a].

Harvey:2000:SEF

[HK00] Andrew Harvey and Siem Jan Koopman. Signal extraction and the formulation of unobserved components models. The Econometrics Journal, 3(1):84–107, June 2000. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hsu:2001:DBT

[HK01] Chih-Chiang Hsu and Chung-Ming Kuan. Distinguishing between trend-break models: method and empirical evidence. The Econometrics Journal, 4(2):171–190, December 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Haile:2019:UHA

[HK19] Philip A. Haile and Yuichi Kitamura. Unobserved heterogeneity in auctions. The Econometrics Journal, 22(1):–, January 2019. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/22/1/
C1/5257424.

Heckman:2021:USM

[HK21] James J. Heckman and Ganesh Karapakula. Using a satisficing model of experimenter decision-making to guide finite-sample inference for compromised experiments. The Econometrics Journal, 24(2):C1–C39, May 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/24/2/C1/6207935.

Haug:2007:MME

[HKLZ07] S. Haug, C. Klüppelberg, A. Lindner, and M. Zapp. Method of moment estimation in the COGARCH(1,1) model. The Econometrics Journal, 10(2):320–341, July 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hadri:2015:NPC

[HKR15] Kaddour Hadri, Eiji Kurozumi, and Yao Rao. Novel panel cointegration tests emending for cross-section dependence with N fixed. The Econometrics Journal, 18(3):363–411, October 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hornok:2000:FSD

[HL00] Attila Hornok and Rolf Larsson. The finite sample distribution of the KPSS test. The Econometrics Journal, 3(1):108–121, June 2000. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hadri:2005:TSH

[HL05a] Kaddour Hadri and Rolf Larsson. Testing for stationarity in heterogeneous panel data where the time dimension is finite. The Econometrics Journal, 8(1):55–69, March 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Harvey:2005:TUR

[HL05b] David I. Harvey and Stephen J. Leybourne. On testing for unit roots and the initial observation. The Econometrics Journal, 8 (1):97–111, March 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Harvey:2007:TTS

[HL07] David I. Harvey and Stephen J. Leybourne. Testing for time series linearity. The Econometrics Journal, 10(1):149–165, March 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hong:2020:IFP

[HLL20] Han Hong, Michael P. Leung, and Jessie Li. Inference on finite-population treatment effects under limited overlap. The Econometrics Journal, 23(1):32–47, January 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/23/1/32/5558232.

Harvey:2001:APU

[HLN01] David I. Harvey, Stephen J. Leybourne, and Paul Newbold. Analysis of a panel of UK macroeconomic forecasts. The Econometrics Journal, 4(1):S37–S55, June 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Horvath:2022:DCB

[HLRZ22] Lajos Horváth, Zhenya Liu, Gregory Rice, and Yuqian Zhao. Detecting common breaks in the means of high dimensional crossdependent panels. The Econometrics Journal, 25(2):362–383, May 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
25/2/362/6364356.

Harvey:2008:SUR

[HLT08] David I. Harvey, Stephen J. Leybourne, and A. M. Robert Taylor. Seasonal unit root tests and the role of initial conditions. The Econometrics Journal, 11(3):409–442, November 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hoderlein:2009:IEL

[HM09] Stefan Hoderlein and Enno Mammen. Identification and estimation of local average derivatives in non-separable models without monotonicity. The Econometrics Journal, 12(1):1–25, March 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Henry:2013:SIL

[HM13] Marc Henry and Ismael Mourifié. Set inference in latent variables models. The Econometrics Journal, 16(1):S93–S105, February 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hoderlein:2011:NPM

[HMY11] Stefan Hoderlein, Enno Mammen, and Kyusang Yu. Nonparametric models in binary choice fixed effects panel data. The Econometrics Journal, 14(3):351–367, October 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hoga:2017:TCE

[Hog17] Yannick Hoga. Testing for changes in (extreme) VaR. The Econometrics Journal, 20(1):23–51, February 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hoover:1999:DMR

[HP99a] Kevin D. Hoover and Stephen J. Perez. Data mining reconsidered: encompassing and the general-to-specific approach to specification search. The Econometrics Journal, 2(2):167–191, December 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). See comments [Han99b, HK99, Han99a].

Hoover:1999:ROD

[HP99b] Kevin D. Hoover and Stephen J. Perez. Reply to our discussants. The Econometrics Journal, 2(2):244–247, December 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Harris:2004:DCR

[HP04] D. Harris and D. S. Poskitt. Determination of cointegrating rank in partially non-stationary processes via a generalised von-Neumann criterion. The Econometrics Journal, 7(1):191–217, June 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Haan:2010:SAE

[HP10] Peter Haan and Victoria Prowse. A structural approach to estimating the effect of taxation on the labour market dynamics of older workers. The Econometrics Journal, 13(3):S99–S125, October 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hendry:1998:EJR

[HS98] David F. Hendry and Neil Shephard. The econometrics journal of the Royal economic society: Foreword. The Econometrics Journal, 1(2):i–ii, December 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Heaton:2004:ICF

[HS04] Chris Heaton and Victor Solo. Identification of causal factor models of stationary time series. The Econometrics Journal, 7 (2):618–627, December 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hsu:2017:MST

[HS17] Yu-Chin Hsu and Xiaoxia Shi. Model-selection tests for conditional moment restriction models. The Econometrics Journal, 20 (1):52–85, February 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hu:2018:IES

[HS18] Yingyao Hu and Ji-Liang Shiu. Identification and estimation of semi-parametric censored dynamic panel data models of short time periods. The Econometrics Journal, 21(1):55–85, February 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hsu:2017:CTC

[Hsu17] Yu-Chin Hsu. Consistent tests for conditional treatment effects. The Econometrics Journal, 20(1):1–22, February 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hu:2015:IES

[HSW15] Yingyao Hu, Ji-Liang Shiu, and Tiemen Woutersen. Identification and estimation of single-index models with measurement error and endogeneity. The Econometrics Journal, 18(3):347–362, October 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Huang:2008:PVA

[Hua08] Xiao Huang. Panel vector autoregression under cross-sectional dependence. The Econometrics Journal, 11(2):219–243, July 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Huang:2011:QML

[Hua11] Xiao Huang. Quasi-maximum likelihood estimation of discretely observed diffusions. The Econometrics Journal, 14(2):241–256, July 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hounyo:2016:VEE

[HV16] Ulrich Hounyo and Bezirgen Veliyev. Validity of Edgeworth expansions for realized volatility estimators. The Econometrics Journal, 19(1):1–32, February 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Hsiao:2007:LAT

[HW07] Cheng Hsiao and Siyan Wang. Lag-augmented two- and threestage least squares estimators for integrated structural dynamic models. The Econometrics Journal, 10(1):49–81, March 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Huang:2008:EGM

[HWY08] Da Huang, Hansheng Wang, and Qiwei Yao. Estimating GARCH models: when to use what? The Econometrics Journal, 11(1): 27–38, March 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Han:2012:NSN

[HZ12] Heejoon Han and Shen Zhang. Non-stationary non-parametric volatility model. The Econometrics Journal, 15(2):204–225, June 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Igami:2020:AIS

[Iga20] Mitsuru Igami. Artificial intelligence as structural estimation: Deep Blue, Bonanza, and AlphaGo. The Econometrics Journal, 23(3):S1–S24, September 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/23/3/S1/5802896.

ilKim:2007:UCR

[iK07] Kyoo il Kim. Uniform convergence rate of the seminonparametric density estimator and testing for similarity of two unknown densities. The Econometrics Journal, 10(1):1–34, March 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Ikefuji:2024:RPP

[IMY24] Masako Ikefuji, Jan R. Magnus, and Takashi Yamagata. Revealing priors from posteriors with an application to inflation forecasting in the UK. The Econometrics Journal, 27(1):151–170, January 2024. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
27/1/151/7288648.

Inoue:2006:BAM

[Ino06] Atsushi Inoue. A bootstrap approach to moment selection. The Econometrics Journal, 9(1):48–75, March 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Ioannidis:2005:RBB

[Ioa05] Evangelos E. Ioannidis. Residual-based block bootstrap unit root testing in the presence of trend breaks. The Econometrics Journal, 8(3):323–351, November 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Iskhakov:2020:E

[IRS20a] Fedor Iskhakov, John Rust, and Bertel Schjerning. Editorial. The Econometrics Journal, 23(3):Si–Siii, September 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/23/3/Si/5925025.

Iskhakov:2020:MLS

[IRS20b] Fedor Iskhakov, John Rust, and Bertel Schjerning. Machine learning and structural econometrics: contrasts and synergies. The Econometrics Journal, 23(3):S81–S124, September 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/23/3/
S81/5899047.

Carrion-i-Silvestre:2005:BPA

[iSDLB05] Josep Lluís Carrion i Silvestre, Tomás Del Barrio-Castro, and Enrique López-Bazo. Breaking the panels: an application to the GDP per capita. The Econometrics Journal, 8(2):159–175, July 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Iskhakov:2010:SDM

[Isk10] Fedor Iskhakov. Structural dynamic model of retirement with latent health indicator. The Econometrics Journal, 13(3):S126– S161, October 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Carrion-i-Silvestre:2006:JHS

[iSS06] Josep Lluís Carrion i Silvestre and Andreu Sansó. Joint hypothesis specification for unit root tests with a structural break. The Econometrics Journal, 9(2):196–224, July 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Jiang:2014:WCQ

[JJS14] Jiancheng Jiang, Xuejun Jiang, and Xinyuan Song. Weighted composite quantile regression estimation of DTARCH models. The Econometrics Journal, 17(1):1–23, February 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Jiang:2010:EEM

[JK10] George J. Jiang and John L. Knight. ECF estimation of Markov models where the transition density is unknown. The Econometrics Journal, 13(2):245–270, July 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Jungbacker:2015:LBD

[JK15] Borus Jungbacker and Siem Jan Koopman. Likelihood-based dynamic factor analysis for measurement and forecasting. The Econometrics Journal, 18(2):C1–C21, June 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Jensen:2001:NSM

[JL01] Morten B. Jensen and Asger Lunde. The NIG–S&ARCH model: a fat-tailed, stochastic, and autoregressive conditional heteroskedastic volatility model. The Econometrics Journal, 4 (2):319–342, December 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Jacobson:2008:IER

[JLLN08] Tor Jacobson, Johan Lyhagen, Rolf Larsson, and Marianne Nessén. Inflation, exchange rates and PPP in a multivariate panel cointegration model. The Econometrics Journal, 11(1): 58–79, March 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Junker:2005:MAR

[JM05] Markus Junker and Angelika May. Measurement of aggregate risk with copulas. The Econometrics Journal, 8(3):428–454, November 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Jochmans:2017:NSB

[JM17] Koen Jochmans and Thierry Magnac. A note on sufficiency in binary panel models. The Econometrics Journal, 20(2):259–269, June 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Johansen:2000:CAP

[JMN00] Søren Johansen, Rocco Mosconi, and Bent Nielsen. Cointegration analysis in the presence of structural breaks in the deterministic trend. The Econometrics Journal, 3(2):216–249, December 2000. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Jochmans:2013:PCE

[Joc13] Koen Jochmans. Pairwise-comparison estimation with nonparametric controls. The Econometrics Journal, 16(3):340–372, October 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Jochmans:2014:FDP

[Joc14] Koen Jochmans. First-differencing in panel data models with incidental functions. The Econometrics Journal, 17(3):373–382, October 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Jakobsen:2022:DRC

[JP22] Martin Emil Jakobsen and Jonas Peters. Distributional robustness of K-class estimators and the PULSE. The Econometrics Journal, 25(2):404–432, May 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.

oup.com/ectj/article/25/2/404/6380481.

Jun:2012:DEV

[JPX12] Sung Jae Jun, Joris Pinkse, and Haiqing Xu. Discrete endogenous variables in weakly separable models. The Econometrics Journal, 15(2):288–303, June 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Jun:2016:ENT

[JPX16] Sung Jae Jun, Joris Pinkse, and Haiqing Xu. Estimating a nonparametric triangular model with binary endogenous regressors. The Econometrics Journal, 19(2):113–149, June 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Johansen:2004:MTE

[JS04] Søren Johansen and Anders Rygh Swensen. More on testing exact rational expectations in cointegrated vector autoregressive models: Restricted constant and linear term. The Econometrics Journal, 7(2):389–397, December 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Juhl:2005:FCM

[Juh05] Ted Juhl. Functional-coefficient models under unit root behaviour. The Econometrics Journal, 8(2):197–213, July 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Jensen:2012:EEV

[JW12] Peter S. Jensen and Allan H. Würtz. Estimating the effect of a variable in a high-dimensional linear model. The Econometrics Journal, 15(2):325–357, June 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Juodis:2019:OPU

[JW19] Arturas Juodis and Joakim Westerlund. Optimal panel unit root testing with covariates. The Econometrics Journal, 22(1):57–72, January 2019. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
22/1/57/5257426.

Kalliovirta:2012:MTB

[Kal12] Leena Kalliovirta. Misspecification tests based on quantile residuals. The Econometrics Journal, 15(2):358–393, June 2012. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kamionka:1998:SML

[Kam98] Thierry Kamionka. Simulated maximum likelihood estimation in transition models. The Econometrics Journal, 1(1):129–153, June 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kang:2014:ESS

[Kan14] Kyu H. Kang. Estimation of state-space models with endogenous Markov regime-switching parameters. The Econometrics Journal, 17(1):56–82, February 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kapetanios:2008:BPP

[Kap08] G. Kapetanios. A bootstrap procedure for panel data sets with many cross-sectional units. The Econometrics Journal, 11(2): 377–395, July 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kasy:2015:NPI

[Kas15] Maximilian Kasy. Non-parametric inference on the number of equilibria. The Econometrics Journal, 18(1):1–39, February 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kaufmann:2000:MBC

[Kau00] Sylvia Kaufmann. Measuring business cycles with a dynamic Markov switching factor model: an assessment using Bayesian simulation methods. The Econometrics Journal, 3(1):39–65, June 2000. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kawakatsu:2007:NIB

[Kaw07] Hiroyuki Kawakatsu. Numerical integration-based Gaussian mixture filters for maximum likelihood estimation of asymmetric stochastic volatility models. The Econometrics Journal, 10(2): 342–358, July 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kheifets:2015:STN

[Khe15] Igor L. Kheifets. Specification tests for nonlinear dynamic models. The Econometrics Journal, 18(1):67–94, February 2015. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kilic:2004:LTS

[Kil04] Rehim Kiliç. Linearity tests and stationarity. The Econometrics Journal, 7(1):55–62, June 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kim:2014:CBT

[Kim14] Dukpa Kim. Common breaks in time trends for large panel data with a factor structure. The Econometrics Journal, 17(3):301– 337, October 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kim:2020:PIN

[Kim20] Dongwoo Kim. Partial identification in nonseparable count data instrumental variable models. The Econometrics Journal, 23(2): 232–250, May 2020. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/23/2/232/5682487.

Kiviet:2013:IIS

[Kiv13] Jan F. Kiviet. Identification and inference in a simultaneous equation under alternative information sets and sampling schemes. The Econometrics Journal, 16(1):S24–S59, February 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kondo:2003:HPI

[KjL03] Yasushi Kondo and Myoung jae Lee. Hedonic price index estimation under mean-independence of time dummies from quality characteristics. The Econometrics Journal, 6(1):28–45, June 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Karanasos:2003:MAE

[KK03] M. Karanasos and J. Kim. Moments of the ARMA–EGARCH model. The Econometrics Journal, 6(1):146–166, June 2003. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kleiber:2005:FSP

[KK05] Christian Kleiber and Walter Krämer. Finite-sample power of the Durbin–Watson test against fractionally integrated disturbances. The Econometrics Journal, 8(3):406–417, November 2005. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kim:2011:DET

[KK11] Yunmi Kim and Chang-Jin Kim. Dealing with endogeneity in a time-varying parameter model: joint estimation and two-step estimation procedures. The Econometrics Journal, 14(3):487– 497, October 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kawakatsu:2009:EAO

[KL09] Hiroyuki Kawakatsu and Ann G. Largey. EM algorithms for ordered probit models with endogenous regressors. The Econometrics Journal, 12(1):164–186, March 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Knaus:2021:MLE

[KLS21] Michael C. Knaus, Michael Lechner, and Anthony Strittmatter. Machine learning estimation of heterogeneous causal effects: Empirical Monte Carlo evidence. The Econometrics Journal, 24(1): 134–161, January 2021. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/24/1/134/5854188.

Kim:2004:TSQ

[KM04] Tae-Hwan Kim and Christophe Muller. Two-stage quantile regression when the first stage is based on quantile regression. The Econometrics Journal, 7(1):218–231, June 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Karaca-Mandic:2003:SEC

[KMT03] Pinar Karaca-Mandic and Kenneth Train. Standard error correction in two-stage estimation with nested samples. The Econometrics Journal, 6(2):401–407, December 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kluppelberg:2002:TRR

[KMVW02] Claudia Klüppelberg, Ross A. Maller, Mark Van De Vyver, and Derick Wee. Testing for reduction to random walk in autoregressive conditional heteroskedasticity models. The Econometrics Journal, 5(2):387–416, December 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Knight:2008:ESC

[KN08] John Knight and Cathy Q. Ning. Estimation of the stochastic conditional duration model via alternative methods. The Econometrics Journal, 11(3):593–616, November 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Keane:2020:CDN

[KN20] Michael Keane and Timothy Neal. Comparing deep neural network and econometric approaches to predicting the impact of climate change on agricultural yield. The Econometrics Journal, 23(3):S59–S80, September 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/23/3/S59/5849000.

Knaus:2022:DML

[Kna22] Michael C. Knaus. Double machine learning-based programme evaluation under unconfoundedness. The Econometrics Journal, 25(3):602–627, September 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/25/3/602/6596870.

Kurita:2011:CMP

[KNR11] Takamitsu Kurita, Heino Bohn Nielsen, and Anders Rahbek. An I(2) cointegration model with piecewise linear trends. The Econometrics Journal, 14(2):131–155, July 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Koenker:2020:IMR

[Koe20] Roger Koenker. The ignorant monopolist redux. The Econometrics Journal, 23(2):316–322, May 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.

oup.com/ectj/article/23/2/316/5625070.

Komarova:2013:PIA

[Kom13] Tatiana Komarova. Partial identification in asymmetric auctions in the absence of independence. The Econometrics Journal, 16 (1):S60–S92, February 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kongsted:2003:CAS

[Kon03] Hans Christian Kongsted. An I(2) cointegration analysis of small-country import price determination. The Econometrics Journal, 6(1):53–71, June 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Koop:2010:RFC

[Koo10] Gary Koop. A review of A First Course in Bayesian Statistical Methods. The Econometrics Journal, 13(1):B1–B5, February 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Korolev:2022:RFE

[Kor22] Ivan Korolev. On reduced form estimation of the effect of policy interventions on the COVID-19 pandemic. The Econometrics Journal, 25(3):762–780, September 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:/

/academic.oup.com/ectj/article/25/3/762/6555449.

Kiviet:1998:DFA

[KP98] Jan F. Kiviet and Garry D. A. Phillips. Degrees of freedom adjustment for disturbance variance estimators in dynamic regression models. The Econometrics Journal, 1(2):44–70, December 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Koop:2001:TOJ

[KP01a] Gary Koop and Dale J. Poirier. Testing for optimality in job search models. The Econometrics Journal, 4(2):257–272, December 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Koop:2001:AFN

[KP01b] Gary Koop and Simon M. Potter. Are apparent findings of nonlinearity due to structural instability in economic time series? The Econometrics Journal, 4(1):37–55, June 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Koop:2004:FDF

[KP04] Gary Koop and Simon Potter. Forecasting in dynamic factor models using Bayesian model averaging. The Econometrics Journal, 7(2):550–565, December 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kiviet:2005:MAL

[KP05] Jan F. Kiviet and Garry D. A. Phillips. Moment approximation for least-squares estimators in dynamic regression models with a unit root. The Econometrics Journal, 8(2):115–142, July 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kyriacou:2017:IIS

[KPR17] Maria Kyriacou, Peter C. B. Phillips, and Francesca Rossi. Indirect inference in spatial autoregression. The Econometrics Journal, 20(2):168–189, June 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Koning:2003:DCS

[KR03] Ruud H. Koning and Geert Ridder. Discrete choice and stochastic utility maximization. The Econometrics Journal, 6(1):1–27, June 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kokoszka:2013:PSI

[KR13] Piotr Kokoszka and Matthew Reimherr. Predictability of shapes of intraday price curves. The Econometrics Journal, 16(3):285– 308, October 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kreiss:2023:IRD

[KR23] Alexander Kreiss and Christoph Rothe. Inference in regression discontinuity designs with high-dimensional covariates. The Econometrics Journal, 26(2):105–123, May 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/26/2/105/6957254.

Kring:2009:MTG

[KRH^{+}09] Sebastian Kring, Svetlozar T. Rachev, Markus Höchstötter, Frank J. Fabozzi, and Michele Leonardo Bianchi. Multi-tail generalized elliptical distributions for asset returns. The Econometrics Journal, 12(2):272–291, July 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kristensen:2012:NPD

[Kri12] Dennis Kristensen. Non-parametric detection and estimation of structural change. The Econometrics Journal, 15(3):420–461, October 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Krief:2017:SLM

[Kri17] Jerome M. Krief. Semi-linear mode regression. The Econometrics Journal, 20(2):149–167, June 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kruiniger:2021:IAM

[Kru21] Hugo Kruiniger. Identification without assuming mean stationarity: quasi-maximum likelihood estimation of dynamic panel models with endogenous regressors. The Econometrics Journal, 24(3):417–441, September 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/24/3/417/6030027.

Kapetanios:2006:URT

[KS06] George Kapetanios and Yongcheol Shin. Unit root tests in threeregime SETAR models. The Econometrics Journal, 9(2):252–278, July 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Komunjer:2010:SPE

[KS10] Ivana Komunjer and Andres Santos. Semi-parametric estimation of non-separable models: a minimum distance from independence approach. The Econometrics Journal, 13(3):S28–S55, October 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Koopman:1999:SAM

[KSD99] Siem Jan Koopman, Neil Shephard, and Jurgen A. Doornik. Statistical algorithms for models in state space using SsfPack 2.2. The Econometrics Journal, 2(1):107–160, June 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kline:2018:ITE

[KT18] Brendan Kline and Elie Tamer. Identification of treatment effects with selective participation in a randomized trial. The Econometrics Journal, 21(3):332–353, October 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kasy:2023:MSB

[KT23] Maximilian Kasy and Alexander Teytelboym. Matching with semi-bandits. The Econometrics Journal, 26(1):45–66, January 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/26/1/
45/6717767.

Kuosmanen:2008:RTC

[Kuo08] Timo Kuosmanen. Representation theorem for convex nonparametric least squares. The Econometrics Journal, 11(2):308–325, July 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Karabiyik:2021:FUC

[KW21] Hande Karabiyik and Joakim Westerlund. Forecasting using cross-section average-augmented time series regressions. The Econometrics Journal, 24(2):315–333, May 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/24/2/315/5921167.

Kleiman-Weiner:2018:NPB

[KWTZ18] Max Kleiman-Weiner, Joshua B. Tenenbaum, and Penghui Zhou. Non-parametric Bayesian inference of strategies in repeated games. The Econometrics Journal, 21(3):298–315, October 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kurozumi:2015:CSB

[KY15] Eiji Kurozumi and Yohei Yamamoto. Confidence sets for the break date based on optimal tests. The Econometrics Journal, 18(3):412–435, October 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Kejriwal:2021:GFA

[KY21] Mohitosh Kejriwal and Xuewen Yu. Generalized forecast averaging in autoregressions with a near unit root. The Econometrics Journal, 24(1):83–102, January 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.

oup.com/ectj/article/24/1/83/5814315.

Kaplan:2023:CLI

[KZ23] David M. Kaplan and Wei Zhao. Comparing latent inequality with ordinal data. The Econometrics Journal, 26(2):189–214, May 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
26/2/189/6840227.

Larsson:1998:DAU

[Lar98] Rolf Larsson. Distribution approximation of unit root tests in autoregressive models. The Econometrics Journal, 1(2):10–26, December 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Lombardi:2008:IES

[LC08] Marco J. Lombardi and Giorgio Calzolari. Indirect estimation of α-stable distributions and processes. The Econometrics Journal, 11(1):193–208, March 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Li:2011:NPT

[LCG11] Degui Li, Jia Chen, and Jiti Gao. Non-parametric time-varying coefficient panel data models with fixed effects. The Econometrics Journal, 14(3):387–408, October 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Lechner:1999:NBE

[Lec99] Michael Lechner. Nonparametric bounds on employment and income effects of continuous vocational training in East Germany. The Econometrics Journal, 2(1):1–28, June 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Li:2008:KNN

[LG08] Rui Li and Guan Gong. K-nearest-neighbour non-parametric estimation of regression functions in the presence of irrelevant variables. The Econometrics Journal, 11(2):396–408, July 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Lee:2021:EDM

[LG21] Sanghyeok Lee and Tue Gørgens. Estimation of dynamic models of recurrent events with censored data. The Econometrics Journal, 24(2):199–224, May 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/24/2/199/5903267.

Lee:2015:RHT

[LHK15] Wei-Ming Lee, Yu-Chin Hsu, and Chung-Ming Kuan. Robust hypothesis tests for M-estimators with possibly non-differentiable estimating functions. The Econometrics Journal, 18(1):95–116, February 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Lind:2005:RSK

[Lin05] Jo Thori Lind. Repeated surveys and the Kalman filter. The Econometrics Journal, 8(3):418–427, November 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Liu:2009:SFG

[Liu09] Ji-Chun Liu. Stationarity of a family of GARCH processes. The Econometrics Journal, 12(3):436–446, November 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Lee:2005:SEO

[LK05] Myoung-Jae Lee and Ayal Kimhi. Simultaneous equations in ordered discrete responses with regressor-dependent thresholds. The Econometrics Journal, 8(2):176–196, July 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Liu:2016:MAP

[LK16] Chu-An Liu and Biing-Shen Kuo. Model averaging in predictive regressions. The Econometrics Journal, 19(2):203–231, June 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Leybourne:2003:TCP

[LKSN03] Stephen Leybourne, Tae-Hwan Kim, Vanessa Smith, and Paul Newbold. Tests for a change in persistence against the null of difference-stationarity. The Econometrics Journal, 6(2):291–311, December 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Li:2023:APM

[LL23] Chaojun Li and Yan Liu. Asymptotic properties of the maximum likelihood estimator in regime-switching models with timevarying transition probabilities. The Econometrics Journal, 26 (1):67–87, January 2023. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/26/1/67/6671600.

Larsson:2001:LBC

[LLL01] Rolf Larsson, Johan Lyhagen, and Mickael Löthgren. Likelihoodbased cointegration tests in heterogeneous panels. The Econometrics Journal, 4(1):109–142, June 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). See corrigendum [ÖD11].

Liu:2022:EGM

[LM22] Hang Liu and Kanchan Mukherjee. R-estimators in GARCH models: asymptotics and applications. The Econometrics Journal, 25(1):98–113, January 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/25/1/98/6359714.

Lutkepohl:2021:TIH

[LMNS21] Helmut Lütkepohl, Mika Meitz, Aleksei Netsunajev, and Pentti Saikkonen. Testing identification via heteroskedasticity in structural vector autoregressive models. The Econometrics Journal, 24(1):1–22, January 2021. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/24/1/1/5820226. See erratum [Ano21b].

Lee:2014:BSB

[LMP14] Young K. Lee, Enno Mammen, and Byeong U. Park. Backfitting and smooth backfitting in varying coefficient quantile regression. The Econometrics Journal, 17(2):S20–S38, June 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Liu:2020:TWE

[LMW20] Shenglong Liu, Ismael Mourifié, and Yuanyuan Wan. Two-way exclusion restrictions in models with heterogeneous treatment effects. The Econometrics Journal, 23(3):345–362, September 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/23/3/
345/5849432.

Leybourne:1999:BDF

[LN99] Stephen J. Leybourne and Paul Newbold. The behaviour of Dickey–Fuller and Phillips-Perron tests under the alternative hypothesis. The Econometrics Journal, 2(1):92–100, June 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Leybourne:2000:BSS

[LN00] Stephen J. Leybourne and Paul Newbold. Behaviour of the standard and symmetric Dickey–Fuller-type tests when there is a break under the null hypothesis. The Econometrics Journal, 3 (1):1–15, June 2000. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Linton:2009:NPR

[LNN09] Oliver Linton, Jens Perch Nielsen, and Søren Feodor Nielsen. Non-parametric regression with a latent time series. The Econometrics Journal, 12(2):187–207, July 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Liu:2013:HRM

[LO13] Qingfeng Liu and Ryo Okui. Heteroscedasticity-robust C_{p} model averaging. The Econometrics Journal, 16(3):463–472, October 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Lieberman:2005:EAM

[LP05] Offer Lieberman and Peter C. B. Phillips. Expansions for approximate maximum likelihood estimators of the fractional difference parameter. The Econometrics Journal, 8(3):367–379, November 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Li:2009:DNF

[LP09] Qiaoling Li and Jiazhu Pan. Determining the number of factors in a multivariate error correction-volatility factor model. The Econometrics Journal, 12(1):45–61, March 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Liu:2017:PEB

[LPR17] Xiaodong Liu, Eleonora Patacchini, and Edoardo Rainone. Peer effects in bedtime decisions among adolescents: a social network model with sampled data. The Econometrics Journal, 20(3): S103–S125, October 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Linton:2014:TSD

[LPW14] Oliver Linton, Thierry Post, and Yoon-Jae Whang. Testing for the stochastic dominance efficiency of a given portfolio. The Econometrics Journal, 17(2):S59–S74, June 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Lucchetti:2005:ART

[LR05] Riccardo Lucchetti and Eduardo Rossi. Artificial regression testing in the GARCH-in-mean model. The Econometrics Journal, 8(3):306–322, November 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Liu:2006:URT

[LR06] Hui Liu and Gabriel Rodríguez. Unit root tests and structural change when the initial observation is drawn from its unconditional distribution. The Econometrics Journal, 9(2):225–251, July 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Lanne:2005:NLG

[LS05] Markku Lanne and Pentti Saikkonen. Non-linear GARCH models for highly persistent volatility. The Econometrics Journal, 8(2): 251–276, July 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Linton:2012:E

[LS12] Oliver Linton and Richard J. Smith. EDITORIAL. The Econometrics Journal, 15(1):Ci–Cii, February 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Lee:2021:CSA

[LS21] Seojeong Lee and Youngki Shin. Complete subset averaging with many instruments. The Econometrics Journal, 24(2):290–314, May 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
24/2/290/5981609.

Luutkepohl:2001:MEV

[LST01] Helmut Lüutkepohl, Pentti Saikkonen, and Carsten Trenkler. Maximum eigenvalue versus trace tests for the cointegrating rank of a VAR process. The Econometrics Journal, 4(2):287–310, December 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Lobato:2006:OFD

[LV06] Ignacio N. Lobato and Carlos Velasco. Optimal fractional Dickey–Fuller tests. The Econometrics Journal, 9(3):492–510, November 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Lobato:2022:SSE

[LV22] Ignacio N. Lobato and Carlos Velasco. Single step estimation of ARMA roots for nonfundamental nonstationary fractional models. The Econometrics Journal, 25(2):455–476, May 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/25/2/
455/6505131.

Lin:2017:STS

[LW17] Juan Lin and Ximing Wu. A sequential test for the specification of predictive densities. The Econometrics Journal, 20(2):190–220, June 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Li:2012:WII

[LX12] Hong Li and Zhijie Xiao. Weak instrument inference in the presence of parameter instability. The Econometrics Journal, 15(3): 395–419, October 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Lin:2017:ESI

[LX17] Zhongjian Lin and Haiqing Xu. Estimation of social-influence-dependent peer pressure in a large network game. The Econometrics Journal, 20(3):S86–S102, October 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Liu:2020:AFT

[LY20] Ruixuan Liu and Zhengfei Yu. Accelerated failure time models with log-concave errors. The Econometrics Journal, 23(2):251– 268, May 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
23/2/251/5663610.

Liu:2021:DDM

[LZZ21] Molei Liu, Yi Zhang, and Doudou Zhou. Double/debiased machine learning for logistic partially linear model. The Econometrics Journal, 24(3):559–588, September 2021. CODEN ????

ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/24/3/559/6296639.

Ma:2017:SOR

[Ma17] Jun Ma. Second-order refinement of empirical likelihood ratio tests of nonlinear restrictions. The Econometrics Journal, 20(1): 139–148, February 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Madsen:2005:ECR

[Mad05] Edith Madsen. Estimating cointegrating relations from a cross section. The Econometrics Journal, 8(3):380–405, November 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Madsen:2010:URI

[Mad10] Edith Madsen. Unit root inference in panel data models where the time-series dimension is fixed: a comparison of different tests. The Econometrics Journal, 13(1):63–94, February 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Magnus:2002:EMU

[Mag02] Jan Magnus. Estimation of the mean of a univariate normal distribution with known variance. The Econometrics Journal, 5 (1):225–236, June 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Magdalinos:2007:IUE

[Mag07] Tassos Magdalinos. On the inconsistency of the unrestricted estimator of the information matrix near a unit root. The Econometrics Journal, 10(2):245–262, July 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Magnusson:2010:ILD

[Mag10] Leandro M. Magnusson. Inference in limited dependent variable models robust to weak identification. The Econometrics Journal, 13(3):S56–S79, October 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Manski:2013:ITR

[Man13] Charles F. Manski. Identification of treatment response with social interactions. The Econometrics Journal, 16(1):S1–S23, February 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Mao:2016:TEC

[Mao16] Guangyu Mao. Testing for error cross-sectional independence using pairwise augmented regressions. The Econometrics Journal, 19(3):237–260, October 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Martin:2012:ROH

[Mar12] Gael Martin. A review of The Oxford Handbook of Bayesian Econometrics edited by Geweke (John), Koop (Gary) and van Dijk (Herman). The Econometrics Journal, 15(3):B11–B15, October 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Marsh:2013:RNP

[Mar13] Patrick Marsh. A review of non-parametric econometrics. The Econometrics Journal, 16(2):B1–B3, June 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Mayoral:2007:MDE

[May07] Laura Mayoral. Minimum distance estimation of stationary and non-stationary ARFIMA processes. The Econometrics Journal, 10(1):124–148, March 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

McElroy:2008:EFH

[McE08] Tucker McElroy. Exact formulas for the Hodrick–Prescott filter. The Econometrics Journal, 11(1):209–217, March 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

McKenzie:2001:EAM

[McK01] David J. McKenzie. Estimation of AR(1) models with unequally spaced pseudo-panels. The Econometrics Journal, 4(1):89–108, June 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Menchetti:2023:CCO

[MCM23] Fiammetta Menchetti, Fabrizio Cipollini, and Fabrizia Mealli. Combining counterfactual outcomes and ARIMA models for policy evaluation. The Econometrics Journal, 26(1):1–24, January 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/26/1/
1/6713620.

Meddahi:2003:ARI

[Med03] Nour Meddahi. ARMA representation of integrated and realized variances. The Econometrics Journal, 6(2):335–356, December 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Meyer:2003:SVB

[MFB03] Renate Meyer, David A. Fournier, and Andreas Berg. Stochastic volatility: Bayesian computation using automatic differentiation and the extended Kalman filter. The Econometrics Journal, 6(2): 408–420, December 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

McElroy:2019:TCV

[MJ19] Tucker S. McElroy and Agnieszka Jach. Testing collinearity of vector time series. The Econometrics Journal, 22(2):97–116, May 2019. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/22/2/
97/5303850.

Moench:2011:HFA

[MN11] Emanuel Moench and Serena Ng. A hierarchical factor analysis of U.S. housing market dynamics. The Econometrics Journal, 14(1):C1–C24, February 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Marriott:2003:EET

[MNT03] J. M. Marriott, J. C. Naylor, and A. R. Tremayne. Exploring economic time series: a Bayesian graphical approach. The Econometrics Journal, 6(1):124–145, June 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Moffatt:2001:GCM

[Mof01] Peter G. Moffatt. Graphical conditional moment tests. The Econometrics Journal, 4(1):56–69, June 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Monfardini:1998:ESV

[Mon98] Chiara Monfardini. Estimating stochastic volatility models through indirect inference. The Econometrics Journal, 1(1):113– 128, June 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Machado:2005:BEC

[MP05] José A. F. Machado and Paulo Parente. Bootstrap estimation of covariance matrices via the percentile method. The Econometrics Journal, 8(1):70–78, March 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Moon:2008:ALP

[MP08] Hyungsik Roger Moon and Benoit Perron. Asymptotic local power of pooled t-ratio tests for unit roots in panels with fixed effects. The Econometrics Journal, 11(1):80–104, March 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Mutl:2011:HTC

[MP11] Jan Mutl and Michael Pfaffermayr. The Hausman test in a Cliff and Ord panel model. The Econometrics Journal, 14(1):48–76, February 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Masten:2023:CEA

[MP23] Matthew A. Masten and Alexandre Poirier. Choosing exogeneity assumptions in potential outcome models. The Econometrics Journal, 26(3):327–349, September 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:/

/academic.oup.com/ectj/article/26/3/327/7005213.

Moon:2014:POP

[MPP14] Hyungsik Roger Moon, Benoit Perron, and Peter C. B. Phillips. Point-optimal panel unit root tests with serially correlated errors. The Econometrics Journal, 17(3):338–372, October 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Manrique:1998:SBL

[MS98] Aurora Manrique and Neil Shephard. Simulation-based likelihood inference for limited dependent processes. The Econometrics Journal, 1(1):174–202, June 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Magnus:2005:TE

[MS05a] Jan R. Magnus and Ashoke K. Sinha. On Theil’s errors. The Econometrics Journal, 8(1):39–54, March 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Moauro:2005:TDU

[MS05b] Filippo Moauro and Giovanni Savio. Temporal disaggregation using multivariate structural time series models. The Econometrics Journal, 8(2):214–234, July 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Munkin:1999:SML

[MT99] Murat K. Munkin and Pravin K. Trivedi. Simulated maximum likelihood estimation of multivariate mixed-Poisson regression models, with application. The Econometrics Journal, 2(1):29– 48, June 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Moscone:2017:SEH

[MTV17] Francesco Moscone, Elisa Tosetti, and Veronica Vinciotti. Sparse estimation of huge networks with a block-wise structure. The Econometrics Journal, 20(3):S61–S85, October 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Magnus:2007:LSD

[MV07] Jan R. Magnus and Andrey L. Vasnev. Local sensitivity and diagnostic tests. The Econometrics Journal, 10(1):166–192, March 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Maasoumi:2008:ERG

[MW08] Esfandiar Maasoumi and Le Wang. Economic reform, growth and convergence in China. The Econometrics Journal, 11(1):128–154, March 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

MacKinnon:2018:WBF

[MW18] James G. MacKinnon and Matthew D. Webb. The wild bootstrap for few (treated) clusters. The Econometrics Journal, 21(2):114– 135, June 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Margaritella:2023:UIC

[MW23] Luca Margaritella and Joakim Westerlund. Using information criteria to select averages in CCE. The Econometrics Journal, 26(3):405–421, September 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/26/3/405/7100953.

Meyer:2000:BBA

[MY00] Renate Meyer and Jun Yu. BUGS for a Bayesian analysis of stochastic volatility models. The Econometrics Journal, 3(2): 198–215, December 2000. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Mynbaev:2011:RAC

[Myn11] Kairat T. Mynbaev. Regressions with asymptotically collinear regressors. The Econometrics Journal, 14(2):304–320, July 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Mu:2018:IEH

[MZ18] Beili Mu and Zhengyu Zhang. Identification and estimation of heteroscedastic binary choice models with endogenous dummy regressors. The Econometrics Journal, 21(2):218–246, June 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Nielsen:2011:FMN

[NF11] Morten Ørregaard Nielsen and Per Frederiksen. Fully modified narrow-band least squares estimation of weak fractional cointegration. The Econometrics Journal, 14(1):77–120, February 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Nicolau:2002:NTS

[Nic02] João Nicolau. A new technique for simulating the likelihood of stochastic differential equations. The Econometrics Journal, 5 (1):91–103, June 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Nielsen:2004:CAP

[Nie04a] Heino Bohn Nielsen. Cointegration analysis in the presence of outliers. The Econometrics Journal, 7(1):249–271, June 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Nielsen:2004:EIM

[Nie04b] Morten Ørregaard Nielsen. Efficient inference in multivariate fractionally integrated time series models. The Econometrics Journal, 7(1):63–97, June 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Nielsen:2008:IOC

[Nie08] Heino Bohn Nielsen. Influential observations in cointegrated VAR models: Danish money demand 1973-2003. The Econometrics Journal, 11(1):39–57, March 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Nielsen:2024:PQL

[NR24] Heino Bohn Nielsen and Anders Rahbek. Penalized quasilikelihood estimation and model selection with parameters on the boundary of the parameter space. The Econometrics Journal, 27(1):107–125, January 2024. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/27/1/107/7287596.

Newbold:2001:FEE

[NS01] Paul Newbold and Richard J. Smith. Forecasting in econometrics: Editors’ introduction. The Econometrics Journal, 4(1):1–2, June 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Nankervis:2012:TUE

[NS12] John C. Nankervis and Nathan E. Savin. Testing for uncorrelated errors in ARMA models: non-standard Andrews–Ploberger tests. The Econometrics Journal, 15(3):516–534, October 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Ng:2024:CHF

[NS24] Serena Ng and Susannah Scanlan. Constructing high frequency economic indicators by imputation. The Econometrics Journal, 27(1):C1–C30, January 2024. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/27/1/C1/7424127.

Nekipelov:2022:ROM

[NSS22] Denis Nekipelov, Vira Semenova, and Vasilis Syrgkanis. Regularised orthogonal machine learning for nonlinear semiparametric models. The Econometrics Journal, 25(1):233–255, January 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/25/1/
233/6327541.

Nakatani:2009:TVI

[NT09] Tomoaki Nakatani and Timo Teräsvirta. Testing for volatility interactions in the constant conditional correlation GARCH model. The Econometrics Journal, 12(1):147–163, March 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Ng:2002:FAT

[NV02] Serena Ng and Timothy Vogelsang. Forecasting autoregressive time series in the presence of deterministic components. The Econometrics Journal, 5(1):196–224, June 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Orsal:2011:CLB

[ÖD11] Deniz Dilan Karaman Örsal and Bernd Droge. Corrigendum to ‘Likelihood-based cointegration tests in heterogeneous panels’ (Larsson R., J. Lyhagen and M. Löthgren, Econometrics Journal, 4, 2001, 109–142). The Econometrics Journal, 14(1):121– 125, February 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). See [LLL01].

Ouyang:2006:CVN

[OLL06] Desheng Ouyang, Dong Li, and Qi Li. Cross-validation and nonparametric k nearest-neighbour estimation. The Econometrics Journal, 9(3):448–471, November 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Olden:2022:TDE

[OM22] Andreas Olden and Jarle Møen. The triple difference estimator. The Econometrics Journal, 25(3):531–553, September 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/25/3/
531/6545797.

Ooms:1999:RSS

[Oom99] Marius Ooms. Review of SsfPack 2.2: statistical algorithms for models in state space. The Econometrics Journal, 2(1):161–166, June 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Otsu:2023:EMD

[OP23] Taisuke Otsu and Martin Pesendorfer. Equilibrium multiplicity in dynamic games: Testing and estimation. The Econometrics Journal, 26(1):C26–C42, January 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.

oup.com/ectj/article/26/1/C26/6576205.

Osborn:2012:RMN

[Osb12] Denise R. Osborn. A review of Modelling Nonlinear Economic Time Series by Teräsvirta (Timo), Tjøstheim (Dag) and Granger (Clive W. J.). The Econometrics Journal, 15(2):B1–B3, June 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Ohn:2004:TDD

[OTP04] Jonathan Ohn, Larry W. Taylor, and Adrian Pagan. Testing for duration dependence in economic cycles. The Econometrics Journal, 7(2):528–549, December 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Otsu:2011:LDG

[Ots11] Taisuke Otsu. Large deviations of generalized method of moments and empirical likelihood estimators. The Econometrics Journal, 14(2):321–329, July 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Orme:2006:ADF

[OY06] Chris D. Orme and Takashi Yamagata. The asymptotic distribution of the F -test statistic for individual effects. The Econometrics Journal, 9(3):404–422, November 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Okui:2020:KEP

[OY20] Ryo Okui and Takahide Yanagi. Kernel estimation for panel data with heterogeneous dynamics. The Econometrics Journal, 23(1): 156–175, January 2020. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/23/1/156/5607791.

Poignard:2023:EHD

[PA23] Benjamin Poignard and Manabu Asai. Estimation of highdimensional vector autoregression via sparse precision matrix.

The Econometrics Journal, 26(2):307–326, May 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/26/2/307/6985002.

Palandri:2022:RIC

[Pal22] Alessandro Palandri. Rank-invariance conditions for the comparison of volatility forecasts. The Econometrics Journal, 25(1): 155–175, January 2022. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/25/1/155/6228830.

Paruolo:2002:MCE

[Par02] Paolo Paruolo. On Monte Carlo estimation of relative power. The Econometrics Journal, 5(1):65–75, June 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Petursson:2000:RHD

[Pét00] Thórarinn G. Pétursson. The representative household’s demand for money in a cointegrated VAR model. The Econometrics Journal, 3(2):162–176, December 2000. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Phillips:2015:EMT

[Phi15] Peter C. B. Phillips. Edmond Malinvaud: a tribute to his contributions in econometrics. The Econometrics Journal, 18(2): A1–A13, June 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Pitarakis:2004:LSE

[Pit04] Jean-Yves Pitarakis. Least squares estimation and tests of breaks in mean and variance under misspecification. The Econometrics Journal, 7(1):32–54, June 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Paparoditis:2002:TBB

[PP02] Efstathios Paparoditis and Dimitris Politis. The tapered block bootstrap for general statistics from stationary sequences. The Econometrics Journal, 5(1):131–148, June 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Panopoulou:2004:CAD

[PP04] Ekaterini Panopoulou and Nikitas Pittis. A comparison of autoregressive distributed lag and dynamic OLS cointegration estimators in the case of a serially correlated cointegration error. The Econometrics Journal, 7(2):585–617, December 2004. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Pedersen:2014:MVT

[PR14] Rasmus S. Pedersen and Anders Rahbek. Multivariate variance targeting in the BEKK–GARCH model. The Econometrics Journal, 17(1):24–55, February 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Perron:2016:RBT

[PR16] Pierre Perron and Gabriel Rodríguez. Residuals-based tests for cointegration with generalized least-squares detrended data. The Econometrics Journal, 19(1):84–111, February 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Proietti:1998:SPA

[Pro98] Tommaso Proietti. Spurious periodic autoregressions. The Econometrics Journal, 1(1):1–22, June 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Proietti:2006:TDS

[Pro06] Tommaso Proietti. Temporal disaggregation by state space methods: Dynamic regression methods revisited. The Econometrics Journal, 9(3):357–372, November 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Petursson:2001:WFE

[PS01] Thórarinn G. Pétursson and Torsten Sløk. Wage formation and employment in a cointegrated VAR model. The Econometrics Journal, 4(2):191–209, December 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Phillips:2003:DPE

[PS03] Peter C. B. Phillips and Donggyu Sul. Dynamic panel estimation and homogeneity testing under cross section dependence. The Econometrics Journal, 6(1):217–259, June 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Preminger:2007:MSM

[PS07] Arie Preminger and Shinichi Sakata. A model selection method for S-estimation. The Econometrics Journal, 10(2):294–319, July 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Poskitt:2009:AMC

[PS09] D. S. Poskitt and C. L. Skeels. Assessing the magnitude of the concentration parameter in a simultaneous equations model. The Econometrics Journal, 12(1):26–44, March 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Perron:2011:RES

[PS11a] Pierre Perron and Richard J. Smith. Royal economic society annual conference 2009 special issue on factor models: Theoretical and applied perspectives. The Econometrics Journal, 14(1): ci–ciii, February 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Phillips:2011:NPR

[PS11b] Peter C. B. Phillips and Liangjun Su. Non-parametric regression under location shifts. The Econometrics Journal, 14(3):457–486, October 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Patton:2016:RES

[PS16] Andrew J. Patton and Richard J. Smith. Royal economic society annual conference 2014 special issue on large dimensional models. The Econometrics Journal, 19(1):Ci–Cii, February 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Preminger:2017:LSE

[PS17] Arie Preminger and Giuseppe Storti. Least-squares estimation of GARCH(1,1) models with heavy-tailed errors. The Econometrics Journal, 20(2):221–258, June 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Psaradakis:2001:MLS

[Psa01] Zacharias Psaradakis. Markov level shifts and the unit-root hypothesis. The Econometrics Journal, 4(2):225–241, December 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Pong:2008:DSL

[PST08] Shiuyan Pong, Mark B. Shackleton, and Stephen J. Taylor. Distinguishing short and long memory volatility specifications. The Econometrics Journal, 11(3):617–637, November 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Ponomareva:2011:MMI

[PT11] Maria Ponomareva and Elie Tamer. Misspecification in moment inequality models: back to moment equalities? The Econometrics Journal, 14(2):186–203, July 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Pesaran:2008:BAL

[PUY08] M. Hashem Pesaran, Aman Ullah, and Takashi Yamagata. A bias-adjusted LM test of error cross-section independence. The Econometrics Journal, 11(1):105–127, March 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Perron:2001:AAN

[PV01] Pierre Perron and Cosme Vodounou. Asymptotic approximations in the near-integrated model with a non-zero initial condition. The Econometrics Journal, 4(1):143–169, June 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Phillips:2011:CGA

[PY11] Peter C. B. Phillips and Jun Yu. Corrigendum to ‘A Gaussian approach for continuous time models of short-term interest rates’ (Yu, J. and P. C. B. Phillips, Econometrics Journal, 4, 210– 224). The Econometrics Journal, 14(1):126–129, February 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). See [YP01].

Qu:2013:SVM

[QP13] Zhongjun Qu and Pierre Perron. A stochastic volatility model with random level shifts and its applications to S&P 500 and NASDAQ return indices. The Econometrics Journal, 16(3):309– 339, October 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Qu:2007:SCD

[Qu07] Zhongjun Qu. Searching for cointegration in a dynamic system. The Econometrics Journal, 10(3):580–604, November 2007. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Qu:2016:IVE

[QWfL16] Xi Qu, Xiaoliang Wang, and Lung fei Lee. Instrumental variable estimation of a spatial dynamic panel model with endogenous spatial weights when T is small. The Econometrics Journal, 19(3):261–290, October 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Raggi:2005:AMM

[Rag05] Davide Raggi. Adaptive MCMC methods for inference on affine stochastic volatility models with jumps. The Econometrics Journal, 8(2):235–250, July 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Read:2022:AIS

[Rea22] Matthew Read. Algorithms for inference in SVARs identified with sign and zero restrictions. The Econometrics Journal, 25(3):699– 718, September 2022. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/25/3/699/6529229.

Robinson:2006:IVE

[RG06] P. M. Robinson and M. Gerolimetto. Instrumental variables estimation of stationary and non-stationary cointegrating regressions. The Econometrics Journal, 9(2):291–306, July 2006. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Rahbek:1999:CRI

[RM99] Anders Rahbek and Rocco Mosconi. Cointegration rank inference with stationary regressors in VAR models. The Econometrics Journal, 2(1):76–91, June 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Robin:2010:RDS

[Rob10] Jean-Marc Robin. Recent developments in structural microeconometrics. The Econometrics Journal, 13(3):Si–Sii, October 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Rodrigues:2002:LTT

[Rod02] Paulo Rodrigues. On LM type tests for seasonal unit roots in quarterly data. The Econometrics Journal, 5(1):176–195, June 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Rossi:2007:EHT

[Ros07] Barbara Rossi. Expectations hypotheses tests at long horizons. The Econometrics Journal, 10(3):554–579, November 2007. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Rossi:2016:REF

[Ros16] Barbara Rossi. A review of economic forecasting. The Econometrics Journal, 19(3):B1–B3, October 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Rose:2017:IPE

[Ros17] Christiern D. Rose. Identification of peer effects through social networks using variance restrictions. The Econometrics Journal, 20(3):S47–S60, October 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Rossi:2021:IEE

[Ros21] Barbara Rossi. Identifying and estimating the effects of unconventional monetary policy: How to do it and what have we learned? The Econometrics Journal, 24(1):C1–C32, January 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/24/1/
C1/5872478.

Rodriguez-Poo:2014:DSP

[RPS14] Juan M. Rodriguez-Poo and Alexandra Soberon. Direct semiparametric estimation of fixed effects panel data varying coefficient models. The Econometrics Journal, 17(1):107–138, February 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Robinson:2014:ILM

[RR14] Peter M. Robinson and Francesca Rossi. Improved Lagrange multiplier tests in spatial autoregressions. The Econometrics Journal, 17(1):139–164, February 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Runstler:2004:MPS

[Rün04] Gerhard Rünstler. Modelling phase shifts among stochastic cycles. The Econometrics Journal, 7(1):232–248, June 2004. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Rivers:2002:MST

[RV02] Douglas Rivers and Quang Vuong. Model selection tests for nonlinear dynamic models. The Econometrics Journal, 5(1):1–39, June 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Romeu:2005:CEB

[RVH05] Andrés Romeu and Marcos Vera-Hernández. Counts with an endogenous binary regressor: a series expansion approach. The Econometrics Journal, 8(1):1–22, March 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Richter:2023:TPC

[RWW23] Stefan Richter, Weining Wang, and Wei Biao Wu. Testing for parameter change epochs in GARCH time series. The Econometrics Journal, 26(3):467–491, September 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/26/3/467/7022314.

Sandberg:2008:CVL

[San08] Rickard Sandberg. Critical values for linearity tests in timevarying smooth transition autoregressive models when data are highly persistent. The Econometrics Journal, 11(3):638–647, November 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Semenova:2021:DML

[SC21] Vira Semenova and Victor Chernozhukov. Debiased machine learning of conditional average treatment effects and other causal functions. The Econometrics Journal, 24(2):264–289, May 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/24/2/
264/5899048.

Scaillet:2016:IPN

[Sca16] Olivier Scaillet. On ill-posedness of nonparametric instrumental variable regression with convexity constraints. The Econometrics Journal, 19(2):232–236, June 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Schluter:2012:DGD

[Sch12a] Christian Schluter. Discussion of S. G. Donald et al. and R. Davidson. The Econometrics Journal, 15(1):C54–C57, February 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). See [DHB12, Dav12].

Schluter:2012:PII

[Sch12b] Christian Schluter. On the problem of inference for inequality measures for heavy-tailed distributions. The Econometrics Journal, 15(1):125–153, February 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Choi:2010:IRM

[sCK10] Hwan sik Choi and Nicholas M. Kiefer. Improving robust model selection tests for dynamic models. The Econometrics Journal, 13(2):177–204, July 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Sentana:1998:RBC

[Sen98] Enrique Sentana. The relation between conditionally heteroskedastic factor models and factor GARCH models. The Econometrics Journal, 1(2):1–9, December 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Sen:2003:LBD

[Sen03] A. Sen. Limiting behaviour of Dickey–Fuller F -tests under the crash model alternative. The Econometrics Journal, 6(2):421– 429, December 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Sentana:2009:EMV

[Sen09] Enrique Sentana. The econometrics of mean-variance efficiency tests: a survey. The Econometrics Journal, 12(3):C65–C101, November 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Sherman:1999:CCM

[SHD99] Robert P. Sherman, Yu-Yun K. Ho, and Siddhartha R. Dalal. Conditions for convergence of Monte Carlo EM sequences with an application to product diffusion modeling. The Econometrics Journal, 2(2):248–267, December 1999. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Shin:2008:SEB

[Shi08] Youngki Shin. Semiparametric estimation of the Box–Cox transformation model. The Econometrics Journal, 11(3):517–537, November 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Song:2014:GDS

[SHR14] Song Song, Wolfgang K. Härdle, and Ya’acov Ritov. Generalized dynamic semi-parametric factor models for high-dimensional non-stationary time series. The Econometrics Journal, 17(2): S101–S131, June 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Silva:2003:NEM

[Sil03] J. M. C. Santos Silva. A note on the estimation of mixture models under endogenous sampling. The Econometrics Journal, 6(1):46– 52, June 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Silva:2011:RME

[Sil11] João M. C. Santos Silva. A review of Micro-Econometrics: Methods of Moments and Limited Dependent Variables (2nd Ed.) by Lee (Myoung-jae). The Econometrics Journal, 14(2):B1–B4, July 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Smith:2003:MSS

[Smi03] Murray D. Smith. Modelling sample selection using Archimedean copulas. The Econometrics Journal, 6(1):99–123, June 2003. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Smith:2008:SFM

[Smi08a] Murray D. Smith. Stochastic frontier models with dependent error components. The Econometrics Journal, 11(1):172–192, March 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Smith:2008:EJR

[Smi08b] Richard J. Smith. The econometrics journal of the Royal economic society. The Econometrics Journal, 11(1):i–iii, March 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Smith:2015:RES

[Smi15] Richard J. Smith. Royal economic society annual conference 2012 special issue on econometrics of forecasting. The Econometrics Journal, 18(2):Ci–Cii, June 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Smith:2016:RES

[Smi16] Richard J. Smith. Royal economic society annual conference 2013special issue on econometrics of heterogeneity. The Econometrics Journal, 19(3):Ciii–Civ, October 2016. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Smith:2017:RES

[Smi17] Richard J. Smith. Royal economic society annual conference 2015 special issue on econometrics of matching. The Econometrics Journal, 20(2):Ci–Cii, June 2017. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Smith:2018:RES

[Smi18] Richard J. Smith. Royal economic society annual conference 2016 special issue on model selection and inference. The Econometrics Journal, 21(1):Ci–Cii, February 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Sorensen:2000:PBE

[Sør00] Michael Sørensen. Prediction-based estimating functions. The Econometrics Journal, 3(2):123–147, December 2000. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Sperlich:2009:NNP

[Spe09] Stefan Sperlich. A note on non-parametric estimation with predicted variables. The Econometrics Journal, 12(2):382–395, July 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Sarafidis:2009:IEC

[SR09] Vasilis Sarafidis and Donald Robertson. On the impact of error cross-sectional dependence in short dynamic panel estimation. The Econometrics Journal, 12(1):62–81, March 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Soberon:2022:NPD

[SRPR22] Alexandra Soberon, Juan M. Rodriguez-Poo, and Peter M. Robinson. Nonparametric panel data regression with parametric cross-sectional dependence. The Econometrics Journal, 25(1): 114–133, January 2022. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/25/1/114/6272425.

Shin:2006:MGT

[SS06] Yongcheol Shin and Andy Snell. Mean group tests for stationarity in heterogeneous panels. The Econometrics Journal, 9(1):123– 158, March 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Singh:2024:DRC

[SS24] Rahul Singh and Liyang Sun. Double robustness for complier parameters and a semi-parametric test for complier characteristics. The Econometrics Journal, 27(1):1–20, January 2024. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL

http://academic.oup.com/ectj/article/27/1/1/7303282.

Strikholm:2006:SPD

[ST06] Birgit Strikholm and Timo Teräsvirta. A sequential procedure for determining the number of regimes in a threshold autoregressive model. The Econometrics Journal, 9(3):472–491, November 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Shin:2021:ECM

[ST21] Youngki Shin and Zvezdomir Todorov. Exact computation of maximum rank correlation estimator. The Econometrics Journal, 24(3):589–607, September 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/24/3/589/6247621.

Stoye:2022:BIP

[Sto22] Jörg Stoye. Bounding infection prevalence by bounding selectivity and accuracy of tests: with application to early COVID-19.

The Econometrics Journal, 25(1):1–14, January 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/25/1/1/6325165.

Sun:2013:HAR

[Sun13] Yixiao Sun. A heteroskedasticity and autocorrelation robust F test using an orthonormal series variance estimator. The Econometrics Journal, 16(1):1–26, February 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Song:2022:EIT

[SY22] Kyungchul Song and Zhengfei Yu. Estimation and inference on treatment effects under treatment-based sampling designs. The Econometrics Journal, 25(3):554–575, September 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/25/3/554/6554242.

Schafgans:2010:SAA

[SZW10] Marcia M. A. Schafgans and Victoria Zinde-Walsh. Smoothness adaptive average derivative estimation. The Econometrics Journal, 13(1):40–62, February 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Takada:2008:AQP

[Tak08] Teruko Takada. Asymptotic and qualitative performance of nonparametric density estimators: a comparative study. The Econometrics Journal, 11(3):573–592, November 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Taylor:2013:RUR

[Tay13] Robert Taylor. A review of unit root tests in time series: Volumes 1 and 2. The Econometrics Journal, 16(3):B5–B8, October 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Tchatoka:2015:BVS

[Tch15] Firmin Doko Tchatoka. On bootstrap validity for specification tests with weak instruments. The Econometrics Journal, 18(1): 137–146, February 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Tchatoka:2014:IRI

[TD14] Firmin Doko Tchatoka and Jean-Marie Dufour. Identificationrobust inference for endogeneity parameters in linear structural models. The Econometrics Journal, 17(1):165–187, February 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Tao:2014:SIM

[TfL14] Ji Tao and Lung fei Lee. A social interaction model with an extreme order statistic. The Econometrics Journal, 17(3):197– 240, October 2014. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Thomas:2006:CEB

[Tho06] Alban Thomas. Consistent estimation of binary-choice panel data models with heterogeneous linear trends. The Econometrics Journal, 9(2):177–195, July 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Tsionas:2004:MSS

[TK04] Efthymios G. Tsionas and Subal C. Kumbhakar. Markov switching stochastic frontier model. The Econometrics Journal, 7(2): 398–425, December 2004. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Tse:2002:RBD

[Tse02] Y. K. Tse. Residual-based diagnostics for conditional heteroscedasticity models. The Econometrics Journal, 5(2):358–374, December 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Tsyawo:2023:FIR

[Tsy23] Emmanuel Selorm Tsyawo. Feasible IV regression without excluded instruments. The Econometrics Journal, 26(2):235–256, May 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
26/2/235/6888009.

Tugan:2021:PVM

[Tug21] Mustafa Tugan. Panel VAR models with interactive fixed effects. The Econometrics Journal, 24(2):225–246, May 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/24/2/225/5873007.

Thompson:2019:SGA

[TW19] Brennan S. Thompson and Matthew D. Webb. A simple, graphical approach to comparing multiple treatments. The Econometrics Journal, 22(2):188–205, May 2019. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.

oup.com/ectj/article/22/2/188/5487076.

Ulrick:2008:USP

[Ulr08] Shawn W. Ulrick. Using semi-parametric methods in an analysis of earnings mobility. The Econometrics Journal, 11(3):478–498, November 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Uematsu:2019:HDM

[UT19] Yoshimasa Uematsu and Shinya Tanaka. High-dimensional macroeconomic forecasting and variable selection via penalized regression. The Econometrics Journal, 22(1):34–56, January 2019. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/22/1/
34/5317155.

Vanhems:2010:NPE

[Van10] Anne Vanhems. Non-parametric estimation of exact consumer surplus with endogeneity in price. The Econometrics Journal, 13(3):S80–S98, October 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

vonBrasch:2024:IES

[vBRV24] Thomas von Brasch, Arvid Raknerud, and Trond C. Vigtel. Identifying the elasticity of substitution with biased technical change: a structural panel GMM estimator. The Econometrics Journal, 27(1):84–106, January 2024. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/27/1/84/7285795.

Vrontos:2003:FFM

[VDP03] I. D. Vrontos, P. Dellaportas, and D. N. Politis. A full-factor multivariate GARCH model. The Econometrics Journal, 6(2): 312–334, December 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

vanderSluis:1998:CAS

[vdS98] Pieter J. van der Sluis. Computationally attractive stability tests for the efficient method of moments. The Econometrics Journal, 1(1):203–227, June 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Veraart:2011:LEL

[Ver11] Almut E. D. Veraart. Likelihood estimation of Lévy-driven stochastic volatility models through realized variance measures. The Econometrics Journal, 14(2):204–240, July 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

vanGarderen:2002:EID

[vGS02] Kees Jan van Garderen and Chandra Shah. Exact interpretation of dummy variables in semilogarithmic equations. The Econometrics Journal, 5(1):149–159, June 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

vanKippersluis:2018:BPE

[vKR18] Hans van Kippersluis and Cornelius A. Rietveld. Beyond plausibly exogenous. The Econometrics Journal, 21(3):316–331, October 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Dijk:2003:EIT

[VST03] Dick Van Dijk, Birgit Strikholm, and Timo Teräsvirta. The effects of institutional and technological change and business cycle fluctuations on seasonal patterns in quarterly industrial production series. The Econometrics Journal, 6(1):79–98, June 2003. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Wang:2020:SEG

[WC20] Xi Wang and Songnian Chen. Semiparametric estimation of generalized transformation panel data models with nonstationary error. The Econometrics Journal, 23(3):386–402, September 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/23/3/
386/5835681.

Westerlund:2018:CPG

[Wes18] Joakim Westerlund. CCE in panels with general unknown factors. The Econometrics Journal, 21(3):264–276, October 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Wang:2007:TSE

[WH07] Liqun Wang and Cheng Hsiao. Two-stage estimation of limited dependent variable models with errors-in-variables. The Econometrics Journal, 10(2):426–438, July 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Wilhelmsson:2009:VRT

[Wil09] Anders Wilhelmsson. Value at Risk with time varying variance, skewness and kurtosis — the NIG–ACD model. The Econometrics Journal, 12(1):82–104, March 2009. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Wilke:2011:REA

[Wil11] Ralf A. Wilke. A review of Econometric Analysis of Cross Section and Panel Data (2nd ed.) by Wooldridge (Jeffrey M.). The Econometrics Journal, 14(3):B5–B9, October 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Windmeijer:2019:TSL

[Win19] Frank Windmeijer. Two-stage least squares as minimum distance. The Econometrics Journal, 22(1):1–9, January 2019. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/22/1/1/5317152.

Westerlund:2022:CHF

[WK22] Joakim Westerlund and Yousef Kaddoura. CCE in heterogeneous fixed-T panels. The Econometrics Journal, 25(3):719– 738, September 2022. CODEN ???? ISSN 1368-4221 (print),

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/25/3/719/6556007.

Wang:2013:ESA

[WL13] Wei Wang and Lung-Fei Lee. Estimation of spatial autoregressive models with randomly missing data in the dependent variable. The Econometrics Journal, 16(1):73–102, February 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Wooldridge:2023:SAN

[Woo23] Jeffrey M. Wooldridge. Simple approaches to nonlinear difference-in-differences with panel data. The Econometrics Journal, 26(3):C31–C66, September 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.

oup.com/ectj/article/26/3/C31/7250479.

Wright:2010:TAC

[Wri10] Jonathan H. Wright. Testing the adequacy of conventional asymptotics in GMM. The Econometrics Journal, 13(2):205–217, July 2010. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Wu:2005:PAE

[WS05] Ximing Wu and Thanasis Stengos. Partially adaptive estimation via the maximum entropy densities. The Econometrics Journal, 8(3):352–366, November 2005. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Wen:2019:GNG

[WW19] Kuangyu Wen and Ximing Wu. A guided nonparametric goodness-of-fit test with application to income distributions. The Econometrics Journal, 22(3):207–222, September 2019. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/22/3/207/5497296.

Wu:2018:TCV

[WX18] Jilin Wu and Zhijie Xiao. Testing for changing volatility. The Econometrics Journal, 21(2):192–217, June 2018. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Wang:2023:BTU

[WY23] Xiaohu Wang and Jun Yu. Bubble testing under polynomial trends. The Econometrics Journal, 26(1):25–44, January 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/26/1/
25/6648704.

Wan:2006:FRO

[WZO06] Alan T. K. Wan, Guohua Zou, and Kazuhiro Ohtani. Further results on optimal critical values of pre-test when estimating the regression error variance. The Econometrics Journal, 9(1):159– 176, March 2006. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Wan:2007:SRL

[WZQ07] Alan T. K. Wan, Guohua Zou, and Huaizhen Qin. On the sensitivity of the restricted least squares estimators to covariance misspecification. The Econometrics Journal, 10(3):471–487, November 2007. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Xiao:1998:ACT

[XP98] Zhijie Xiao and Peter C. B. Phillips. An ADF coefficient test for a unit root in ARMA models of unknown order with empirical applications to the US economy. The Econometrics Journal, 1(2): 27–43, December 1998. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Xu:2008:BAU

[Xu08] Ke-Li Xu. Bootstrapping autoregression under non-stationary volatility. The Econometrics Journal, 11(1):1–26, March 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Xu:2014:EDG

[Xu14] Haiqing Xu. Estimation of discrete games with correlated types. The Econometrics Journal, 17(3):241–270, October 2014. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Xu:2015:TSC

[Xu15] Ke-Li Xu. Testing for structural change under non-stationary variances. The Econometrics Journal, 18(2):274–305, June 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Xu:2021:POF

[Xu21] Ruonan Xu. Potential outcomes and finite-population inference for M-estimators. The Econometrics Journal, 24(1):162–176, January 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
24/1/162/5873407.

Yang:2002:LLM

[Yan02] Minxian Yang. Lag length and mean break in stationary VAR models. The Econometrics Journal, 5(2):374–387, December 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Yang:2012:BPE

[Yan12] Jingjing Yang. Break point estimators for a slope shift: levels versus first differences. The Econometrics Journal, 15(1):154– 169, February 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Yan:2022:FAF

[YC22] Yayi Yan and Tingting Cheng. Factor-augmented forecasting regressions with threshold effects. The Econometrics Journal, 25(1):134–154, January 2022. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/25/1/134/6212225.

Yoshida:2002:ESF

[YG02] Atsushi Yoshida and Alessandra Guariglia. Estimating saving functions in the presence of excessive-zeros problems. The Econometrics Journal, 5(2):435–456, December 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Yuan:2023:SPI

[YLW23] Meng Yuan, Pengfei Li, and Changbao Wu. Semi-parametric inference on Gini indices of two semi-continuous populations under density ratio models. The Econometrics Journal, 26(2):174–188, May 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X

(electronic). URL http://academic.oup.com/ectj/article/
26/2/174/6840233.

Yu:2001:GAC

[YP01] Jun Yu and Peter C. B. Phillips. A Gaussian approach for continuous time models of the short-term interest rate. The Econometrics Journal, 4(2):210–224, December 2001. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). See corrigendum [PY11].

Yamamoto:2013:ETM

[YP13] Yohei Yamamoto and Pierre Perron. Estimating and testing multiple structural changes in linear models using band spectral regressions. The Econometrics Journal, 16(3):400–429, October 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Yang:2008:GLT

[YT08] Zhenlin Yang and Yiu-Kuen Tse. Generalized LM tests for functional form and heteroscedasticity. The Econometrics Journal, 11(2):349–376, July 2008. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Yang:2011:FBA

[YV11] Jingjing Yang and Timothy J. Vogelsang. Fixed-b analysis of LM-type tests for a shift in mean. The Econometrics Journal, 14(3):438–456, October 2011. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Yu:2013:ATR

[YZ13] Ping Yu and Yongqiang Zhao. Asymptotics for threshold regression under general conditions. The Econometrics Journal, 16(3): 430–462, October 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Yang:2023:TWG

[YZ23] Yimin Yang and Huili Zhang. Three-way gravity models with multiplicative unobserved effects. The Econometrics Journal, 26 (3):422–443, September 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://academic.oup.

com/ectj/article/26/3/422/7153324.

Yang:2021:PKT

[YZLC21] Lixiong Yang, Chunli Zhang, Chingnun Lee, and I-Po Chen. Panel kink threshold regression model with a covariate-dependent threshold. The Econometrics Journal, 24(3):462–481, September 2021. CODEN ???? ISSN 1368-4221 (print), 1368-423X (elec-

tronic). URL http://academic.oup.com/ectj/article/24/3/
462/6027470.

Zaman:2002:MLE

[Zam02] Asad Zaman. Maximum likelihood estimates for the Hildreth– Houck random coefficients model. The Econometrics Journal, 5 (1):237–262, June 2002. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Zhang:2004:SED

[ZfL04] Wei Zhang and Lung fei Lee. Simulation estimation of dynamic discrete choice panel models with accelerated importance samplers. The Econometrics Journal, 7(1):120–142, June 2004. CO-DEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Zhang:2013:SPE

[Zha13] Zhengyu Zhang. Semi-parametric estimation of a generalized threshold regression model under conditional quantile restriction.

The Econometrics Journal, 16(2):250–277, June 2013. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Zhang:2015:IEP

[ZL15] Zhengyu Zhang and Bing Liu. Identification and estimation of partially linear censored regression models with unknown heteroscedasticity. The Econometrics Journal, 18(2):242–273, June 2015. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Zhang:2023:MSV

[ZLS23] Xiao Zhang, Xu Liu, and Xingjie Shi. Model selection for varying coefficient nonparametric transformation model. The Econometrics Journal, 26(3):492–512, September 2023. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http://

academic.oup.com/ectj/article/26/3/492/7075872.

Zhang:2012:TCT

[ZSP12] Yonghui Zhang, Liangjun Su, and Peter C. B. Phillips. Testing for common trends in semi-parametric panel data models with fixed effects. The Econometrics Journal, 15(1):56–100, February 2012. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic).

Zhu:2020:NSB

[ZSS20] Huanjun Zhu, Vasilis Sarafidis, and Mervyn J. Silvapulle. A new structural break test for panels with common factors. The Econometrics Journal, 23(1):137–155, January 2020. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:/

/academic.oup.com/ectj/article/23/1/137/5593950.

Zhao:2024:ATS

[ZW24] Puying Zhao and Changbao Wu. Augmented two-step estimating equations with nuisance functionals and complex survey data. The Econometrics Journal, 27(1):37–61, January 2024. CODEN ???? ISSN 1368-4221 (print), 1368-423X (electronic). URL http:

//academic.oup.com/ectj/article/27/1/37/7227331.

Zhu:2021:MAE

[ZZMZ21] Rong Zhu, Xinyu Zhang, Yanyuan Ma, and Guohua Zou. Model averaging estimation for high-dimensional covariance matrices with a network structure. The Econometrics Journal, 24(1): 177–197, January 2021. CODEN ???? ISSN 1368-4221 (print), REFERENCES 121

1368-423X (electronic). URL http://academic.oup.com/ectj/
article/24/1/177/5912831.